A Complete Bibliography of Publications in Mathematics of Computation, 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

    E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
               WWW URL: https://www.math.utah.edu/~beebe/

02 January 2024 Version 2.35

Title word cross-reference

(+1, −1) [Yan66, Yan68, Yan69]. (−1, 1) [Rot61]. (−∞, ∞) [Hab64]. (Mod2) [Wat62b]. (p − 1)! ≡ −1 [Pea63]. −1/2 [CT67a]. −12, 12 [CT67b]. 0 [GV68]. 1 [GV68, Lin65a, NDT69]. 1/2 [CT67a]. 10 [FC67, NDT69]. 10, 000 [Pat62]. 10, 000 < p < 15, 000 [Kra61]. $12.50 [Hou63b]. 1271 [Knu62]. $17.00 [Hou60]. 2 [KK64, SS65, Wat62a]. 200, 000 [Ste64b]. 2^{132} + 1 [Ise60]. 2^{159} − 1 [Ise61]. 2^{2p} + 1, p [Bri62]. 2^{n}(n ≤ 6) [Sha65b]. 2^{n} ± 1 [BS67b, Ise65]. 2^{p} − 1, pP rime < 10^{4} [Rie62]. 2^{p−1} ≡ (modp^{2}) [Pea63]. 2^{p−1} ≡ 1(modp^{2})forp < 100, 000 [Kra60]. 2π^{−1/2} ∫_{0}^{∞} exp(−x^{2})f(x)dx [Gal69].

2π ∫_{0}^{∞} (sin tt)^{n}dt [HKW60]. 3 [FC67, Sho60, Str61b, Str64]. 3/2 [CT67a]. 32 [CT67b]. 4 [Rob64, Wag67]. 401 [Ste64b]. 5 [New67b, Wag67]. 6 [Sho60]. 6000 < p < 7000 [KB64]. 8 [Ber68, KK64]. a, b [DJ67]. a^{p−1} ≡ 1(modp^{2}) [Rie64]. Arccosx [PG60]. Arcsinx [PG60]. Arctanx [PG60]. B(x) [Sha64c]. B_{14}(x)

[Car65]. {amj^{2}}, {am(j + τ)^{2}} [Jag64]. E

[Cod65a, Cod66]. e^{iα} [Lin65a]. E^{n} [Day66]. e^{x} [SM63]. E_{1}(x) [CT68]. Ei(x) [TM68, CT69]. Eig [Var66]. Eig3 [For64]. Erfc(x) [Ric64b]. f(z)/z [Gau66]. F_{1/2}(x) [WR63]. F_{19} [Rie63]. G [GA68, GS69, Ver65, Ver67a, Wim67]. G_{n} = 6^{2^{n}} + 1 [Rie69b]. Γ(p) [Atk68]. Γ(x) [Ric64b]. GF (p) [Col69]. H_{n} = 10^{2^{n}} + 1 [Rie69b]. I_{0}(x), I_{1}(x) [Bur62b]. I_{n}(b) = 2π^{−1} ∫_{0}^{∞} (sinxx)^{n} cos(bx)dx [Tho66]. I_{n}(b) = 2π ∫_{0}^{∞} (sin xx)^{n} cos bxdx [Fet67].

I_{n}(b) = 2π ∫_{0}^{∞} (sin xx)^{n} cos(bx)dx [MR65].

∫_{0}^{1} −ln(x)f(x)dx [And65]. ∫_{0}^{1} f(x)x^{2} dx [SH65].

∫_{0}^{b} exp(−x^{2})f(x)dx [SBG69].

∫_{0}^{∞} e^{−x}J_{0}(ηxξ)J_{1}(xξ)x^{−n}dx [Fet60b].

∫_{0}^{∞} exp(−x^{2})f(x)dx [SBG69]. ∫_{0}^{∞} x^{β}e^{−x}f(x)dx [CCJM63]. ∫_{0}^{∞} e−xJ_{0} (ηxϵ) J_{1} (xϵ) x−ndx [Fet60a].

J_{0}(x) [Ger64]. J_{n}(λ)Y_{n}(ηλ) − J_{n}(ηλ)Y_{n}(λ) [WMR67]. J′_{n}(x)Y′_{n}(βx) − J′_{n}(βx)Y′_{n}(x) = 0 [BA62]. J′_{p}(ξ)Y′_{p}(kξ) − J′_{p}(kξ)Y′_{p}(ξ) = 0 [Bau64]. k [GV68, Mit66, Cod65a, Cod66]. K_{ν}(t)/I_{ν}(t) [Rob65a]. L [Spi69, Swa68]. L^{p} [Bur67]. L_{1} [Ric64a]. L_{∞} [Ric64b]. [c^{2k} + (x − x^{−1})^{2k}]^{−1} [WG60]. ≤ 452, 479, 659 [SL68]. M = (6a + 1)2^{2m−1} − 1 [WZ68]. M′ = (6a − 1)2^{2m} − 1 [WZ68]. N [Cla62, Mul59, APR65, DKV63, Mus64, Rei69, Ber66]. n = 2 [Lev61]. N = h · 2^{n} − 1 [Rie69a]. n^{2} + a [Sha60b]. n^{4} + 1 [Glo60, Glo62, Glo65, Lal67, Sha61]. n ≡ 2(mod4) [Yan66, Yan68, Yan69]. ν [Low66]. O(h^{2}) [Swa68]. p [GM67, JC69, LD62]. P_{ν−1/2}^{(−m)}(cos θ) [Low66]. P_{ν}^{1}(cos θ) [Wil68]. P_{r}(n_{1}, n_{2}) [CG68]. ∂^{n}/∂x_{1}∂x_{2} ⋯ ∂x_{n} + a(x_{1}, x_{2}, ⋯, x_{n}) [SD65]. ∂I_{ν}(z)∂ν, ν = ±13 [EG63]. ∂∂θP_{μ}^{1}(cos θ) [Wil68]. π [Pat62, SW62]. ψ [Gla66, Moo67]. q < 10^{8} [Rie62]. Q_{x} ≡ (x^{3} + 5x)/6, x ≥ 0 [SL68]. QR [Par66]. Shi(x) [TM68]. σ_{2s} [Lin65b]. ∑_{ν=0}^{∞}(t^{ν}/[uν + m]!) [Les65]. t [AB69]. u_{xy} = f(x, y, u) [Day63]. x^{2} + y^{2} = l^{2}, y^{2} + z^{2} = m^{2}, z^{2} + x^{2} = n^{2} [LB66]. x^{2} − Dy^{4} = k [LD68]. x^{3} + y^{3} = z^{3} − d [GLS64]. x^{n} [Rot61]. x exp x^{2}erfcx [Old68]. y′ = f(x, y) [LS65a, LS66]. y′′ = fy + g [Fro62]. y^{3} − x^{2} = k [LJB66]. |x^{3} + y^{3} + z^{3}| ≤ 1 [Spo69].

-Bases [SS65]. -Determinant [Yan66]. -Dimensional [Ber66, Cla62, Mit66, Mus64, Mul59]. -Distribution [AB69]. -Function [Gla66]. -Functions [Spi69, Ver65, Ver67a, Wim67]. -Groups [JC69]. -j [Sho60]. -Matrices [Yan68, Yan69]. -Person [APR65]. -Transform [GA68]. -Transformation [GS69].

16 [Wim72].

49 [Ric65].

50th [Gau94].

A. [Hou61]. Abel [Fet64]. Abelian [PS60]. Abscissas [Cas65, Mic63, Rab60]. Academic [Ham64]. Acceleration [Wyn62]. Accuracy [GB68]. Accurate [Ger64]. Adams [BRB65]. Addendum [Fla68]. Addition [How62, Pat68b]. Additional [Ise65]. Additions [Glo62]. Additive [SS65]. ADI [LR68]. Adjacent [Joh63]. Adjoint [CK68]. Adjusting [LS66]. Aid [Kee65]. Algebra [Bes61, Kan66, Moo66]. Algebraic [Lay67, Lev61, New66, Rob64]. Algebraic-Exponential [Lay67]. Algèbre [Hou61]. ALGOL [For64, Var66]. Algorithm [Ber68, Bro69, Cha67, CT65, GH67, HL69, KM68, La 64, Lan67, McA65, Olv64, Par64a, Par68, RU68, Smi65, Sta62, Wil69, Wyn61]. Allocation [BKK63]. Alternating [DKV63, KS65, Kel69, Lee62, Pop60, Wid66]. Alternative [Ell67, Sal60a]. Amicable [AØS67, BM68, Hag69, Lee68, Lee69]. Among [Pat62, SD69]. Analogue [Thu69b]. Analogues [Bra63, DG63]. Analyse [Hou63a]. Analysis [Bar68c, Bar68d, BH62b, FL61, Mur67, Olv64, Tai67, Wyn61, Hou60, Hou63b, Ham65]. Analytic [BK62, CJ68c, DL67, Lyn68, Val68, vW69]. Annihilation [Ede66]. Anniversary [Gau94]. Antenna [Den69]. Appearance [LP67c]. Application [Bro67, McA65, PS60, Pri65, Rig64]. Applications [BF67a, Cor61, HL69, Lea66, Lev61, S.64, Smi65, SL69, Tra65]. Applied [BKK64, Gar65b, Par64b, Pri68]. Approach [Ell67, YP61]. Approximate [Dan69a, Dav67, LS65b, Pet63, RS61, Sec65, Spi67, Str61a, SS63b, Str64, Thu69a, Hou60]. Approximation [BKK63, BKK64, Bur67, Cad64, Day66, Den66, DM69, Fai64, Flo60, Fos65, GTM69, HO69, Har66, Hol69, Joh68, KR68, Klo64, LS65a, Lay64, Lea67, Mon64, Rea61, Ric69, SG68, Shi66, Sta63, Sto61, Val68, WR63, Wyn60, Shi60, Ric65]. Approximations [BP67, BH64, Bur62b, Bur63a, CK68, Cho69, Cod65a, Cod66, CH67, CT67b, Cod68, CT68, CT69, Cod69, Ell65, Ell67, FL66, FL67, FG67, Fie65, Kam61, KL68, Kre68, Lar64, Luk68, Man67, Moo67, Mou64, Old68, Osb67, Osh69b, Pri68, Ric64a, Spi61, Spi62, Spi67, SF69, WC61, Wim61, CT67a]. Arbitrary [BI65, Ber67, Del65, Hub66, La 63, WG64]. Archimedes [WGZ65]. Arctangent [Gla68].

Argument [Bar61, Mec66, vZW64, vZW65]. Arguments [GB68, TM68]. Arising [BK62, Den69, Dra61, Gau69, Kel64, WG60, vW69]. Arithmetic [Han68, LP67a, MA65, Sha62]. Arithmetical [Wag67]. Arrays [Ric64b]. Artin [Wre61]. Ascent [Mei69]. Assignments [Por63]. Assisted [Coh69]. Associated [Jor68, KG68, Mil63, Mil68a, Por63, SCF64, Wim62, Wim72]. Asymptotic [BH62a, CJ68b, DB60, LN67, NL69, Pea69, Sha64c, Ste69a, Tho65, vZW64, vZW65, Wil64, Wim67]. Atomic [OD65]. Attainable [But65]. August [Gau94]. Autocorrelation [Jag64]. Averages [DSS68]. Axisymmetric [Mac65].

Balanced [Bus69b]. Base [Ber68]. Based [But64b, LS65a, LS66, Lyn69, Pat68a]. Bases [SS65]. Basic [Par68, Sar68, Ver66a]. Basis [LB61]. Bateman [Sch63, Sha65c]. Behandlung [Ric65]. Behavior [Bar62, Pea69]. Belge [Hou63a]. Bell [LD62]. Benster [Hou60]. Berlin [Ric65]. Bernoulli [Car65, KB67]. Bernstein [Sta63]. Bessel [Wim72, AM61, Bau64, DB60, Hun64, KKI67, LL62a, LS67a, Lew69, Mac65, MM67, Mec66, Mec68, NF63, Raz65, SD69, Tun66, Van64, Wim62]. Bessel-Function [Mac65, MM67]. Best [Day66, PR63, Sec65, Shi66, Shi60]. Beta [DJ67, OM68]. Beta-Function [DJ67]. Between [Joh65, Kam69, LF68, Sha64a]. Bi [Fre61]. Bi-Variate [Fre61]. Bibliography [Str61a]. Biharmonic [Sim67]. Bilinear [Lee68]. Binary [GV68]. Biquadrates [LP66]. Bivariate [Ahl64, AB69, Amo69, Bur68, DJ61]. Bivins [Sho60]. Block [Cha67, Pri65, SW69a]. Block-Symmetric [Cha67]. Book [G.68, H.65, Ham64, Ham65, Hou60, Hou61, Hou63a, Hou63b, Ric65, S.64, Sho60]. Books [Ano60d, Ano61d, Ano62e, Ano63e, Ano63f, Ano63g]. Bose [NDT69]. Bossett [Hou61]. Bound [Gol63]. Boundary [Bat63, CK68, Fro62, Jam68b, KL68, Kre68, Low60a, Low61, Mit61, Mur67, Osh69b, Osh69a, Pop60, Ros60, Thu69a, Thu69b]. Boundary-Layer [Mur67]. Bounded [DM69]. Boundedness [Tun66]. Bounds [AH64, BLL64, Cha68b, DP65, Gol62, Kin66, Man66, McN64, Rab68, Ral62, Sch62, Sec65, Ste68a, Swa68, Var68b, Wen65]. Bridge [Dic68]. British [Gau94]. Broadening [WKG66]. Budde [Par64a]. Byrne [Pin69].

Cable [PWMD63]. Cables [WP60]. Calcul [Hou61]. Calculate [Com60]. Calculating [Ros60]. Calculation [BHK68, CT65, DJ67, Fet65, Fet67, FG62, GW69, Hun64, Mec66, Met69, OD65, PG60, Rot60, Rud66, SW62, SW63, Spi69, Str68a, TM68, Var68a, WP60, Woo68, dF66, vv66]. Calculations [KKI67, Wil63]. Calculus [Tra65]. Canonical [Par67]. Carlo [Hab66, Hab67b]. Cartesian [OB64]. Case [Fos68, Lev61, Wid66]. Cases [GS69]. Catalytic [KG68]. Cattle [WGZ65]. Cauchy [Cho68a, Gin63, Mil68a]. Central [Coo61, Sal62a]. Centre [Hou63a]. century [Gau94]. Certain [BJ60, Bri64, Bur63b, CJ68a, Coh67b, Ehr67, Ell67, Gra64, Gre60, Hun64, JLB67, Kil60, Lan60, Lan67, Lav64b, Low60b, Low61, Lyn65a, Lyn65b, May65, Mil68b, SK61, SW63, Sha64b, SCF64, Sta62, SS63b, Ver66a, WM65, Wun62]. Certification [Var66]. Change [Sol67]. Characteristic [Day63, Swi60, Tam62]. Characters [Com60]. Chebyshev [AM61, Bar61, BDN69, Bur63b, Cha68b, Cod65a, Cod65b, Cod66, CH67, CT67a, CT67b, Cod68, CT68, CT69, Cod69, Ell64, ES65, Ell65, Hum64, Mei69, Mou64, NDT69, Scr69, Shi60, Shi66, Ste68a, WC61, Woo67]. Chi [GTM69, Tho65]. Chi-Square [GTM69]. Christoffel [Gau68]. Circle [DJ61, FG62]. Circles [Mel69]. Circuit [WG60]. Circulants [CJ68a]. Circular [Cad64, DJ62, Kam69, WD61]. Class [Ayo67, Bro65, Coh62, Coo67, Fos65, Her69, Hun68, LS65a, Met69, Mil68b, Nas65, Sha67a, Sha69, Tra66, Val68, Ver65, Wim67]. Classes [JC69, MM67]. Clausen [Woo68]. Clenshaw [Cha68c]. Close [Har66]. Closed [Cha68b, LD67]. cm [Hou60, Hou61, Hou63b, Ric65, Ham64]. cm. [Hou63a]. Co [Hou63b]. Cods [McA65]. Coefficient [Sta62]. Coefficients [BF66, Car62, Cas65, Eis68, Ell64, ES65, Fie60, Jam68b, Lyn65c, Mic63, Raz65, Str63]. collaboration [Hou61]. Colloque [Hou63a]. Color [YP61]. Columbia [Gau94]. Combinations [Lot62]. Combinatorial [Che64, Erd64]. Comments [LS67c, LS67b]. Commonly [Ker69]. Comp [Wim72]. Compared [JJ65]. Comparison [DSS68, LSL64, WB64]. Complement [Wil64]. Complete [Cod65a, Cod65b, Cod66, Ise60, Ise61, OM68, SK61]. Complex [CT65, Cor61, Dav64, Dor66, GH67, Lan60, Lon60b, Lyn69, Mec66, Str65, Var68a, Var68b, Yan65]. Components [von67]. Composite [Asc60]. Compositeness [Pax61]. Compressible [Ric69]. Computable [LS67a]. Computation [AB69, Amo69, BK68, Boe60, Boe62, Coc69a, Coc69b, Fle68, GH67, Gar69, Gau61, Gau66, JC69, Jam68a, KB67, LWF72, Man61, Mec68, Ric64a, Rob65a, Sha62, Sie65, SD65, Str65, Swe63, Wil64, WS66, Gau94]. Computational [BKK63, Kan66, YP61, Gau94]. Computations [AØS67, Jen62]. Computed [Gol62]. Computer [Bau60, Chi69, Coh69, Fra69, Gre61, Hay65, JJ65, Kee65, LB61, Mit61, PS60, SS68, TS61]. Computer-Assisted [Coh69]. Computers [Bar62, Lev61]. Computing [APR65, BI65, Col69, DJ62, Gar65a, Lon60b, Map63, Mar68, SM63, Sch62, Van69, WB64, Wil67]. Concerning [BH62a, Bri62, DGL66, Eps62, Jag65, Mil68b, Sha63a, WS66, Wre68, Sha60b]. Condensation [Dan69b]. Condition [CJ68a, Dur66, Fro62, Gau69, Pea65]. Conditional [Klo64]. Conditionally [Kro67]. Conditions [Bat63, CK68, Eis68, Low60a, Low61, Osh69a, PD64]. Conduction [Low60a, Low61, MP63]. Conductors [Kam69]. Cones [Str61b]. Configurations [Gol69]. Confluent [Fie67]. Conformal [BC67]. Congruence [BF66, Kra60, McA65, Rie64]. Congruences [Pea63, WM65]. Congruential [Bar62]. Conjecture [LP67b, Sha60b, Sha63a, Wil63]. Conjectures [Rob65b]. Conjugate [Mag60]. Connected [BC67]. Consecutive [BLL64, Coh67a, KK64, LP67a]. Considered [Low66]. Consistency [Eis68]. Constant [BK64, Chu65, Knu62, LFiCM65, Mel68, Mel69, Mit68, Sha67b, Spo69, Swe63, Wre60, Wre61, WS66]. Constants [Sha64b, Tre66]. Constrained [Kin66]. Construction [BP67, Dav67, Gau68, Lyn65e, Swi60]. Containing [BS67a, NF63]. Continued [Fra69, GH67, Spi61, Spi62, WS66]. Continued-Fraction [Spi62]. Continuity [Rab68]. Continuous [SL61]. Contour [LD67, Woo66]. Controlled [Bau60, Rom64]. Convective [RW66]. Convergence [BHT69, Bro64, But66, Cho69, Gar67, Lee62, LR68, New67a, Par68, Wid66]. Convergent [Den69, Tra66]. Convergents [SM63]. Converging [Lea67]. Conversion [MA65, Sch67]. convex [Coc69b]. Convolution [BKK64, Jon61]. Coordinates [OB64]. Corner [Wig69]. Correct [Bat63, DG63]. Correcting [McA65]. Correction [Sat66]. Corrections [Cla62, Pop60]. Corrector [Cas69, Coo61, KM68]. Correctors [Kro67, Ste65]. Correlation [Gre61, Jen62]. Correlations [KF62]. Corrigenda [Ano60a, Ano60b, Ano61a, Ano61b, Ano61c, Ano62a, Ano62b, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano65a, Ano66a, Ano67a, Ano67b, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Cod66]. Corrigendum [Ano60c, Ano62c, Ano62d, Ano64b, Ano64c, Ano65b, Ano65c, Ano66b, Ano67c, Ano69d, CT67a, Ste71, Wim72]. Cosine [Tha61]. Coulomb [Boe69]. Counterexample [LP67b]. Coupled [Dal63, PD61, Tre66]. Coverage [DJ62]. Criteria [Bro64, Low61, PD64, Rie69a]. Cross [Bau64, LL62a]. Cross-Product [LL62a]. Cryptography [Lev61]. Cubes [LRB69]. Cubic [LLMS62, Mag60]. Cumulative [LWF72]. Cunningham [Glo62]. Curtis [Hou60, Cha68c]. Curve [Bur62a, Smi65]. Curve-Fitting [Smi65]. Curves [PR63, Rea61, Sat66, Sto61]. Cusps [Atk68]. Cyclic [Rig64]. Cyclotomic [BF67a, Met69, Rob65b]. Cylinder [Dor66]. Cylindrical [MP63].

D [Hou60, Pin69]. Data

[Bur62a, Jen62, Mon64, PR63, Sha63a]. Davidon [Bar68a]. Davidon-Like [Bar68a]. Davis [LLS65]. Dawson [Hum64]. Decimals [SW62]. Decision [MTK62]. Decomposition [Cho68a, Par67]. Decompositions [de 64]. Deferment [Sie65]. Defined [Hol69, Hun68, Moo66, Wyn60]. Definite [LSL64, Van64, vH63]. Deflection [Dic68]. Degree [Bur66, RR68, Str60a, Str61b, Str64, Str66]. Derivative [CK68]. Derivatives [Gau66, HW60, Joh68, Mou64, Pat69, Sha65a, Smi65, Str60b]. Deriving [Val68]. Descriptions [Ano60d, Ano61d, Ano62e, Ano63e, Ano63f, Ano63g]. Designs [Yan66, Yan68, Yan69]. Determinant [Yan66]. Determinants [Lav64b, PRS63, Yan65]. Determination [Bro69, Gre61, Olu65]. Deterministic [Fra63]. Development [S.64]. Diagonal [Bus69b, La 63, SM63]. Diagonalizing [La 64]. Diagonally [Pri65]. Difference [AZ69, Ape68, BF67a, BK62, Bra63, Bra66, CK68, Car69, Dal63, DG63, Eis68, FG67, Fra65b, Gor68, GM68a, GM68b, Gra64, H.65, HW61, Hay65, Hub66, Jam68a, Kel64, KL68, Kre68, Lar64, Lay64, Lee66, Low61, MP63, Osb62, Osb67, Osh69b, Osh69a, PM62, Pea69, Pri68, Rig64, RW66, Ros64, Sim67, SW69b, Thu69a, Tun66, Wid65, Yam67, vW69]. Differences [AH64, Bur63c, Del66, Eps62, Gar65a, LP67c, Lay67, Sal62a]. Different [Tre66]. Differentiable [SL61]. Differential [BP67, BK62, BKK64, BC69, But66, Coo67, Coo68, Day64, DGL66, FG67, Fra65b, G.68, Gar65a, Gea67, Jam68b, Kat68, KM68, LS67c, ML67, Mur67, Nor62, Osb62, Osb67, PD61, Ros64, Sha66, Ste69a, Tre66, Wen65, Wit64, vW69, LS67b]. Differential-Difference [BK62, vW69]. Differentiation [Lyn68]. Diffusion [HW61, Lar64, Pea65]. Digamma [Moo67]. Digital [CEFT62, Kee65]. Digits [Pat62]. Dimensional [Asc60, Ber66, Cla62, Gin63, Hub66, Lyn65b, Man66, Mit66, Mul59, Mus64, RR69, SS64]. Dimensions [Bes64, Bra63, Mil60b, Mil60a, Mil60c]. Diophantine [GLS64, LJB66, LB66, LD68, Lee68]. Dirac [CT67a, CT67b]. Direct [de 64]. Direction [DKV63, KS65, Kel69, Lee62, Wid66]. Dirichlet [Bra63, BHT69, Eps62, Kam69, SW63, Sha64b, Spi69, Wig69]. Disc [BS67a]. Discontinuity [Nav63]. Discontinuous [Ape68]. Discrete [BP67, BHT69, Bur62a, Cho69, Man66, Thu69b, Tra65]. Displacements [Pet65]. Dissipative [Osh69b]. Distributed [Geb64]. Distribution [AB69, Amo69, BH62a, Cad64, DJ61, Geb64, GV68, Jag64, PS67, SK67]. Distributions [Kon65, Kra66]. Divisibility [Lee69]. Division [JJ65, Moo66]. Divisors [Ehr67, Kra61, Kra66, Mus66, SK67]. DM [Ric65]. Domain [Mil60b, Mil60a, Mil60c]. Domains [BC67, Coh62, Coh65, Coh67b, Coh69]. Dominant [Pri65]. Donald [G.68]. Doppler [WKG66]. Double [BC69, Lan67, Lin65a]. Douglas [Fai69]. DuFort [Gor68]. Dunod [Hou61]. Dynamic [BKK63, Ste64a].

E. [Pin69]. Eberlein [Sar68, SS69]. Economical [Fyf66]. Economization [Bur62a]. edition [Hou60]. Editor [Hou63b, G.68]. Edward [Ham64]. Effect [Kro67]. Efficiencies [Gre67]. Efficient [Cha67, DJ67, Spi61, Woo68, WS66]. Eigensystem [Var68b]. Eigenvalue [BH62b, Car69, Coo61, Dra61, For64, Gar65b, Gea69, God66, Goo65, Osb62, Par64b, Var66]. Eigenvalues [BK68, FC67, Gar69, Gar65a, Gol62, LF68, Olu65, Swa68, Wen65, WB64]. Eigenvectors [FC67, LF68, Var68a, WB64]. Eighteen [BF67a]. Eighteenth [SO60]. Einstein [NDT69]. Elastic [PWMD63]. Elasticity [Van64]. Electronic [PS60]. Electrons [Gol69]. Elementary [Erd64, Joh65, Spi61]. Elements [Coc69a, Ham65]. Elimination [Bar68b, Sha63b, Wil67]. Ellipse [Cad64]. Ellipses [LS65b]. Ellipsoids [Fis64]. Elliptic [Boe61, BH64, Cod65a, Cod65b, Cod66, FL67, Fet65, Fra65a, Jam68a, Lan60, LL62b, Lin60, LT64, Lin65a, Lin65b, Luk68, NC66, Van69]. End [Pea65]. Epsilon [Wyn61]. Equal [Bar66, Bru69, LP66, LPS67]. Equally [Mil60b]. Equation [Asc60, BK62, BC69, Bra63, CK68, Cha68a, Coc69a, Con67, Coo61, DG63, Fai64, FL66, FG67, GLS64, Gin63, Gor68, LJB66, LD68, Lar64, Lee68, Low60c, Low60d, Mit61, MP63, PM62, Pea65, Pea69, Rad60, RS61, Sat66, Voi66, vW69]. Equations [Ape68, Bat63, BKK64, Ber67, BF66, Bra66, Bro65, But66, Cho68b, Cho69, Coo67, Coo68, Dal63, Day64, Day67, DGL66, Dic68, Fet64, FL61, Fra65b, Gar65a, Gea67, HW61, Han68, Hub66, Jam68b, Jar66, Jon61, Kat68, Kel64, KM68, LB66, Lee66, LS67c, Lin69, Lon60b, ML67, Mur67, Nas65, Nor62, Osb62, Osb67, Osh69a, PD61, Pet63, Pet65, PRS63, Rig64, RBM67, Rob64, Ros60, Ros64, Sal60b, Scr69, Sha66, Sta62, Ste69a, Tai67, Tra66, Tre66, Wal69, Wit64, LS67b, G.68]. Equidistant [Mon64]. Equidistribution [Fra64]. Erlang [Whi63]. Errata [Ano61e, Ano63h, Ano63i, Ano63j, Ano65d, Ano65e, Ano66c]. Erratum [Ano60e]. Error [AH64, Bar68c, Bar68d, Cas65, CJ68b, Cha68b, Cha68c, CJ68c, CR68, Cod69, Gau61, Har66, Jor68, Kin66, LR68, Lyn65c, Lyn64, Man67, McA65, McN64, NL69, Olv64, Osb67, Rab68, Ral62, Rei67, Sch62, SG68, Sec65, Sim65, Ste68a, Str65, Str68a, Tai67, vZW64, WD61, Wil64, Woo67, H.65]. Error-Bounds [McN64]. Error-Correcting [McA65]. Errors [Des63, Ell65, Ell67, LLS65, Low60b, Low60c, Low60d]. Essentially [BHT69]. Estimate [Sim65]. Estimates [BHT69, Bra66, CJ68b, Cha68c, CJ68c, ES65, Kan66, Rei67, Whi65]. Estimating [HK68, LLS65]. Estimation [Ell64, Man67, Rig65]. Euler [Knu62, KB67, LP67b, McN64, Nav63, Sha67a, Swe63]. Evaluating [Hir68, Whi63]. Evaluation [Bar61, CR68, CCJM63, Ell64, Fra65a, Fyf66, Gal69, Gla66, Gre60, Hir68, Hun68, KG68, KM68, KM67, Lan60, LL62a, Lav64b, Lin60, LT64, Lin65a, Lon60a, McN64, MR65, Mel68, Mit68, NC66, SS64, SO60, Sta63, Tak66, Tho66, Van64, Wre60, Wre61]. Evaluations [Lyn65b, Sha66]. Even [Lee69]. Every [SL68]. Exact [BF66]. Existence [Bur63b, Jam68b, Wun63]. Expansion [AM61, DB60, Ell64, ES65, LC61, NDT69, Sha64c, Tho65, Van69, Ver66c, vZW64, vZW65, Wyn60]. Expansions [Boe69, Cod65b, FW61, Fie67, Fra69, Hol69, Hum64, LL62b, LW63, LN67, NL69, OB64, Tha61, Ver65, Ver66a, Ver66b, Ver67a, Wim62, Woo67, Wim72]. Experimental [LB61, PR63, SS65]. Experiments [Jor68, KS63, Ric66]. Explicit [Asc60, Atk68, CJ68a, Ker69, Lar64, Low61, Lut68, MP63, Swa68]. Exponential [CT68, CT69, Cor61, Del66, Lay64, Lay67, Led61, LD62, vZW65]. Exponentials [Ric64a]. Exponents [Ste64b, Tam62]. Expository [Str61a]. Expressions [BP67]. Extended [BA62]. Extension [Hea65, Lyn65d, Sri69]. Extensions [HL69, New69, RU68, Str68b]. Extrapolated [Pet65]. Extrapolation [Let66]. Extremal [Bus69b].

Factor [Glo62, Mag60, Met69, Rie63, Rig65].

Factorization [Ise60, Ise61, McE69].

Factorizations [Bri63, BS67b, Ise65]. Factors [BJ60, Bri64, Glo60, Kar61, Rie62, Rie69b, Wra64]. Fairing [TS61]. Fast [Map63, Ber67, Ber68]. Features [Coh67b, GC68]. Fermat [Pax61, Rie63, SH64, Wra64]. Fermi [CT67a, CT67b, WR63]. Fibonacci [Geb67, Wun63]. Field [Met69, Yan65]. Fields [Ayo67, Coo61, McE69, Swi60]. Fifth [Str66]. Figurate [Wun62]. Figures [Ger64, GV68]. Filon [Fos68, Tuc67]. Filon-Trapezoidal [Tuc67]. Finding [BS67a, Ste69b, Ste71]. Finite [AZ69, AH64, Bra63, Car69, Eps62, Gar65a, GM68a, GM68b, Gra64, Jam68a, Lav64a, Lay64, Lay67, Lew69, Lon60a, McE69, Osb62, Osb67, PM62, Pea69, Pri68, Ros64, Sim67, SS68, SW69b, Thu69a, Yam67]. Finite-Difference [AZ69, Car69, GM68a, GM68b, Pea69, Sim67, SW69b, Thu69a, Yam67]. Finitely [Kam61]. First [FL67, Ger64, KK64, LP67c, LD62, Low66, Met69, Pat62, Tre66, Van69]. First-Order [LD62, Tre66]. Fit [PR63]. Fitting [Bur62a, Smi65]. Five [SL68]. Fixed [SS63a]. Floating [LS67c, LS67b]. Floating-Point [LS67c, LS67b]. Floors [Coh65]. Flow [BHK68, DG63, KG68, Ric69, von67]. Fluid [PWMD63, Ric69]. Flux [von67]. Folding [Lun68]. Ford [S.64]. Form [Bus69a, Glo65, La 63, Lal67, SL68, Sha60b, Sha61, WG64]. Formal [Hol69]. Formally [Wyn60]. Forms [WZ68]. Formula [Ahl64, BH62a, Bau60, Cas65, CJ68b, Cha68b, Coo61, Fos68, Jag66, Lut68, McN64, Sch67, Sri69]. Formula-Controlled [Bau60]. Formulae [And65, Fyf66, LS65a, McN64, Pat68b, Pat68a, Pat69, Sar68, SS69, Wim68]. Formulas [Bes64, Boe61, Bur63b, Bur63c, Bus68, But64b, CJ68c, Dav64, Dur66, Gau68, Hab64, Hab67a, Hab69, HW60, Har62b, LS65b, Lyn68, Mil63, New63, RR69, RS61, Sal60a, SK61, Sal62c, Sec65, Sha65a, Sta63, SS63a, Str60a, Str61b, SS63b, Str63, Str64, Str66, Str68b, Str60b, Wil69, Wyn61]. Formulation [GM68c]. Four [Bra63, LRB69, YP61]. Four-Color [YP61]. Fourier [Ber67, Ber68, CT65, Rud66, Tun66, WG60, dF66]. Fourth [Blu62, Bra63, FG67, Jar66].

Fourth-Order [Blu62, Bra63, FG67]. Fraction [Fra69, GH67, Spi61, Spi62]. Fractions [WS66]. Frankel [Gor68]. Franklin [Uch66]. Free [Ros60]. Fresnel [Boe60, Cod68, Fle68]. full [Coc69b]. Function [Bar61, Bro67, Car62, CR68, CH67, Cod69, DL67, DB60, DJ62, DJ67, Ell64, ES65, GB68, Gau61, Geb64, Gla66, Gla68, HP62, Har66, Hea65, Hum64, Joh68, Lal69, LS67a, Leh60, Leh66, Lin60, LT64, Lin65b, Low66, LW63, Lyn65b, Lyn69, Mac65, Man66, MM67, Moo67, Mou64, NF63, Old68, OM68, PS67, Rei69, Ric69, Rot61, SL61, SG68, Shi66, Spi66, Spi68, SD65, Str65, Str68a, Tak66, vZW64, WC61, Wil64, Woo67, Wre68, Shi60]. Functional [Cha68a, Nav63, NL69, Spi67]. Functionals [Dan69a]. Functions [AM61, Bau64, Boe62, Boe69, Bur62b, Bur63a, Bur64, Bur68, CJ68c, DM69, Dor66, Ell67, EG63, FW61, Fie65, Fle68, GH67, GS62, Hol69, Hun64, Hun68, Jag64, Kam61, KF62, KKI67, Kir60, LL62a, Led61, Lew69, LL62b, Lin65a, Lon60a, LNP69, LC61, Lyn68, Mec66, Mec68, Mil63, MTK62, Nav63, NDT69, OD65, PR63, Rab68, Rei67, Ric64a, SD69, Sha66, Spi61, Spi62, Spi69, Tam62, Tra66, Van64, Val68, Ver65, Ver66a, Ver66b, Ver66c, Ver67a, Ver67b, Whi63, Wim62, Wim67, Wim68, Wyn60, vH63, Wim72, Wim72]. Fundamental [Coh65, Coh67b, Coh69, Sim67]. Funktionen [Ric65]. Further [NL69, Wre60].

G [S.64]. G. [Pin69]. Galois [Coc69a, Swi60]. Games [APR65]. Gamma [Bar61, CH67, GB68, Tak66, WC61, Whi63, Wre68]. Gap [Atk68]. Gaps [Joh65, Sha64a]. Gases [S.64]. Gauss [Cas65, CJ68b, Cha68b, CJ68c, CCJM63, Fet65, Gal69, Gau68, GW69, KS63, McN64, Pat68a, Sha69, Sha63b, SH65, Ste68a, Whi65]. Gauss-Type [McN64, Whi65]. Gaussian [And65, Bar68b, Cad64, Day63, DJ61, Pin69, Rab68, RS61, Rot61, Sha60a, SCF64, SS63a, SBG69]. Gauthier [Hou63a]. Gauthier-Villars [Hou63a]. Gelbard [Pin69]. General [DJ61, Gar65b, Kel64, Nas65, Pet63, Var68a, Var68b, Wil69]. Generalisation [Fai69]. Generalised [Ver67b]. Generalization [Ahl64, Her69, Ste68b]. Generalizations [Sha67b]. Generalized [BI65, BI66, Bur64, Cas65, Fie65, GA68, HP62, Lav64a, LW63, Sha67a, SCF64, SW69b, Tra65, WD61, Wil67]. Generated [Bar62, Gre61, Tau65]. Generating [Cla62, Geb64, Geb67, Mul59]. Generation [Bow67, BF67b, How62, JJ65, Joh63, Lot62, Wel61]. Given [BS67a, Map63, Shi66, Shi60]. Global [Par68]. Globally [Tra66]. Golomb [Mit68]. Goursat [AH64]. Gradient [Gre67]. Gram [Ric66]. Graph [May65]. Graphs [Bow67, S.64]. Gravity [Mac65]. Greatly [Tre66]. Green [Man66]. Greenspan [G.68]. Groboillot [Hou61]. Ground [Kam69]. Group [Coc69a, Com60]. Groups [Bro69, JC69, PS60, Sha65b, SS68]. Growth [Jor68]. Günter [Ric65].

Hadamard [BH65]. Half

[DJ67, Lin60, LT64, Lin65a, Swa68, Gau94]. Half- [Swa68]. half-century [Gau94]. Half-Integer [DJ67]. Hall [Sha65b]. Hardy [Sha60b, Sha63a, Sri69]. Harmonic [Man66, Mil60c, OB64]. Harmonics [Rot60]. Hasse [ZL69]. Having [Wal69]. Heat [Asc60, DG63, Dra61, Gin63, Low60a, Low60c, Low61, MP63]. Helpful [Bes64]. Henrici [H.65, Ham65]. Hermite [Ahl64, SCF64]. Hessenberg [Bus69a, Par67, Par68]. Heuristic [BH62a]. High [CEFT62, DG63, GB68, Lev61, Lyn65b, Mit61, PG60, Rab60, RS61, RR68, SS64, TS61]. High-Dimensional [Lyn65b, SS64]. High-Order [DG63, RS61]. High-Speed [CEFT62, Lev61, Mit61, TS61]. Higher [BLL64, GH67, Hou60]. Hilbert [Coh65, Coh67b, Coh69, DE62, FC67, Lav64a, SL69]. Hill [Hou63b, Sat66]. Hille [Sri69]. Homogeneous [PR63]. Horn [Sch63, Sha65c]. Householder [HL69]. Howland [Nel61]. Hundred [Ger64]. Hydrodynamics [Van64]. Hyman [Gar65b]. Hyperbolic [AZ69, Ape68, GM67, GM68a, GM68b, GM68c, Gra64, Tha61, Yam67]. Hypercircle [Val68]. Hypercubes [Lyn65c, Lyn65d, Lyn65e]. Hypercubic [Lyn65a, Lyn65b]. Hypergeometric [FW61, Fie65, LC61, LW63, Ver66a, Ver66b, Ver66c, Ver67b, Wim68]. Hyperosculatory [Sal60a, Sal62c]. Hypersphere [Mit66]. Hypothesis [GC68].

Ideal [Ste64b]. Identities [Che64, Mac65]. Identity [Bar68b]. ihre [Ric65]. II [Sha65c, BW67b, Bar68d, Bri64, GM68b, Hab67b, Lyn65d, SS69, Spi68, Yan69]. III [Del65, Lyn65e]. Imaginary [Ayo67, Bar61, LS67a, vZW64]. Immersed [PWMD63]. Implicit [But64a, SW69a, Wid66]. Improved [Com60, Coo61, Dic68, SK61, vZW64, vZW65]. Improvement [Mec68]. Impulsive [Pea65]. Incidence [Nik60]. Incomplete [Bar61, DJ67, FL67, OM68, Tak66, Van69, Whi63]. Index [LWF72]. Inequality [Val68]. Infinite [Bur63c, Har62b, LW65, LW63]. Infinity [Fro62]. Initial [Ape68, Day63, Kre68, LR68, Osh69b, Sim65]. Initial-Boundary [Osh69b]. Instability [Bar68a]. Integer [Bar68b, DJ67, SL68]. Integer-Preserving [Bar68b]. Integers [Coh67a, GV68, LD62, Moo66, Rob65b, Wag67, de 64]. Integral [Cad64, CT68, CT69, Coh62, Cor61, Day67, Fet67, Fra65a, GTM69, Hun68, Jon61, LS67a, Lin69, Mec68, MR65, NF63, Raz65, Rei69, RS61, SD69, Scr69, Voi66, vZW65, WR63, Wim61, Woo68, vH63]. Integrals [Boe60, Boe61, CR68, Cod65a, Cod65b, Cod66, CT67b, Cod68, FL67, Fet65, Fle68, Fos65, Gau61, GS62, Gla66, Gla68, Gre60, HO69, Har62b, Hir68, KKI67, KM67, Lan60, Lew69, LW65, Lon60a, Lor60, Luk68, McN64, MM67, Mil63, NC66, Nel61, OD65, Ric69, SS64, SBG69, Tho65, Tho66, Van64, Van69, Wil64, WKG66, Woo66, Woo67, dF66, vv66, CT67a]. Integrand [HW60, Mil60c, Str60b, vv66]. Integrating [Bes64]. Integration [Bar68c, Bar68d, But64b, DJ61, Feu68, FL61, Gea67, Has61, Led61, LS65b, LS67c, Low60c, Lyn65a, Lyn65b, Lyn65c, Lyn65d, Lyn65e, LD67, McL63, ML67, Mus64, New63, Nor62, Pat68a, Pat69, Rab68, RR69, RBM67, Sec65, Str60a, Str61a, Str61b, SS63b, Str64, Str66, Str68b, Tre66, Wit64, LS67b]. Integro [PD61]. Integro-Differential [PD61]. Interpolant [LS66]. Interpolation [Ahl64, Bur68, Coo68, Fre61, Kah69, Lay67, New66, Sal60a, Sal62b, Sal62c, Smi65]. Interscience [Hou60]. Interval [Fre61, Hab64, Han68, Rot61, Shi66, Shi60]. Intervals [Bur63c]. Introduction [Ham64]. Invariant [Par66]. Inverse [BI65, Sal60a, Str68a, Tha61, Var68a, Wil67]. Inverses [CJ68a, Col69, Ker69, Lav64a]. Inversion [BI66, Ber66, DM62, Ede66, Gau69, Low60b]. Inverting [Cha67, Geb64, LSL64, May65]. Investigation [Chi69, Hay65, Jag66]. Involving [Gla66, HW60, Lew69, Pat69, Str60b, Van64, Ver66b, Ver67a]. Irreducibility [Car65]. Irreducible [Swi60, Sha65c]. Isomorphism [JC69]. Isoperimetric [Mel68]. Iterated [Ric69, Tha61, Wyn62]. Iterates [Lal69]. Iteration [DKV63, Sen64, Tra66, Var68a]. Iterations [Ber68]. Iterative [AZ69, BI65, BI66, Des63, Dor69, God66, Jar66, Mar68, Nas65, Pet63].

J [Hou61, Sho60]. J. [Uch66]. Jacobi [LW63, Pet65]. Jacobian [LL62b]. James [Sha65b]. John [G.68, Sho60, Hou63b]. Johnson [Hou61]. Joint [Jag64]. Jr [Sha65b]. Jr. [Sho60].

Kantorovich [Hou60]. Kelvin [Bur63a]. Kernel [APR65, Jam68a, KG68]. Kernels [Tun66]. Khintchine [Wre60, WS66]. Kind [Boe61, DB60, Fet65, Fra65a, LS67a, Lin69, Mec66]. Kinds [FL67, Van69]. Korganoff [Hou61]. Krylov [Hou60]. Kutta [Blu62, But64a, But65, Fyf66, HJ64, Kin66, Lut68, Ral62].

L [Ham64, Hou61]. L. [Hou60]. Laguerre [Bur63b, Cas65, CCJM63, GW60, GS62, Lor60, Mil63, Par64b, SCF64]. Lal [Sha67b]. Laminar [KG68]. Landau [SS66, Chi69]. Laplace [Ber66, Bus68, Gau69, Ver65]. Large [JLB67, LS67a, LSL64, Wim67]. Largest [Gra63]. Latent [DE62]. Latin [OV61]. Lattice [FG62, KS63, Mit66, Spo69]. Lawson [RU68]. Lax [GM68c, Yam67]. Layer [Mit61, Mur67]. Least [DSS68, HL69, Jor68, Klo64, Sie65]. Least-squares [Jor68]. Legendre [Lew69, Mil63, New69]. Lehmer [Ste71, Ste69b]. Lemniscate [Car62]. Length [Ber67, Coc69b]. Lens [Lea67]. Lens-Shaped [Lea67]. Less [Map63, Rob64].

Level [Lee66]. Like [Bar68a, Bar66, Bru69, LPS67]. Limit [Map63]. Limiting [GS69, Man67]. Limits [Lyn65b]. Line [Dor69, Gar67, Rea61, Sal62a, Sto61]. linéaire [Hou61]. Linear [BP67, BF66, Dal63, Day64, HL69, Jor68, Kan66, Lee66, Mil68a, Pet63, PRS63, S.64, Sch62, Ste69a, Tau65, vH63]. Lines [Kam69, TS61]. Liouville [BK68, Car69, Leh60]. Littlewood [Sha60b, Sha63a]. Lobatto [Mic63, Pat68a, Rab60]. Locally [Hub66]. Locating [DL67]. Logarithm [CH67, KR68]. Logarithmic [Spi62]. Logical [Bau60]. Lommel [Boe62]. Low [Rab68]. Low-Order [Rab68]. Lower [Gol63]. LR [Gol62]. Lucas [KB64]. Lucasian [Rie69a].

M [Pin69]. Machine [CT65, LLMS62, Ste64a, Var68b]. Maclaurin [McN64, Nav63]. Maehly [SM63]. Majority [MTK62]. Mann [OV61]. Mantissa [Kon65]. Manuel [Sho60]. Map [Lun68]. Map-Folding [Lun68]. Mapping [BC67]. Marshall [Sha65b]. Mass [Dra61]. Math [Wim72]. Mathematics [LWF72, Gau94, Gau94, Ham64]. Mathématiques [Hou63a]. Mathieu [BC69, Kir60, Tam62]. Matrices [BH65, BI66, Bus69b, Cho68a, DM62, FC67, Gea69, Gol62, Her69, Hof65, KS65, Ker69, LF68, LSL64, LNP69, Low60b, May65, Mil68b, Nik60, Par66, Par67, Par68, Pri65, Pri68, Sen64, Sha63b, Yan68, Yan69]. Matrix [BI65, Bes61, Bes68, BH62b, Bus69a, Cha67, DE62, DP65, Ebe64, Ede66, Fra64, Gar69, Gau69, Gol63, La 63, La 64, Lav64a, Mar68, MM66, PS60, Par64b, Pet65, Por63, Sch67, Var68a, Var68b, Wal69, WG64, WB64, Wyn62]. Matrix-Power [Fra64]. Max [Por63]. Maxima [CG68]. Maximal [Joh65, Mar68, Sha64a, Yan66, Yan68, Yan69]. Maximization [Bur66]. Maximum [Gor68, Yan65]. McGraw [Hou63b]. McGraw-Hill [Hou63b]. Means [Fet65, PRS63]. Measure [Sar68, SS69]. Mechanical [Sar68, SS69]. Meijer [Ver67a]. Meinardus [Ric65]. Meissner [Sat66]. Merits [SM63]. Mersenne [BJ60, Bri64, Ehr67, Gil64, Hur62, Kar61, Kra61, KB64, Kra66, Rie62, SH64, SK67]. Mervin [Cla62]. Mesh [Kel64]. Method [AZ69, APR65, Bar62, BI65, BI66, Blu62, Bro64, Cla62, DSS68, Day63, Day64, Day67, DL67, DJ62, Dic68, Fai69, Fle68, Gar65b, God66, Gol62, GM68c, Has61, Hol69, Kel69, LS66, Lay64, Lin69, LLS65, Lon60a, Lon60b, Lyn64, Mag60, Map63, Mei69, Mul59, OD65, Par64b, Pet63, Pet65, Pop60, Ros60, Sie65, SF69, Ste69b, Ste71, Str65, Tre66, Van69, Wen65, Wid66, Wig69, Wil67, Wit64]. Méthodes [Hou61]. Methods [Bar68a, BRB65, Bro65, Bro67, But65, Car69, Com60, Coo67, Coo68, DKV63, Fra65b, GM67, GM68a, GM68b, Gre67, Hou60, HJ64, Jam68a, Jam68b, Jar66, Kin66, Lea66, Lee62, Let66, LSL64, LR68, Lyn69, ML67, MP63, Nas65, Osb62, PM62, Ral62, SW69a, Thu69a, WB64, H.65]. Metropolis [Sho60]. Mid [Jag66]. Mid-Point [Jag66]. Midpoint [Hab67a, Ste68b]. Miller [Olv64]. Min [Por63]. Minimal [Bow67, Con67, RR69, Coc69b]. Minimax [Den66]. Minimisation [Bro67]. Minimising [Osb67]. Minimization [Dan69a, PD64]. Minimum [BW67a, BW67b, BS67a, Fis64, Kin66, LD62, Ral62]. Miscellaneous [Bri63]. Mises [Mit61]. Missing [Jen62]. Mixed [Bat63, Kre68, Lay67, Low60a, Osh69b, Sch67, Thu69a, Thu69b]. Möbius [GC68]. mod [Wat62a]. Models [Coh69]. Modification [Blu62]. Modified [DB60, Hab66, Hab67b, Jag66, Kir60, LS67a, Mec66]. Modular [Man61, Sch67, Sha62]. Modulo [Fra64, LD62, New67b, Tau65]. Modulus [LL62b]. Moments [Rob65a]. Monotone [BH64, Pri68, Thu69b]. Monte [Hab66, Hab67b]. Monte-Carlo [Hab66, Hab67b]. Motion [FL61, PWMD63, WP60]. Moulton [BRB65]. Muller [Cla62]. Multidimensional [DG63]. Multiple [Bar68c, Bar68d, HO69, Hir68, Sal62a, SS63a]. Multiplication [JJ65, LF68]. Multiplicative [Bar62, Col69]. Multipoint [Jar66]. Multistep [Bar68b, GM68c, New63]. Multivariate [Mon64].

N [Sho60, Uch66]. N. [Pin69]. Natural

[BB68, CH67]. Navier [Cho68b, Cho69]. Needed [Mac65, MM67]. Negative [LB61]. Neumann [Dor69, Wig69]. Newton [Bro67, Lon60b, SF69]. NF [Hou61]. Nine [Kil60, Lee69, PK64]. No [Wag67]. Nodes [SS63a]. Non [Bur63b, Dal63, KF62, LS66, SL61, Wun63, Hou61].

Non-Differentiable [SL61]. Non-Existence [Bur63b, Wun63]. Non-Linear [Dal63]. Non-Polynomial [LS66]. Non-Stationary [KF62]. Noncommutative [Wid66]. Nonexistence [Kil60]. Nonlinear [Bro65, Coo67, Day67, G.68, GM68a, GM68b, GM68c, Kel69, Lin69, Nas65]. Nonnegative [Dav67, Wal69, Wil69]. Nonselfadjoint [Car69]. Nonsingularity [Hof65]. Nonsymmetric [Yam67]. Nordsieck [LS67c, LS67b]. Norm [BW67a, BW67b]. Normal [Amo69, Geb64]. Normally [Geb64]. Note [Ayo67, Bar66, BI66, Bur62a, Cas69, Cav63, Cla62, Day66, DB60, DE62, Des63, Fai69, Fet60a, Fet60b, Flo60, Glo60, Gor68, Hab64, Kil60, Kro67, LRB69, Lee62, Mil68b, Mul59, New67b, Osb62, PK64, Par64a, Rea61, Rie64, Rud66, Sal60b, Sal62b, SM63, Sha60a, Uch66, Ver66c, Ver67a, Ver67b, Yan65, de 67]. Notes [AØS67, And65, Atk68, Bar68a, BDN69, Bar66, BS67a, BH62a, Bau60, Bau64, BK62, BKK64, BI65, BI66, Ber66, Ber67, Bes64, Bes68, BK64, Boe60, Boe61, Boe62, Bow67, BF67b, BM68, BJ60, Bri63, Bro64, Bur62a, Bur62b, Bur63a, Bur63c, Bur64, Bur67, Bur66, Bus68, Bus69b, But61, Car62, Car65, Cav63, Cha68a, Cha67, CJ68a, Cha68b, CG68, Chi69, Chu65, Coc69a, Cod68, Col69, Day63, Day64, Day66, DGL66, DB60, DE62, Dor69, DKV63, DM62, Ebe64, Ede66, Eps62, Erd64, FL66, Fai69, Fet60b, Fet64, Fet67, Fis64, For64, Fra69, Fra65a, Fre61, Fro62, GB68, Gar69, Gar65b, Gar67, Geb64, Geb67, Ger64, Gla66, Gla68, Glo60, Glo62, Glo65, Goo65, Gor68, Gre60, GV68, Hab64, Hab67a, HO69, Har62a, Har66, HKW60, Hea65]. Notes [How62, HJ64, Hum64, Hun64, Hun68, Hur62, Ise61, Ise65, JJ65, Jar66, Jen62, Joh63, Joh65, Joh68, JLB67, Kar61, KK64, Kee65, KS65, Ker69, Klo64, Kon65, Kra60, Kra61, KB64, Kra66, Kro67, La 63, La 64, LJB66, LB66, Lal67, LD68, Lal69, LRB69, LS65a, LP66, LP67b, LP67a, LP67c, Lan60, Lan67, LL62a, Lav64b, Lav64a, LF68, Lay64, Lay67, LS67a, Lea67, Led61, LS65b, Les65, LL62b, Lin65a, Lin65b, LW65, LLS65, Lon60b, Lor60, Lot62, Low66, Low60a, LFiCM65, Lun68, Lut68, Lyn65a, Mac65, Mag60, Man61, Mar68, May65, McA63, Mec68, Mel69, Mil63, Mon64, Moo67, MM66, Mus66, New63, New66, New67a, New69, New67b, Nik60, Old68, Olu65, OD65, OV61, PS60, PK64, Par64a, Par66, Par67, Pat68a, Pat69, PG60, Pin69, Por63, PRS63]. Notes [Pri65, Rab68, Rad60, Raz65, Ric66, Rie62, Rie63, Rie64, Rie69b, Rig64, Rig65, Rob65a, RS61, RR68, Rot60, Rud66, Sal60b, Sal62a, Sal62b, SL68, SO60, Sch63, Sch67, SG68, SH64, Sen64, Sha65a, Sha60a, Sha64a, Sha64b, SK67, Sha67b, Sha62, Sha63b, Sie65, SS68, SH65, Sta62, Ste64b, Ste68a, Ste68b, Str65, Str63, Tak66, Tam62, Tho65, Tho66, TM68, Tuc67, Tun66, Uch66, Var66, Ver65, Ver66a, Ver67a, Ver67b, Voi66, vZW64, vZW65, WG60, WG64, Wat62b, WMR67, Wel61, WM65, WC61, Wil68, WGZ65, WZ68, Wil67, Wit64, WKG66, Woo66, Woo67, Woo68, Wra64, Wre60, WS66, Wun62, Wun63, Yan65, Yan66, Yan69, de 67, vW69]. Notices [Bur67, JLB67, LP67b, Moo67, SK67, Ver67a]. Null [Lyn65e]. Number [Ayo67, BK62, CEFT62, Ehr67, Erd64, FG62, Lal69, Lyn65b, Map63, Met69, Mit66, Pax61, Rie63, SO60, Sha60b, Sha69, WM65, vW69]. Number-Theoretic [Lal69]. Numbers [AØS67, BH62a, BF67a, BM68, BJ60, Bri62, Bri64, But61, CJ68a, Coh62, Ehr67, Geb64, Geb67, Gre61, Hag69, Jag65, Kar61, KB67, Kra61, KB64, Lee68, Mus66, RR69, Rie62, Rie69b, SL68, Sch67, SH64, Sha61, Sha67a, Tau65, Wag67, WZ68, Wra64, Wun62, Sha65c]. Numerical [Bar68a, Bar68c, Bar68d, Bat63, BK62, BHK68, But66, Cho68b, Coh62, Coh65, Coh67b, Con67, Day63, Day64, DGL66, DL67, Dic68, Fet64, Feu68, Fra65a, Gau69, Gea67, Goo65, Ham64, Ham65, Has61, Hou63b, Jam68b, Jon61, Kam69, Kat68, KM68, KM67, LJB66, LS65a, LS66, Led61, LS67c, Lon60a, Low60a, Low60c, Low60d, LD67, LN67, McL63, Mel68, Mil60b, Mil60a, Mil60c, ML67, Mur67, Mus64, Nor62, PD61, Pop60, RBM67, SS64, SS69, SO60, Sim65, SL69, Str60a, Str61b, Tre66, Val68, Voi66, Wit64, Wyn61, vW69, vv66, von67, LS67b, G.68]. Numérique [Hou61, Hou63a]. numerische [Ric65].

Occur [Van64]. Occurring [Ker69]. Octic

[Cav63]. Odd [Hag69, Mus66]. Off

[Des63, Low60c, Low60d, Lyn64, Wig69]. Offset [Cad64, DJ61]. One [Asc60, Day64, Fra64, Ger64, Gin63, Hub66, KM68, Let66, Sal62a, SW69a, Sha65c]. One-Dimensional [Asc60, Gin63, Hub66]. One-Evaluation [KM68]. One-Step [Day64, Let66, SW69a]. Operational [Wyn61]. Operations [Fie60]. Operator [Del65, GM67, Mil68a, Nas65, Pet63, Pet65, Sim67, SD65]. Operators [Car69, Jam68a, Wen65]. Optimal [KS65, McL63, SF69]. Optimized [GM68c]. Optimum [Fre61, HK68, HJ64, Pat68b, Sal62c]. Optimum-Interval [Fre61]. Optimum-Point [Sal62c]. Order [Bat63, BF67a, Blu62, Bra63, But65, Day64, DG63, FL66, FG67, FC67, Jar66, Kil60, LS67a, LD62, Lut68, Mec68, Osb62, PK64, Rab60, Rab68, RS61, Sha64c, Sha65b, Ste69a, Tre66, Yan66, Yan68, Yan69]. Orders [CT67b, GV68, NDT69, SS68, CT67a]. Ordinary [But66, Coo68, Day64, Fra65b, Gar65a, Gea67, KM68, ML67, Mur67, Nor62, Osb67]. Ordinate [Cas65]. Ordinates [SH65]. Orthogonal [Hir68, Hol69, Smi65]. Orthogonalization [Ric66]. Oscillating [Kam61, vv66]. Oscillation [Pri68]. Oscillatory [Lon60a]. Osculatory [Kah69, Sal60a, Sal62b, Sal62c]. Other [LC61, Ric64a]. Over-Relaxation [Bro64, Dor69, Lyn64, Rig65]. Overrelaxation [Gar67, HW61, HK68].

p [Ham64, Hou60, Hou61, Hou63a, Hou63b]. Packing [Mel69]. Padé [Ell67, Fai64, SM63]. Pair [Dal63]. Pairs [BLL64, Lee69]. Paper [Sch63]. Papers [Str61a, AØS67, And65, Atk68, Bar68a, BDN69, Bar66, BS67a, BH62a, Bau60, Bau64, BK62, BKK64, BI65, BI66, Ber66, Ber67, Bes64, Bes68, BK64, Boe60, Boe61, Boe62, Bow67, BF67b, BM68, BJ60, Bri63, Bro64, Bur62a, Bur62b, Bur63a, Bur63c, Bur64, Bur66, Bus68, Bus69b, But61, Car62, Car65, Cav63, Cha68a, Cha67, CJ68a, Cha68b, CG68, Chi69, Chu65, Coc69a, Cod68, Col69, Day63, Day64, Day66, DGL66, DB60, DE62, Dor69, DKV63, DM62, Ebe64, Ede66, Eps62, Erd64, FL66, Fai69, Fet60b, Fet64, Fet67, Fis64, For64, Fra69, Fra65a, Fre61, Fro62, GB68, Gar69, Gar65b, Gar67, Geb64, Geb67, Ger64, Gla66, Gla68, Glo60, Glo62, Glo65, Goo65, Gor68, Gre60, GV68, Hab64, Hab67a, HO69, Har62a, Har66, HKW60, Hea65]. Papers [How62, HJ64, Hum64, Hun64, Hun68, Hur62, Ise61, Ise65, JJ65, Jar66, Jen62, Joh63, Joh65, Joh68, Kar61, KK64, Kee65, KS65, Ker69, Klo64, Kon65, Kra60, Kra61, KB64, Kra66, Kro67, La 63, La 64, LJB66, LB66, Lal67, LD68, Lal69, LRB69, LS65a, LP66, LP67a, LP67c, Lan60, Lan67, LL62a, Lav64b, Lav64a, LF68, Lay64, Lay67, LS67a, Lea67, Led61, LS65b, Les65, LL62b, Lin65a, Lin65b, LW65, LLS65, Lon60b, Lor60, Lot62, Low66, Low60a, LFiCM65, Lun68, Lut68, Lyn65a, Mac65, Mag60, Man61, Mar68, May65, McA63, Mec68, Mel69, Mil63, Mon64, MM66, Mus66, New63, New66, New67a, New69, New67b, Nik60, Old68, Olu65, OD65, OV61, PS60, PK64, Par64a, Par66, Par67, Pat68a, Pat69, PG60, Por63, PRS63, Pri65, Rab68, Rad60]. Papers [Raz65, Ric66, Rie62, Rie63, Rie64, Rie69b, Rig64, Rig65, Rob65a, RS61, RR68, Rot60, Rud66, Sal60b, Sal62a, Sal62b, SL68, SO60, Sch63, Sch67, SG68, SH64, Sen64, Sha65a, Sha60a, Sha64a, Sha64b, Sha67b, Sha62, Sha63b, Sie65, SS68, SH65, Sta62, Ste64b, Ste68a, Ste68b, Str65, Str63, Tak66, Tam62, Tho65, Tho66, TM68, Tuc67, Tun66, Uch66, Var66, Ver65, Ver66a, Ver67b, Voi66, vZW64, vZW65, WG60, WG64, Wat62b, WMR67, Wel61, WM65, WC61, Wil68, WGZ65, WZ68, Wil67, Wit64, WKG66, Woo66, Woo67, Woo68, Wra64, Wre60, WS66, Wun62, Wun63, Yan65, Yan66, Yan69, de 67, vW69]. Parabolic [Bat63, CK68, FG67, Hub66, Jam68b, Lee66, Ros60, Wid65]. Parallel [Kam69, ML67, Sha62]. Parallelogram [Lin60, LT64]. Parameter [Ebe64, PR63, Wim67]. Parameters [DJ67, HK68, KS65, Sta62]. Paris [Hou61, Hou63a]. Parity [PS67]. Parlett [For64, Var66]. Part [Mil60b, Mil60c, SS66]. Partial [Dal63, FG67, SK61]. particular [Sha65c]. Partition [But61, McA63, PS67]. Partitioned [Gol63, LNP69]. Partitions [Che64, New67b]. Patterns [BHK68]. Paul [Sha65c]. Perfect [Mus66, SO60]. Perfectly [RR69]. Period [Lin60, LT64]. Period-Parallelogram [Lin60, LT64]. Periods [LD62, Lin60, LT64, Lin65a]. Permanents [Nik60]. Permutation [BB68, Cav63]. Permutations [Har62a, How62, Joh63, Wel61, de 67]. Person [APR65]. Peter [H.65, Ham65]. Phase [Sol67]. Physical [MM67]. Piatetsky [HO69]. Piatetsky-Shapiro [HO69]. Places [Knu62]. Plane [Dav64, GH67]. Planes [Kee65, Kil60, PK64]. Point [Jag66, LS67c, Pop60, Sal62c, LS67b]. Points [BS67a, BC69, Cla62, FG62, Mil60b, Mit66, Mul59, Pat68b, RR69]. Poisson [Bra63, Bra66, Jam68a]. Polya [Wil63]. polygon [Coc69b]. Polygons [Gol60]. Polyhedra [Gol60, Gra63]. Polylogarithms [Sha64b]. Polynomial [BKK63, Bur62b, Bur67, Bur66, Car65, Cha68a, Coc69a, Cod65b, Den66, Kam61, Klo64, LS66, Lan67, Lea67, LW63, Mon64, NDT69, Shi66, Smi65, Spi62, Sta62, Ste69b, Ste71, Tra66, Wim61, Wim62, Shi60, Wim72]. Polynomials [AM61, Bar61, Bes64, Cav63, GW60, Hir68, Hol69, Hum64, Lor60, McE69, New66, SCF64, Sta63, Swi60, Tho65, Wat62b, Sha65c, Wat62a]. Polynomic [McA65]. Posed [Mil68a]. Position [Har62a]. Positive [BHT69, Dan69b, Kat68, LSL64, Low66, Mar68, TM68, vZW65, Wal69]. Possible [Lot62]. Post [LF68]. Potential [Bur68, von67]. Power [BLL64, Fra64, KK64, Lea66, Man67, Rei69, Wyn60]. Powers [Bar66, Bes68, Bru69, GV68, LP67b, LPS67, LL62b]. pp [Ric65, Sha65c]. pp. [G.68]. Practical [Bur67, Pri65]. Pre [LF68]. Precision [PG60]. Predictor [Cas69, KM68]. Predictor-Corrector [Cas69, KM68]. Predictors [Ste65]. Preserving [Bar68b]. Press [Ham64]. Price [Hou60, Hou61, Hou63b, Ric65]. Primality [Rie69a]. Prime [BH62a, Bri62, Ehr67, Hag69, LP67c, WZ68, Wre61]. Primes [Gil64, Glo65, Hur62, Joh65, JLB67, Lal67, LP67a, Map63, Sha60b, Sha60a, Sha64a, Sha65c]. Primitive [Lin60, LT64, RR68, Wat62a, Wat62b]. Principle [Gor68]. Principles [Del65]. Priori [Gre61]. Probabilities [Whi63]. Probability [Gre60]. Problem [AH64, Bat63, Bra63, Car69, Day63, Dor69, Dra61, Eps62, Erd64, Gar65b, Gin63, Kam69, KS63, Lun68, McA63, Mei69, Mil68a, MM66, PS60, Par64b, Por63, Sim65, Sol66, Sol67, Ste64a, Thu69b, Wig69, WGZ65, YP61, ZL69].

Problems [BH64, BHT69, BH62b, God66, Goo65, HL69, Kre68, Low60a, Low61, MM67, Mur67, Osb62, Osh69b, Pop60, Sha69, Thu69a, Wyn62]. Procedure [BC67, For64, Mar68, Var66]. Procedures [Jor68, Man67, New67a]. Process [Whi63]. Processes [BKK63, But64b, But64a, Des63, Fra63]. Product [Bau64, GS62, LL62a, Str61b, vH63]. Products [GW60, KKI67, Lor60, Mil63, OD65, OB64]. Programming [BKK63, Ste64a]. Programs [Gea69]. Progression [LP67a, Wag67]. Projection [Lyn65d]. Projective [Kee65, Kil60, PK64]. Proof [LLMS62, SL68]. Propagation [H.65, Low60b, Low60c, Low60d]. Properties [BRB65, Bus69b, Dal63, Mil68b, S.64, Wun62]. Pseudo [Bar62, But61, Geb67, Jag65]. Pseudo-Random [Bar62, But61, Geb67, Jag65]. Pseudorandom [GC68]. Psi [Tho65]. Publishers [Hou60]. Pyramidal [Wun62].

QR [Par68]. Quadratic [Ayo67, Coh62].

Quadrature

[And65, BDN69, Bur63b, Bur63c, But64b, Cas65, CJ68b, Cha68b, Cha68c, CJ68c, CCJM63, Coo68, Day63, Dur66, Fos68, Gal69, Gau68, GW69, Hab64, Hab66, Hab67b, Hab67a, Hab69, HW60, Har62b, Jag66, LLS65, LN67, Lyn69, Mic63, Mil60b, Mil60a, Mil60c, New69, Pat68b, Pin69, Rab60, Rei67, RS61, Rom64, Rot61, Sal62a, Sar68, SS69, Sha65a, SL69, SS63a, Ste68a, Str63, Str60b, Whi65, Wil69]. Quadratures [BW67a, BW67b, Dav67, SBG69]. Quartic [Sal60b]. Quasi [Bro67, Nav63]. Quasi-Newton [Bro67]. Quasi-Step [Nav63]. Quasilinear [Lee66]. Questions [WS66]. Quotients [Bra66].

R [Sho60]. Radau [But64b]. Radial [OD65]. Radius [BS67a, DP65]. Radix [MA65, Sch67]. Ramanujan [Woo66]. Random [Bar62, But61, Fra63, Geb64, Geb67, Gre61, Jag65, KF62, Tau65, de 67]. Randomness [Pat62]. Rank [Ede66, Gol63]. Rapid [JJ65]. Rate [Wid66]. Rates [Gar67, LR68]. Ratio [DJ67, OM68]. Rational [BF66, CT67b, CT68, Cod69, FL66, FL67, Fie65, Flo60, GB68, KR68, LS65a, Sal62b, Wyn60, CT67a]. Rationals [Ric64a]. Ray [LW63]. Rayleigh [Wen65]. Reactors [KG68]. Real [Coh62, DM69, Hof65, La 63, La 64, WG64, WB64]. Recherches [Hou63a]. Records [Ber67]. Rectangle [Dor69]. Rectangular [Fre61, Lin60, Mil60b, Mil60a, Mil60c]. Recurrence [Mec68, Olv64, Tai67, Tau65]. Recursion [Wim68]. Recursive [Ber67, Gau61]. Reducing [Bus69a]. Reduction [La 63, LNP69, MM66, NC66, Rig64, WG64]. Region [Hub66, Lea67]. Regions [Str61b, SS63b, Str66]. Regular [WS66]. Related [BS67b, Bur62b, CR68, Fle68, Har66, Nel61, SG68, SCF64, Woo66]. Relating [Boe61, Lin65b, Wil63]. Relations [HP62, LF68, SD69]. Relationship [Hir68]. Relative [Coh62, Gre67, SM63]. Relatively [Hag69]. Relaxation [Bro64, Dor69, Lyn64, Rig65]. Remainder [Sta63]. Remainders [BW67a, BW67b]. Remark [Eps62, Joh65, Sch63]. Remarks [Dor69, Sha63a, Sha62, Sol66, Yam67]. Repeated [Gau61, Wil64]. Report [WZ68]. Representation [LS67a, Raz65, Sch67, Wim67]. Representations [GV68]. Represented [Sha65c]. Required [Lyn65b]. Residues [BLL64, Fra64, LLMS62]. Resolvent [Mil68a]. Restricted [Har62a]. Resulting [LF68]. Results [BS67b, Bru69, SS69, SS65]. Review [G.68, H.65, Ham64, Ham65, Hou60, Hou61, Hou63a, Hou63b, Ric65, S.64, Sha65b, Sha65c, Sho60]. Reviews [Ano60d, Ano61d, Ano62e, Ano63e, Ano63f, Ano63g]. Rhombic [LT64]. Ricatti [Fai64]. Riccati [FL66]. Richardson [Let66]. Riemann [GC68, HP62, Leh66, SD65]. Rigorous [Var68b]. Ritz [Rig64, Wen65]. Roger [Sha65c]. Root [Mar68, Tha61]. Roots [BA62, DE62, JJ65, Lon60b]. Rotenberg [Sho60]. Rotors [Gol60]. Round [Des63, Low60c, Low60d, LD67, Lyn64]. Round-Off [Des63, Low60c, Low60d, Lyn64]. Rule [Lyn65d, Ste68b, Tuc67]. Rules [Gal69, GW69, Lyn65a, Lyn65b, Lyn65c, Lyn65d, Lyn65e, Val68]. Runge [Blu62, But64a, But65, Fyf66, HJ64, Kin66, Lut68, Ral62]. Russian [Hou60].

Satisfying [Cha68a]. Scheme [Lee66, LS67c, Yam67, LS67b]. Schemes [Ape68, Eis68, Gra64, Hub66, Low61, RW66, Ros64, SW69b, Wid65]. Schmidt [Ric66]. Schoenberg [Wil63]. Schrödinger [Coo61, Rad60]. Search [Gra63, Kee65]. Second [Bat63, Bur66, Day64, DB60, FL66, FL67, Lin69, Mec66, Osb62, Sha64c, Ste69a, Van69]. Second-Degree [Bur66]. Second-Order [Bat63, Osb62, Sha64c]. Sections [Spi66, Spi68]. Segments [DE62, Lav64a, Rea61, Sto61]. Self [CK68, LS66]. Self-Adjoint [CK68]. Self-Adjusting [LS66]. Semi [LW63]. Semi-Infinite [LW63]. Senior [Sha65b]. Separation [Leh66]. Sequence [Geb67, GC68]. Sequences [Bar62, Jag64, Moo66, Tra65, Wag67]. Serial [Gre61]. Series [AM61, Atk68, BK64, Chu65, CT65, Dan69b, Den69, Ell64, ES65, Ell65, Fie60, Hol69, Hum64, Lan67, Lea66, Les65, LC61, LFiCM65, Man67, Rud66, SK61, Scr69, SW63, Sha64b, Smi65, Tun66, Van69, Wre68, Wyn60]. Set [BS67a, DM69, Gea69, de 64]. Sets [BF67a, Hay65]. Seventh [BLL64]. Several [LSL64]. Shaped [Lea67]. Shapiro [HO69]. Shell [Mus64]. Ship [TS61]. Short [AØS67, And65, Atk68, Bar68a, BDN69, Bar66, BS67a, BH62a, Bau60, Bau64, BK62, BKK64, BI65, BI66, Ber66, Ber67, Bes64, Bes68, BK64, Boe60, Boe61, Boe62, Bow67, BF67b, BM68, BJ60, Bri63, Bro64, Bur62a, Bur62b, Bur63a, Bur63c, Bur64, Bur67, Bur66, Bus68, Bus69b, But61, Car62, Car65, Cav63, Cha68a, Cha67, CJ68a, Cha68b, CG68, Chi69, Chu65, Coc69a, Cod68, Col69, Day63, Day64, Day66, DGL66, DB60, DE62, Dor69, DKV63, DM62, Ebe64, Ede66, Eps62, Erd64, FL66, Fai69, Fet60b, Fet64, Fet67, Fis64, For64, Fra69, Fra65a, Fre61, Fro62, GB68, Gar69, Gar65b, Gar67, Geb64, Geb67, Ger64, Gla66, Gla68, Glo60, Glo62, Glo65, Goo65, Gor68, Gre60, GV68, Hab64, Hab67a, HO69, Har62a, Har66, HKW60, Hea65]. Short [How62, HJ64, Hum64, Hun64, Hun68, Hur62, Ise61, Ise65, JJ65, Jar66, Jen62, Joh63, Joh65, Joh68, JLB67, Kar61, KK64, Kee65, KS65, Ker69, Klo64, Kon65, Kra60, Kra61, KB64, Kra66, Kro67, La 63, La 64, LJB66, LB66, Lal67, LD68, Lal69, LRB69, LS65a, LP66, LP67b, LP67a, LP67c, Lan60, Lan67, LL62a, Lav64b, Lav64a, LF68, Lay64, Lay67, LS67a, Lea67, Led61, LS65b, Les65, LL62b, Lin65a, Lin65b, LW65, LLS65, Lon60b, Lor60, Lot62, Low66, Low60a, LFiCM65, Lun68, Lut68, Lyn65a, Mac65, Mag60, Man61, Mar68, May65, McA63, Mec68, Mel69, Mil63, Mon64, Moo67, MM66, Mus66, New63, New66, New67a, New69, New67b, Nik60, Old68, Olu65, OD65, OV61, PS60, PK64, Par64a, Par66, Par67, Pat68a, Pat69, PG60, Por63, PRS63]. Short [Pri65, Rab68, Rad60, Raz65, Ric66, Rie62, Rie63, Rie64, Rie69b, Rig64, Rig65, Rob65a, RS61, RR68, Rot60, Rud66, Sal60b, Sal62a, Sal62b, SL68, SO60, Sch63, Sch67, SG68, SH64, Sen64, Sha65a, Sha60a, Sha64a, Sha64b, SK67, Sha67b, Sha62, Sha63b, Sie65, SS68, SH65, Sta62, Ste64b, Ste68a, Ste68b, Str65, Str63, Tak66, Tam62, Tho65, Tho66, TM68, Tuc67, Tun66, Uch66, Var66, Ver65, Ver66a, Ver67a, Ver67b, Voi66, vZW64, vZW65, WG60, WG64, Wat62b, WMR67, Wel61, WM65, WC61, Wil68, WGZ65, WZ68, Wil67, Wit64, WKG66, Woo66, Woo67, Woo68, Wra64, Wre60, WS66, Wun62, Wun63, Yan65, Yan66, Yan69, de 67, vW69]. Shuffling [Geb67]. Sieve [CEFT62]. Sign [BH62b]. Significant [Ger64]. Similar [Tho66]. Similarity [La 63]. Simple [Gea69, LB61, Rei67, SS68, SS63a, Tuc67]. Simplex [Dur66]. Simplexes [Str64]. Simplified [TM68]. Simulation [Fra63]. Simultaneous [Bro65, DM69, Pet65]. Single [Coo67]. Single-Step [Coo67]. Singular [BK68, Eis68, Jam68b, KS65, Par66, Ric69, Sha63b]. Singularities [Voi66]. Singularity [Wig69]. Six [MTK62]. Sixteen [Mil60b]. Sixth [Lut68]. Sixth-Order [Lut68]. Size [Let66]. Slowly [Den69]. Smooth [LR68]. Smoothing [PD64]. Solid [Mel69]. Solid-Packing [Mel69]. Solution [Bat63, BK62, BKK64, Cho68b, Con67, Day63, Day64, DGL66, Dic68, Dor69, Eps62, Fai64, FL66, Fet64, God66, Goo65, Jon61, Kam69, Kat68, KM68, LS65a, LS66, Mag60, Mit61, PM62, PD61, Pet63, Pet65, Pop60, RS61, Sal60b, Scr69, Sim65, Sim67, Ste69a, Tra66, Voi66, WGZ65, vW69, von67]. Solutions [Asc60, BF66, Bra66, But66, G.68, GLS64, LJB66, LB66, LD68, Nas65, Pea69, PRS63, Ros60, Sha66, Thu69a, Wal69, WM65]. Solving [Bro65, Coo61, Day67, Fro62, HW61, Han68, HL69, Jar66, Lin69, McA65, MP63, Osb62, Sch62, Sta62]. Some [Bri63, BS67b, Bru69, Coh69, ES65, FG67, GB68, Gla66, Gla68, GS69, Hab64, HP62, Hub66, Ise65, Jag65, Jar66, Kan66, Lar64, Lev61, LNP69, Man67, New69, Pat68a, Rie69b, Rob65b, Sha62, Sol66, Str66, Van64, Wen65, Wim62, Woo66, Yam67, Yan66, de 64, Jor68, Wim72]. Sons [G.68]. Sorting [BB68]. Space [Bro69, DKV63, GM67, GM68c, Lyn65a]. Spaced [Mil60b]. Spaces [BB68]. Span [Rob64]. Sparse [DM62]. Special [Bes68, Cha68a, Fos68, Kee65, Olu65]. Spectra [KF62, WG60]. Spectral [DP65]. Speed [CEFT62, Lev61, Mit61, TS61]. Sphere [BF67b, Bur66, Feu68, FG62, Gol69, McL63]. Spheres [Mul59, Cla62]. Spherical [AM61, Mus64]. Spherically [Lyn65a, SS63b]. Split [GM67]. Springer [Ric65]. Springer-Verlag [Ric65]. Square [Gol63, GTM69, JJ65, La 63, Mil60b, NF63, Tha61, WG64]. Squares [Coh67a, DSS68, HL69, Klo64, OV61, Sie65, Wun63, Jor68]. Stability [BRB65, CK68, Cas69, Dal63, FL61, Gol69, KM68, Kre68, Low61, Osh69b, Sat66, Wid65]. Stabilizing [Ste65]. Stable [FG67, Kro67, Lar64]. Stanislaus [Bau60]. Starting [Day67, New67a, SF69]. State [Sol67]. Stationary [KF62]. Statistical [Gil64, Pat62]. Statistics [JLB67]. Steady [Sol67]. Steen [Pin69]. Steepest [Mei69]. Stefan [Sol66]. Steiner [Coc69b]. Step [Coo67, Day64, Let66, Nav63, SW69a]. Stepwise [Lot62]. Stiefel [Ham64]. Stochastic [Fra65b, Hab69]. Stokes [Cho68b, Cho69]. Strategy [Ste64a]. Strongly [vv66]. Studies [Dra61]. Study [Coh62, Coh67b, Den69, Pat62, SS68, WG60]. Sturm [BK68, Car69]. Submerged [WP60]. Substitution [Bus68, New67a]. Subtract [Wig69]. Successive [Bro64, Dor69, Gau66, HW61, Joh65, Lyn64, New67a, Rig65, Sha64a]. Such [Jon61]. Sufficiency [Dur66]. Sum [Coh67a, LP67b, Les65, SL68]. Summation [Dan69b, Den69, SK61, Ver67b]. Summing [Smi65]. Sums [Bar66, Bru69, LRB69, LP66, LPS67, Lan67, Led61, Lee69]. Supplementary [Sha63a]. Surface [Con67, OB64]. Survey [Coh65, Hou63b, LPS67]. Surveys [Str61a]. Suspension [Dic68]. Sylvester [Bar68b]. Symbols [Sho60]. Symmetric [Cha67, Com60, Gol62, HW61, Kat68, La 64, LSL64, Lyn65a, Lyn65c, Lyn65d, Lyn65e, RR69, SS63b, Str66, Str68b, WB64]. Symmetry [Lyn65a, Lyn65b]. Symposium [Gau94]. System [Cha68a, MA65]. Systematic [AØS67]. Systems [AZ69, BK68, Coo67, GM67, GM68a, GM68b, GM68c, Gra64, Han68, Jon61, Kam69, Osh69a, Sch62, Yam67].

T [Sha65c]. Table [Ano60e, Ano61e, Ano63h, Ano63i, Ano63j, BA62, Glo62, Hea65, SL61, TM68, WD61]. Tables [Ano60d, Ano61d, Ano62e, Ano63e, Ano63f, Ano63g, Bau64, Bes61, CCJM63, Kir60, Lin65b, LW65, Nel61, Pin69, SCF64, Str60b, Sha65c]. Tabulation [EG63, Fie60]. Tangent [KB67]. Taylor [Fie60]. Tchebycheff [Tho65]. Technical [Bur67, JLB67, LP67b, Moo67, SK67, Ver67a, AØS67, And65, Atk68, Bar68a, BDN69, Bar66, BS67a, BH62a, Bau60, Bau64, BK62, BKK64, BI65, BI66, Ber66, Ber67, Bes64, Bes68, BK64, Boe60, Boe61, Boe62, Bow67, BF67b, BM68, BJ60, Bri63, Bro64, Bur62a, Bur62b, Bur63a, Bur63c, Bur64, Bur66, Bus68, Bus69b, But61, Car62, Car65, Cav63, Cha68a, Cha67, CJ68a, Cha68b, CG68, Chi69, Chu65, Coc69a, Cod68, Col69, Day63, Day64, Day66, DGL66, DB60, DE62, Dor69, DKV63, DM62, Ebe64, Ede66, Eps62, Erd64, FL66, Fai69, Fet60b, Fet64, Fet67, Fis64, For64, Fra69, Fra65a, Fre61, Fro62, GB68, Gar69, Gar65b, Gar67, Geb64, Geb67, Ger64, Gla66, Gla68, Glo60, Glo62, Glo65, Goo65, Gor68, Gre60, GV68, Hab64, Hab67a]. Technical [HO69, Har62a, Har66, HKW60, Hea65, How62, HJ64, Hum64, Hun64, Hun68, Hur62, Ise61, Ise65, JJ65, Jar66, Jen62, Joh63, Joh65, Joh68, Kar61, KK64, Kee65, KS65, Ker69, Klo64, Kon65, Kra60, Kra61, KB64, Kra66, Kro67, La 63, La 64, LJB66, LB66, Lal67, LD68, Lal69, LRB69, LS65a, LP66, LP67a, LP67c, Lan60, Lan67, LL62a, Lav64b, Lav64a, LF68, Lay64, Lay67, LS67a, Lea67, Led61, LS65b, Les65, LL62b, Lin65a, Lin65b, LW65, LLS65, Lon60b, Lor60, Lot62, Low66, Low60a, LFiCM65, Lun68, Lut68, Lyn65a, Mac65, Mag60, Man61, Mar68, May65, McA63, Mec68, Mel69, Mil63, Mon64, MM66, Mus66, New63, New66, New67a, New69, New67b, Nik60, Old68, Olu65, OD65, OV61, PS60, PK64, Par64a, Par66, Par67, Pat68a, Pat69, PG60, Por63]. Technical [PRS63, Pri65, Rab68, Rad60, Raz65, Ric66, Rie62, Rie63, Rie64, Rie69b, Rig64, Rig65, Rob65a, RS61, RR68, Rot60, Rud66, Sal60b, Sal62a, Sal62b, SL68, SO60, Sch63, Sch67, SG68, SH64, Sen64, Sha65a, Sha60a, Sha64a, Sha64b, Sha67b, Sha62, Sha63b, Sie65, SS68, SH65, Sta62, Ste64b, Ste68a, Ste68b, Str65, Str63, Tak66, Tam62, Tho65, Tho66, TM68, Tuc67, Tun66, Uch66, Var66, Ver65, Ver66a, Ver67b, Voi66, vZW64, vZW65, WG60, WG64, Wat62b, WMR67, Wel61, WM65, WC61, Wil68, WGZ65, WZ68, Wil67, Wit64, WKG66, Woo66, Woo67, Woo68, Wra64, Wre60, WS66, Wun62, Wun63, Yan65, Yan66, Yan69, de 67, vW69]. Technique [BKK63, May65, Olu65, Sim65]. Techniques [BF66, Cas69, Kan66, Mec68, Wyn62]. Term [Cas65, Sha64c, Sta63]. Terms [Bur63c, Dan69b, Tho65]. Test [But61, Ebe64, Gea69, Her69, KB64, Mil68b]. Tests [For64]. th [GV68, Rei69]. their [Bro67, Sch67, Smi65]. Theorem [Bes61, Chi69, HO69, LLMS62, OV61, SS66, Ste64b, Uch66]. Theorems [Jag65, Jam68b]. Theoretic [Lal69]. Theory [BK62, BK64, Chu65, Erd64, Gil64, Kre68, LFiCM65, Mac65, S.64, Sar68, Sha65c, SL69, vW69]. Third [Boe61, Fet65, Fra65a, LS67a, Hou60]. Thirteenth [Pax61]. Three [Bes64, Bra63, Gil64, Lee66, LW65]. Three-Level [Lee66]. Time [Tre66]. Todd [Hou63b]. Tome [Hou61]. Topological [Coh67b, Coh69]. Toroidal [Rot60]. Toronto [Hea65]. Transcendental [GH67, Rad60]. Transfer [Dra61]. Transform [Ber66, Gau69, GA68, SL69, Ver65, WG60, Ber67, Ber68]. Transformation [Fet65, GS69, Par66]. Transformations [Bus68, Dan69b, La 63]. Transforms [Wim61]. Transient [PWMD63, WP60]. Translated [Hou60]. Transonic [BHK68]. Transport [PM62]. Transpose [LF68]. Transposition [Joh63, PS60, Wel61]. Trapezoidal [Tuc67].

Treatment [Low60a, Low60d]. trees [Coc69b]. Tri [Bus69b, La 63]. Tri-Diagonal [Bus69b, La 63]. Triangle [Bow67, Dav64]. Triangular [Cho68a, Sen64]. Triangulation [Kel64]. Triangulations [BF67b]. Tridiagonal [Gar69, Gol62, Low60b, WG64]. Trigonometric [BK64, Bur64, Chu65, Lew69, LFiCM65, New66]. Trinomial [WM65]. Trinomials [RR68]. Triple [GS62]. Triply [BC67]. Triply-Connected [BC67]. Truncation [AH64, Ell65, Ell67, Osb67]. Tubular [KG68]. Tuned [WG60]. Turning [Ste69a]. Twin [Sha60a, Wre61]. Twin-Prime [Wre61]. Two [Bes61, Boe61, Boe62, Coh67a, DG63, Ebe64, Ell65, GM68c, Ker69, Man66, Man67, MM67, Mil60b, Mil60a, Mil60c, Pop60, PR63, RR69, Ste64a, Str60a, Swi60, Tau65, Ver66b, Ver66c, Wre68]. Two-Dimensional [Man66, RR69]. Two-Parameter [PR63]. Two-Point [Pop60]. Type [BDN69, BH65, Bes68, BH64, BHT69, Bra66, BRB65, Cha68b, Fet64, Kee65, McN64, Osh69b, Rig64, Thu69b, Whi65].

Uhlenbeck [S.64]. Uncommon [Low61]. Unicity [Joh68]. Uniform [Ste69a]. Uniformly [Cla62, Mul59]. Unit [Bur66]. Unnormalized [MA65]. Unstable [Ste65]. Use [Fra69, Let66, Str60b, Val68]. Using [Ber68, Cas65, Han68, La 63, Lay67, Led61, Lyn65e, Mil60b, Mit61, PR63, RS61, Sha65a]. Utility [Lon60b].

V [Hou60, Sha65c]. Value [Day63, Kre68, Osh69b, Pop60, Sim65, Thu69a, Thu69b, Yan65]. Values [Ape68, DJ67, GB68, HP62, HKW60, Kir60, KL68, Lin65b, LW65, Lyn69, Man67]. Vancouver [Gau94]. Variable [Let66, Str65, Sha65c]. Variables [Boe62, DKV63, GM67, GM68c, MTK62, Ver66b, Ver66c]. Variate [Fre61]. Variation [Del65]. Variational [RBM67]. Variations [SS66]. Varieties [Ste69a]. Various [Coh65, GV68]. Vector [Che64, Wyn62]. Vectors [DE62]. Verlag [Ric65]. Version [LS67c, LS67b]. vertices [Coc69b]. Very [CEFT62]. viii [Ric65]. Villars [Hou63a]. Virial [S.64]. Viscous [FL61]. Voigt [Rei67, Rei69]. Volterra [Day67, Lin69].

Volumes [LWF72].

W [S.64]. Wall [Asc60]. Walsh [Ric64b]. Waring [Ste64b]. Wave [Boe69, BH62b, Low60d, OD65, Pea69]. Waves [Mac65]. Weakly [Ste65]. Weierstrass [Lin60, LT64, Lin65a, Lin65b, SL61]. Weight [Mic63, Rot61]. Weights [Rab60, SCF64, SH65, Whi65]. Well [Mil68a]. Well-Posed [Mil68a]. Wendroff [GM68c, Yam67]. Which [Van64, Wyn60]. whose [Jam68b]. Wiener [Fos65, KM67]. Wiley [G.68]. Williamson [BH65]. without [Ste69a]. Wooten [Sho60]. Wrong [KL68, Osh69a].

x [G.68, Ham64]. xv [Hou60]. xvi [Hou63b]. xxvii [Hou61].

York [G.68, Ham64, Hou60].

Zero [Cas65, Low66]. Zeros [Bau64, DL67, Dor66, Ger64, KK64, LL62a, Leh66, SCF64, Spi66, Spi67, Spi68, Ste69b, Ste71, WMR67, Wil68]. Zeta [HP62, Leh66, Spi66, Spi68]. Zeta-Function [Leh66].

References

Amos:1969:CBD

[AB69] D. E. Amos and W. G. Bulgren. On the computation of a bivariate tdistribution. Mathematics of Computation, 23(106):319–333, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aziz:1964:BTE

[AH64] A. K. Aziz and B. E. Hubbard. Bounds on the truncation error by finite differences for the Goursat problem. Mathematics of Computation, 18(85):19–35, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ahlin:1964:BGH

[Ahl64] A. C. Ahlin. A bivariate generalization of Hermite’s interpolation formula. Mathematics of Computation, 18(86): 264–273, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Arthurs:1961:ESB

[AM61] A. M. Arthurs and R. McCarroll. Expansion of spherical Bessel functions in a series of Chebyshev polynomials. Mathematics of Computation, 15(74): 159–162, April 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amos:1969:CBN

[Amo69] D. E. Amos. On computation of the bivariate normal distribution. Mathematics of Computation, 23(107):655– 659, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anderson:1965:GQF

[And65] Donald G. Anderson. Gaussian quadrature formulae for ∫_{0}^{1} −ln(x)f(x)dx (in Technical Notes and Short Papers). Mathematics of Computation, 19(91): 477–481, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1960:Ca

[Ano60a] Anonymous. Corrigenda. Mathematics of Computation, 14(69):97, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1960:Cb

[Ano60b] Anonymous. Corrigenda. Mathematics of Computation, 14(70):222, April 1960.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1960:Cc

[Ano60c] Anonymous. Corrigendum. Mathematics of Computation, 14(71):310, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1960:RDT

[Ano60d] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 14(69):75–95, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1960-14-069/S0025-5718-60-99238-
3;   http://www.ams.org/journals/
mcom/1960-14-069/S0025-5718-60-99238-
3/S0025-5718-60-99238-3.pdf.

Anonymous:1960:TE

[Ano60e] Anonymous. Table erratum. Mathematics of Computation, 14(69):96, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/
journals/mcom/1960-14-069/S0025-
5718-60-99237-1;   http://www.ams.
org/journals/mcom/1960-14-069/S0025-
5718-60-99237-1/S0025-5718-60-99237-
1.pdf.

Anonymous:1961:Ca

[Ano61a] Anonymous. Corrigenda. Mathematics of Computation, 15(73):113–114, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1961:Cb

[Ano61b] Anonymous. Corrigenda. Mathematics of Computation, 15(74):225, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1961:Cc

[Ano61c] Anonymous. Corrigenda. Mathematics of Computation, 15(76):436, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1961:RDT

[Ano61d] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 15(75):299–318, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1961-15-075/S0025-5718-61-99215-
8;   http://www.ams.org/journals/
mcom/1961-15-075/S0025-5718-61-99215-
8/S0025-5718-61-99215-8.pdf.

Anonymous:1961:TE

[Ano61e] Anonymous. Table errata. Mathematics of Computation, 15(75):319– 321, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/
journals/mcom/1961-15-075/S0025-
5718-61-99214-6;   http://www.ams.
org/journals/mcom/1961-15-075/S0025-
5718-61-99214-6/S0025-5718-61-99214-
6.pdf.

Anonymous:1962:Ca

[Ano62a] Anonymous. Corrigenda. Mathematics of Computation, 16(77):126, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1962:Cb

[Ano62b] Anonymous. Corrigenda. Mathematics of Computation, 16(78):262–263, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1962:Cc

[Ano62c] Anonymous. Corrigendum. Mathematics of Computation, 16(79):406, July 1962.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1962:Cd

[Ano62d] Anonymous. Corrigendum. Mathematics of Computation, 16(80):513, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1962:RDT

[Ano62e] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 16(78):252–260, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/journals/mcom/
                                              1962-16-078/S0025-5718-62-99201-
                                              3/home.html.

Anonymous:1963:Ca

[Ano63a] Anonymous. Corrigenda. Mathematics of Computation, 17(81):104, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1963:Cb

[Ano63b] Anonymous. Corrigenda. Mathematics of Computation, 17(82):215, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1963:Cc

[Ano63c] Anonymous. Corrigenda. Mathematics of Computation, 17(83):335, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1963:Cd

[Ano63d] Anonymous. Corrigenda. Mathematics of Computation, 17(84):487–488, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1963:RDTa

[Ano63e] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 17(81):88–99, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/1963-17-081/S0025-5718-1963-
1781088-8;   http://www.ams.org/
journals/mcom/1963-17-081/S0025-
5718-1963-1781088-8/S0025-5718-1963-
1781088-8.pdf.

Anonymous:1963:RDTb

[Ano63f] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 17(81):88–99, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1963-17-081/S0025-5718-63-99186-
5;   http://www.ams.org/journals/
mcom/1963-17-081/S0025-5718-63-99186-
5/S0025-5718-63-99186-5.pdf.

Anonymous:1963:RDTc

[Ano63g] Anonymous. Reviews and descriptions of tables and books. Mathematics of Computation, 17(82):202–213, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1963-17-082/S0025-5718-63-99182-
8;   http://www.ams.org/journals/
mcom/1963-17-082/S0025-5718-63-99182-
8/S0025-5718-63-99182-8.pdf.

Anonymous:1963:TEa

[Ano63h] Anonymous. Table errata. Mathematics of Computation, 17(81):100–103, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/
journals/mcom/1963-17-081/S0025-

                                              5718-63-99185-3;   http://www.ams.
                                              org/journals/mcom/1963-17-081/S0025-
                                              5718-63-99185-3/S0025-5718-63-99185-
                                              3.pdf.

Anonymous:1963:TEb

[Ano63i] Anonymous. Table errata. Mathematics of Computation, 17(81):100– 103, January 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

                                              /www.ams.org/journals/mcom/1963-
                                              17-081/S0025-5718-1963-1781100-6;
                                               http://www.ams.org/journals/mcom/
                                              1963-17-081/S0025-5718-1963-1781100-
                                              6/S0025-5718-1963-1781100-6.pdf;
                                               https://www.ams.org/mathscinet-
                                              getitem?mr=1781100.

Anonymous:1963:TEc

[Ano63j] Anonymous. Table errata. Mathematics of Computation, 17(82):214– 215, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

                                              tronic). URL http://www.ams.org/
                                              journals/mcom/1963-17-082/S0025-
                                              5718-63-99181-6;   http://www.ams.
                                              org/journals/mcom/1963-17-082/S0025-
                                              5718-63-99181-6/S0025-5718-63-99181-
                                              6.pdf.

Anonymous:1964:Cb

[Ano64a] Anonymous. Corrigenda. Mathematics of Computation, 18(87):536, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1964:Ca

[Ano64b] Anonymous. Corrigendum. Mathematics of Computation, 18(86):348, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1964:Cc

[Ano64c] Anonymous. Corrigendum. Mathematics of Computation, 18(88):688, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1965:Ca

[Ano65a] Anonymous. Corrigenda. Mathematics of Computation, 19(89):175, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1965:Cb

[Ano65b] Anonymous. Corrigendum. Mathematics of Computation, 19(91):529, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1965:Cc

[Ano65c] Anonymous. Corrigendum. Mathematics of Computation, 19(92):706, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1965:Ea

[Ano65d] Anonymous. Errata. Mathematics of Computation, 19(91):527–528, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1965:Eb

[Ano65e] Anonymous. Errata. Mathematics of Computation, 19(92):705–706, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1966:Cc

[Ano66a] Anonymous. Corrigenda. Mathematics of Computation, 20(96):643, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1966:Cb

[Ano66b] Anonymous. Corrigendum. Mathematics of Computation, 20(94):344, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1966:E

[Ano66c] Anonymous. Errata. Mathematics of Computation, 20(95):468–471, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1967:Ca

[Ano67a] Anonymous. Corrigenda. Mathematics of Computation, 21(97):130, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1967:Cb

[Ano67b] Anonymous. Corrigenda. Mathematics of Computation, 21(98):296, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1967:Cc

[Ano67c] Anonymous. Corrigendum. Mathematics of Computation, 21(100):751, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1968:Ca

[Ano68a] Anonymous. Corrigenda. Mathematics of Computation, 22(101):246–247, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1968:Cb

[Ano68b] Anonymous. Corrigenda. Mathematics of Computation, 22(102):474, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1968:Cc

[Ano68c] Anonymous. Corrigenda. Mathematics of Computation, 22(103):699, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1968:Cd

[Ano68d] Anonymous. Corrigenda. Mathematics of Computation, 22(104):911, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1969:Ca

[Ano69a] Anonymous. Corrigenda. Mathematics of Computation, 23(105):219, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1969:Cb

[Ano69b] Anonymous. Corrigenda. Mathematics of Computation, 23(106):473, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1969:Cd

[Ano69c] Anonymous. Corrigenda. Mathematics of Computation, 23(108):893, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1969:Cc

[Ano69d] Anonymous. Corrigendum. Mathematics of Computation, 23(107):693, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alanen:1967:SCA

[AØS67] J. Alanen, Ø. Øre, and J. Stemple. Systematic computations on amicable numbers (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):242–245, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Apelkrans:1968:DSH

[Ape68] Mats Y. T. Apelkrans. On difference schemes for hyperbolic equations with discontinuous initial values. Mathematics of Computation, 22(103):525– 539, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Aumann:1965:MCK

[APR65] R. J. Aumann, B. Peleg, and P. Rabinowitz. A method for computing the kernel of n-person games. Mathematics of Computation, 19(92):531–551, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ascher:1960:ESO

[Asc60] Marcia Ascher. Explicit solutions of the one-dimensional heat equation for a composite wall. Mathematics of Computation, 14(72):346–353, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Atkin:1968:EGS

[Atk68] A. O. L. Atkin. Explicit gap series at cusps of Γ(p) (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):416–419, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ayoub:1967:NCN

[Ayo67] Raymond G. Ayoub. A note on the class number of imaginary quadratic fields. Mathematics of Computation, 21(99): 442–445, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Abarbanel:1969:IFD

[AZ69] S. Abarbanel and G. Zwas. An iterative finite-difference method for hyperbolic systems. Mathematics of Computation, 23(107):549–565, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bridge:1962:ETR

[BA62] John F. Bridge and Stanley W. Angrist. An extended table of roots of J′_{n}(x)Y′_{n}(βx) − J′_{n}(βx)Y′_{n}(x) = 0. Mathematics of Computation, 16(78):198– 204, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barakat:1961:EIG

[Bar61] Richard Barakat. Evaluation of the incomplete gamma function of imaginary argument by Chebyshev polynomials. Mathematics of Computation, 15(73):7– 11, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnett:1962:BPR

[Bar62] V. D. Barnett. The behavior of pseudorandom sequences generated on computers by the multiplicative congruential method. Mathematics of Computation, 16(77):63–69, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barrodale:1966:NES

[Bar66] Ian Barrodale. A note on equal sums of like powers (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):318–322, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bard:1968:NID

[Bar68a] Yonathan Bard. On a numerical instability of Davidon-like methods (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):665– 666, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bareiss:1968:SIM

[Bar68b] Erwin H. Bareiss. Sylvester’s identity and multistep integer-preserving Gaussian elimination. Mathematics of Computation, 22(103):565–578, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1968:EANa

[Bar68c] Robert E. Barnhill. An error analysis for numerical multiple integration. I. Mathematics of Computation, 22(101): 98–109, January 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1968:EANb

[Bar68d] Robert E. Barnhill. An error analysis for numerical multiple integration. II. Mathematics of Computation, 22(102): 286–292, April 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Batten:1963:SOC

[Bat63] George W. Batten, Jr. Second-order correct boundary conditions for the numerical solution of the mixed boundary problem for parabolic equations. Mathematics of Computation, 17(84):405–413, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1960:FCL

[Bau60] F. L. Bauer. The formula-controlled logical computer “Stanislaus” (in Technical Notes and Short Papers). Mathematics of Computation, 14(69):64–67, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bauer:1964:TZC

[Bau64] Helmut F. Bauer. Tables of zeros of cross product Bessel functions J′_{p}(ξ)Y′_{p}(kξ) − J′_{p}(kξ)Y′_{p}(ξ) = 0 (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):128–135, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baer:1968:NSP

[BB68] R. M. Baer and P. Brock. Natural sorting over permutation spaces. Mathematics of Computation, 22(102):385– 410, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bergman:1967:PCM

[BC67] Stefan Bergman and Bruce Chalmers.

A procedure for conformal mapping of triply-connected domains. Mathematics of Computation, 21(100):527–542, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blanch:1969:DPM

[BC69] G. Blanch and D. S. Clemm. The double points of Mathieu’s differential equation. Mathematics of Computation, 23(105): 97–108, January 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1969:NTC

[BDN69] R. E. Barnhill, J. E. Dennis, Jr., and G. M. Nielson. A new type of Chebyshev quadrature (in Technical Notes and Short Papers). Mathematics of Computation, 23(106):437–441, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berger:1966:IDL

[Ber66] Bruce S. Berger. Inversion of the Ndimensional Laplace transform (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):418– 421, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bergland:1967:FFT

[Ber67] G. D. Bergland. The Fast Fourier Transform recursive equations for arbitrary length records (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):236–238, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bergland:1968:FFT

[Ber68] G. D. Bergland. A Fast Fourier Transform algorithm using base 8 iterations. Mathematics of Computation, 22(102): 275–279, April 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Best:1961:TTT

[Bes61] Gilbert C. Best. Two theorem tables of matrix algebra. Mathematics of Computation, 15(73):19–22, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Best:1964:HFI

[Bes64] Gilbert C. Best. Helpful formulas for integrating polynomials in three dimensions (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):310–312, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Best:1968:PMS

[Bes68] Gilbert C. Best. Powers of a matrix of special type (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):667–668, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borosh:1966:ESL

[BF66] I. Borosh and A. S. Fraenkel. Exact solutions of linear equations with rational coefficients by congruence techniques. Mathematics of Computation, 20 (93):107–112, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baumert:1967:CNO

[BF67a] L. D. Baumert and H. Fredricksen.

The cyclotomic numbers of order eighteen with applications to difference sets. Mathematics of Computation, 21(98): 204–219, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bowen:1967:GTS

[BF67b] Robert Bowen and Stephen Fisk. Generation of triangulations of the sphere (in Technical Notes and Short Papers). Mathematics of Computation, 21(98): 250–252, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bateman:1962:HAF

[BH62a] Paul T. Bateman and Roger A. Horn.

A heuristic asymptotic formula concerning the distribution of prime numbers (in Technical Notes and Short Papers). Mathematics of Computation, 16(79): 363–367, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1962:SWA

[BH62b] K. M. Brown and P. Henrici. Sign wave analysis in matrix eigenvalue problems. Mathematics of Computation, 16(79): 291–300, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1964:NMT

[BH64] James H. Bramble and Bert E. Hubbard. New monotone type approximations for elliptic problems. Mathematics of Computation, 18(87):349–367, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Baumert:1965:HMW

[BH65] L. D. Baumert and Marshall Hall, Jr. Hadamard matrices of the Williamson type. Mathematics of Computation, 19(91):442–447, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bergman:1968:NCT

[BHK68] S. Bergman, J. G. Herriot, and T. G. Kurtz. On numerical calculation of transonic flow patterns. Mathematics of Computation, 22(101):13–27, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1969:CEE

[BHT69] J. H. Bramble, B. E. Hubbard, and Vidar Thomee. Convergence estimates for essentially positive type discrete Dirichlet problems. Mathematics of Computation, 23(108):695–709, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ben-Israel:1965:IMC

[BI65] Adi Ben-Israel. An iterative method for computing the generalized inverse of an arbitrary matrix (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):452–455, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ben-Israel:1966:NIM

[BI66] Adi Ben-Israel. A note on an iterative method for generalized inversion of matrices (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):439–440, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1960:FCM

[BJ60] John Brillhart and G. D. Johnson. On the factors of certain Mersenne numbers (in Technical Notes and Short Papers). Mathematics of Computation, 14(72): 365–369, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bellman:1962:NSD

[BK62] R. Bellman and B. Kotkin. On the numerical solution of a differentialdifference equation arising in analytic number theory (in Technical Notes and Short Papers). Mathematics of Computation, 16(80):473–475, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boas:1964:CTT

[BK64] R. P. Boas, Jr. and Virginia C. Klema. A constant in the theory of trigonometric series (in Technical Notes and Short Papers). Mathematics of Computation, 18 (88):674, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Banks:1968:CES

[BK68] D. O. Banks and G. J. Kurowski. Computation of eigenvalues of singular Sturm–Liouville systems. Mathematics of Computation, 22(102):304–310, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bellman:1963:PAN

[BKK63] Richard Bellman, Robert Kalaba, and Bella Kotkin. Polynomial approximation — a new computational technique in dynamic programming: Allocation processes. Mathematics of Computation, 17(82):155–161, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bellman:1964:DAA

[BKK64] Richard Bellman, Robert Kalaba, and Bella Kotkin. Differential approximation applied to the solution of convolution equations (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):487–491, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1964:BPC

[BLL64] John Brillhart, D. H. Lehmer, and Emma Lehmer. Bounds for pairs of consecutive seventh and higher power residues. Mathematics of Computation, 18(87):397–407, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blum:1962:MRK

[Blu62]  E. K. Blum.  A modification of

the Runge–Kutta fourth-order method. Mathematics of Computation, 16(78): 176–187, April 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bratley:1968:MAN

[BM68] Paul Bratley and John McKay. More amicable numbers (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):677–678, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1960:CFI

[Boe60] J. Boersma. Computation of Fresnel integrals (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):380, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1961:TFR

[Boe61] J. Boersma. Two formulas relating to elliptic integrals of the third kind (in Technical Notes and Short Papers). Mathematics of Computation, 15(75): 296–298, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1962:CLF

[Boe62] J. Boersma. On the computation of Lommel’s functions of two variables (in Technical Notes and Short Papers). Mathematics of Computation, 16(78): 232–238, April 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1969:ECW

[Boe69] J. Boersma. Expansions for Coulomb wave functions. Mathematics of Computation, 23(105):51–59, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bowen:1967:GMT

[Bow67] Robert Bowen. The generation of minimal triangle graphs (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):248–250, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ballester:1967:CDA

[BP67] C. Ballester and V. Pereyra. On the construction of discrete approximations to linear differential expressions. Mathematics of Computation, 21(99):297– 302, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bramble:1963:FOF

[Bra63] James H. Bramble. Fourth-order finite difference analogues of the Dirichlet problem for Poisson’s equation in three and four dimensions. Mathematics of Computation, 17(83):217–222, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brandt:1966:EDQ

[Bra66] Achi Brandt. Estimates for difference quotients of solutions of Poisson type difference equations. Mathematics of Computation, 20(96):473–499, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1965:SPA

[BRB65] Robert R. Brown, James D. Riley, and Morris M. Bennett. Stability properties of Adams–Moulton type methods. Mathematics of Computation, 19(89): 90–96, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1962:CNP

[Bri62] John Brillhart. Concerning the numbers 2^{2p} + 1, p prime. Mathematics of Computation, 16(80):424–430, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1963:SMF

[Bri63] John Brillhart. Some miscellaneous factorizations (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):447–450, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1964:FCM

[Bri64] John Brillhart. On the factors of certain Mersenne numbers, II. Mathematics of Computation, 18(85):87–92, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broyden:1964:CCM

[Bro64] C. G. Broyden. On convergence criteria for the method of successive overrelaxation (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):136–141, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broyden:1965:CMS

[Bro65] C. G. Broyden. A class of methods for solving nonlinear simultaneous equations. Mathematics of Computation, 19 (92):577–593, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Broyden:1967:QNM

[Bro67] C. G. Broyden. Quasi-Newton methods and their application to function minimisation. Mathematics of Computation, 21(99):368–381, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brown:1969:ADS

[Bro69] Harold Brown. An algorithm for the determination of space groups. Mathematics of Computation, 23(107):499– 514, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brudno:1969:SNR

[Bru69] Simcha Brudno. Some new results on equal sums of like powers. Mathematics of Computation, 23(108):877– 880, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bass:1967:FDM

[BS67a] L. J. Bass and S. R. Schubert. On finding the disc of minimum radius containing a given set of points (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):712– 714, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1967:SFR

[BS67b] John Brillhart and J. L. Selfridge. Some factorizations of 2^{n} ± 1 and related results. Mathematics of Computation, 21(97):87–96, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1962:NCF

[Bur62a] F. D. Burgoyne. A note on the curve fitting of discrete data by economization (in Technical Notes and Short Papers). Mathematics of Computation, 16(80): 498–499, October 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1962:PAR

[Bur62b] F. D. Burgoyne. Polynomial approximations to I_{0}(x), I_{1}(x) and related functions (in Technical Notes and Short Papers). Mathematics of Computation, 16 (80):497–498, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1963:AKF

[Bur63a] F. D. Burgoyne. Approximations to Kelvin functions (in Technical Notes and Short Papers). Mathematics of Computation, 17(83):295–298, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1963:NEC

[Bur63b] F. D. Burgoyne. The non-existence of certain Laguerre–Chebyshev quadrature formulas. Mathematics of Computation, 17(82):196–197, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1963:QFI

[Bur63c] F. D. Burgoyne. Quadrature formulas over infinite intervals in terms of differences (in Technical Notes and Short Papers). Mathematics of Computation, 17(83):298–301, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1964:GTF

[Bur64] F. D. Burgoyne. Generalized trigonometric functions (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):314–316, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burrows:1966:MSD

[Bur66] James W. Burrows. Maximization of a second-degree polynomial on the unit sphere (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):441–444, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1967:PPA

[Bur67] F. D. Burgoyne. Practical L^{p} polynomial approximation (in technical notes and short notices). Mathematics of Computation, 21(97):113–115, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1968:BIP

[Bur68] F. D. Burgoyne. Bivariate interpolation of potential functions. Mathematics of Computation, 22(103):589–594, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buschman:1968:SFL

[Bus68] R. G. Buschman. Substitution formulas for Laplace transformations (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):209– 211, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Businger:1969:RMH

[Bus69a] P. A. Businger. Reducing a matrix to Hessenberg form. Mathematics of Computation, 23(108):819–821, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Businger:1969:EPB

[Bus69b] Peter A. Businger. Extremal properties of balanced tri-diagonal matrices (in Technical Notes and Short Papers).

Mathematics of Computation, 23(105): 193–195, January 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1961:PTP

[But61] J. C. Butcher. A partition test for pseudo-random numbers (in Technical Notes and Short Papers). Mathematics of Computation, 15(74):198–199, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1964:IRK

[But64a] J. C. Butcher. Implicit Runge–Kutta processes. Mathematics of Computation, 18(85):50–64, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1964:IPB

[But64b] J. C. Butcher. Integration processes based on Radau quadrature formulas. Mathematics of Computation, 18(86): 233–244, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1965:AOR

[But65] J. C. Butcher. On the attainable order of Runge–Kutta methods. Mathematics of Computation, 19(91):408–417, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Butcher:1966:CNS

[But66] J. C. Butcher. On the convergence of numerical solutions to ordinary differential equations. Mathematics of Computation, 20(93):1–10, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1967:QRMa

[BW67a] R. E. Barnhill and J. A. Wixom. Quadratures with remainders of minimum norm. I. Mathematics of Computation, 21(97):66–75, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Barnhill:1967:QRMb

[BW67b] R. E. Barnhill and J. A. Wixom. Quadratures with remainders of minimum norm. II. Mathematics of Computation, 21(99):382–387, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cadwell:1964:AIC

[Cad64] J. H. Cadwell. An approximation to the integral of the circular Gaussian distribution over an offset ellipse. Mathematics of Computation, 18(85):106–112, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlitz:1962:CLF

[Car62] L. Carlitz. The coefficients of the lemniscate function (in Technical Notes and Short Papers). Mathematics of Computation, 16(80):475–478, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlitz:1965:IBP

[Car65] L. Carlitz. The irreducibility of the Bernoulli polynomial B_{14}(x) (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):667–670, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carasso:1969:FDM

[Car69] Alfred Carasso. Finite-difference methods and the eigenvalue problem for nonselfadjoint Sturm–Liouville operators. Mathematics of Computation, 23(108): 717–729, October 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cassity:1965:ACE

[Cas65] C. R. Cassity. Abscissas, coefficients, and error term for the generalized Gauss–Laguerre quadrature formula using the zero ordinate. Mathematics of Computation, 19(90):287–296, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Case:1969:NSP

[Cas69] James Case. A note on the stability of predictor-corrector techniques. Mathematics of Computation, 23(108):741– 749, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cavior:1963:NOP

[Cav63] S. R. Cavior. A note on octic permutation polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):450–452, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Concus:1963:TEG

[CCJM63] P. Concus, D. Cassatt, G. Jaehnig, and E. Melby. Tables for the evaluation of ∫_{0}^{∞} x^{β}e^{−x}f(x)dx by Gauss–Laguerre quadrature. Mathematics of Computation, 17(83):245–256, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cantor:1962:VHS

[CEFT62] D. G. Cantor, G. Estrin, A. S. Fraenkel, and R. Turn. A very high-speed digital number sieve. Mathematics of Computation, 16(78):141–154, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cheema:1968:MTN

[CG68] M. S. Cheema and H. Gupta. The maxima of P_{r}(n_{1}, n_{2}) (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):199–200, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1967:CAN

[CH67] W. J. Cody and K. E. Hillstrom. Chebyshev approximations for the natural logarithm of the gamma function. Mathematics of Computation, 21(98):198– 203, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Charmonman:1967:EAI

[Cha67] S. Charmonman. An efficient algorithm for inverting a block-symmetric matrix (in Technical Notes and Short Papers). Mathematics of Computation, 21(100): 715–717, October 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chak:1968:PSS

[Cha68a] A. M. Chak. Polynomial system satisfying a special functional equation (in Technical Notes and Short Papers). Mathematics of Computation, 22(103): 673–676, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chawla:1968:EBG

[Cha68b] M. M. Chawla. Error bounds for the Gauss–Chebyshev quadrature formula of the closed type (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):889–891, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chawla:1968:EEC

[Cha68c] M. M. Chawla. Error estimates for the Clenshaw–Curtis quadrature. Mathematics of Computation, 22(103):651– 656, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cheema:1964:VPC

[Che64] M. S. Cheema. Vector partitions and combinatorial identities. Mathematics of Computation, 18(87):414–420, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chiang:1969:CIL

[Chi69] P. S. Chiang. Computer investigation of Landau’s theorem (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):185–188, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cho:1968:TDC

[Cho68a] Choong Yun Cho. On the triangular decomposition of Cauchy matrices. Mathematics of Computation, 22(104):819– 825, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chorin:1968:NSN

[Cho68b] Alexandre Joel Chorin. Numerical solution of the Navier–Stokes equations. Mathematics of Computation, 22(104):

745–762, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chorin:1969:CDA

[Cho69] Alexandre Joel Chorin. On the convergence of discrete approximations to the Navier–Stokes equations. Mathematics of Computation, 23(106):341–353, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Church:1965:CTT

[Chu65] Robert F. Church. On a constant in the theory of trigonometric series (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):501, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Charmonman:1968:EIC

[CJ68a] S. Charmonman and R. S. Julius. Explicit inverses and condition numbers of certain circulants (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):428–430, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chawla:1968:AEE

[CJ68b] M. M. Chawla and M. K. Jain. Asymptotic error estimates for the Gauss quadrature formula. Mathematics of Computation, 22(101):91–97, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chawla:1968:EEG

[CJ68c] M. M. Chawla and M. K. Jain. Error estimates for Gauss quadrature formulas for analytic functions. Mathematics of Computation, 22(101):82–90, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Campbell:1968:SDA

[CK68] C. M. Campbell and P. Keast. The stability of difference approximations to a self-adjoint parabolic equation, under derivative boundary conditions. Mathematics of Computation, 22(102):336– 346, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Clark:1962:CMM

[Cla62] G. Miller Clark. Corrections to Mervin E. Muller, A Note on a Method for Generating Points Uniformly on N-Dimensional Spheres. Mathematics of Computation, 16(78):261, 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Mul59].

Cockayne:1969:CGG

[Coc69a] E. J. Cockayne. Computation of Galois group elements of a polynomial equation (in Technical Notes and Short Papers). Mathematics of Computation, 23(106): 425–428, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). SUPPLEMENT on microfiche.

Cockayne:1969:CML

[Coc69b] E. J. Cockayne. Computation of minimal length full Steiner trees on the vertices of a convex polygon. Mathematics of Computation, 23(107):521–531, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1965:CAC

[Cod65a] W. J. Cody. Chebyshev approximations for the complete elliptic integrals K and E. Mathematics of Computation, 19(89): 105–112, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigenda [Cod66].

Cody:1965:CPE

[Cod65b] W. J. Cody. Chebyshev polynomial expansions of complete elliptic integrals. Mathematics of Computation, 19(90): 249–259, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1966:CCA

[Cod66] W. J. Cody. Corrigenda: “Chebyshev approximations for the complete elliptic integrals K and E”. Mathematics of Computation, 20(93):207, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Cod65a].

Cody:1968:CAF

[Cod68] W. J. Cody. Chebyshev approximations for the Fresnel integrals (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):450–453, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1969:RCA

[Cod69] W. J. Cody. Rational Chebyshev approximations for the error function. Mathematics of Computation, 23(107): 631–637, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1962:NSR

[Coh62] Harvey Cohn. A numerical study of the relative class numbers of real quadratic integral domains. Mathematics of Computation, 16(78):127–140, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1965:NSF

[Coh65] Harvey Cohn. A numerical survey of the floors of various Hilbert fundamental domains. Mathematics of Computation, 19(92):594–605, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohen:1967:SST

[Coh67a] Edward L. Cohen. On the sum of the squares of two consecutive integers. Mathematics of Computation, 21(99): 460–465, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1967:NST

[Coh67b] Harvey Cohn. A numerical study of topological features of certain Hilbert fundamental domains. Mathematics of Computation, 21(97):76–86, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cohn:1969:SCA

[Coh69] Harvey Cohn. Some computer-assisted topological models of Hilbert fundamental domains. Mathematics of Computation, 23(107):475–487, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Collins:1969:CMI

[Col69] George E. Collins. Computing multiplicative inverses in GF(p) (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):197– 200, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Comet:1960:IMC

[Com60] Stig Comet. Improved methods to calculate the characters of the symmetric group. Mathematics of Computation, 14(70):104–117, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Concus:1967:NSM

[Con67] Paul Concus. Numerical solution of the minimal surface equation. Mathematics of Computation, 21(99):340–350, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooley:1961:IEC

[Coo61] J. W. Cooley. An improved eigenvalue corrector formula for solving the Schrödinger equation for central fields. Mathematics of Computation, 15(76): 363–374, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1967:CSS

[Coo67] G. J. Cooper. A class of single-step methods for systems of nonlinear differential equations. Mathematics of Computation, 21(100):597–610, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooper:1968:IQM

[Coo68] G. J. Cooper. Interpolation and quadrature methods for ordinary differential equations. Mathematics of Computation, 22(101):69–76, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Corrington:1961:ACE

[Cor61] Murlan S. Corrington. Applications of the complex exponential integral. Mathematics of Computation, 15(73):1– 6, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chiarella:1968:EIR

[CR68] C. Chiarella and A. Reichel. On the evaluation of integrals related to the error function. Mathematics of Computation, 22(101):137–143, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cooley:1965:AMC

[CT65] James W. Cooley and John W. Tukey.

An algorithm for the machine calculation of complex Fourier series. Mathematics of Computation, 19(90):297– 301, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1967:CRC

[CT67a] W. J. Cody and Henry C. Thacher, Jr. Corrigendum: “Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2”. Mathematics of Computation, 21(99): 525, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/
stable/2003289. See [CT67b].

Cody:1967:RCA

[CT67b] W. J. Cody and Henry C. Thacher, Jr.

Rational Chebyshev approximations for Fermi–Dirac integrals of orders −\frac{1}{2}, \frac{1}{2} and \frac{3}{2}. Mathematics of Computation, 21(97):30–40, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2003468. See

also [CT67a].

Cody:1968:RCA

[CT68] W. J. Cody and Henry C. Thacher, Jr.

Rational Chebyshev approximations for the exponential integral E_{1}(x). Mathematics of Computation, 22(103):641– 649, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1969:CAE

[CT69] W. J. Cody and Henry C. Thacher, Jr. Chebyshev approximations for the exponential integral Ei(x). Mathematics of Computation, 23(106):289–303, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daly:1963:SPC

[Dal63] Bart J. Daly. The stability properties of a coupled pair of non-linear partial difference equations. Mathematics of Computation, 17(84):346–360, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daniel:1969:AMF

[Dan69a] James W. Daniel. On the approximate minimization of functionals. Mathematics of Computation, 23(107):573– 581, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Daniel:1969:SSP

[Dan69b] James W. Daniel. Summation of series of positive terms by condensation transformations. Mathematics of Computation, 23(105):91–96, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Davis:1964:TFC

[Dav64] Philip J. Davis. Triangle formulas in the complex plane. Mathematics of Computation, 18(88):569–577, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Davis:1967:CNA

[Dav67] Philip J. Davis. A construction of nonnegative approximate quadratures. Mathematics of Computation, 21(100):

578–582, October 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Day:1963:GQM

[Day63] J. T. Day. A Gaussian quadrature method for the numerical solution of the characteristic initial value problem u_{xy} = f(x, y, u) (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):438–441, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Day:1964:OSM

[Day64] J. T. Day. A one-step method for the numerical solution of second order linear ordinary differential equations (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):664–668, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Day:1966:NBA

[Day66] J. T. Day. A note on best approximation in E^{n} (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):599–600, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Day:1967:SMS

[Day67] J. T. Day. A starting method for solving nonlinear Volterra integral equations. Mathematics of Computation, 21(98): 179–188, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dempsey:1960:NAE

[DB60] E. Dempsey and G. C. Benson. Note on the asymptotic expansion of the modified Bessel function of the second kind (in Technical Notes and Short Papers).

Mathematics of Computation, 14(72): 362–365, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBruijn:1964:SDD

[de 64] N. G. de Bruijn. Some direct decompositions of the set of integers. Mathematics of Computation, 18(88):537–546, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBalbine:1967:NRP

[de 67] Guy de Balbine. Note on random permutations (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):710–712, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Denman:1962:NLR

[DE62] H. H. Denman and R. C. W. Ettinger. Note on latent roots and vectors of segments of the Hilbert matrix (in Technical Notes and Short Papers). Mathematics of Computation, 16(79):370–371, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delves:1965:VPA

[Del65] L. M. Delves. Variation principles for an arbitrary operator. III. Mathematics of Computation, 19(91):380–386, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delves:1966:ED

[Del66] L. M. Delves. Exponential differences. Mathematics of Computation, 20(94): 246–256, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Denman:1966:MPA

[Den66] Harry H. Denman. Minimax polynomial approximation. Mathematics of Computation, 20(94):257–265, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Den:1969:SSC

[Den69] Chi Fu Den. Summation of a slowly convergent series arising in antenna study. Mathematics of Computation, 23(107): 651–654, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Descloux:1963:NRE

[Des63] J. Descloux. Note on the round-off errors in iterative processes. Mathematics of Computation, 17(81):18–27, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

deBalbine:1966:CFI

[dF66] Guy de Balbine and Joel N. Franklin.

The calculation of Fourier integrals. Mathematics of Computation, 20(96): 570–589, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1963:THO

[DG63] Jim Douglas, Jr. and James E. Gunn.

Two high-order correct difference analogues for the equation of multidimensional heat flow. Mathematics of Computation, 17(81):71–80, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Decell:1966:CNS

[DGL66] H. P. Decell, Jr., L. F. Guseman, Jr., and R. N. Lea. Concerning the numerical solution of differential equations (in Technical Notes and Short Papers).

Mathematics of Computation, 20(95): 431–434, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dickey:1968:IMN

[Dic68] R. W. Dickey. An improved method for the numerical solution of the suspension bridge deflection equations. Mathematics of Computation, 22(102):298– 303, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1961:IGB

[DJ61] A. R. DiDonato and M. P. Jarnagin. Integration of the general bivariate Gaussian distribution over an offset circle. Mathematics of Computation, 15(76): 375–382, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1962:MCC

[DJ62] A. R. DiDonato and M. P. Jarnagin. A method for computing the circular coverage function. Mathematics of Computation, 16(79):347–355, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1967:ECI

[DJ67] A. R. DiDonato and M. P. Jarnagin. The efficient calculation of the incomplete beta-function ratio for half-integer values of the parameters a, b. Mathematics of Computation, 21(100):652– 662, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Douglas:1963:ADI

[DKV63] J. Douglas, Jr., R. B. Kellogg, and R. S. Varga. Alternating direction iteration methods for n space variables (in Technical Notes and Short Papers). Mathematics of Computation, 17(83): 279–282, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Delves:1967:NML

[DL67] L. M. Delves and J. N. Lyness. A numerical method for locating the zeros of an analytic function. Mathematics of Computation, 21(100):543–560, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dulmage:1962:ISM

[DM62] A. L. Dulmage and N. S. Mendelsohn.

On the inversion of sparse matrices (in Technical Notes and Short Papers). Mathematics of Computation, 16(80): 494–496, October 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Diaz:1969:SAS

[DM69] J. B. Diaz and H. W. McLaughlin. Simultaneous approximation of a set of bounded real functions. Mathematics of Computation, 23(107):583–593, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Doring:1966:CZC

[Dor66] Boro Doring. Complex zeros of cylinder functions. Mathematics of Computation, 20(94):215–222, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dorr:1969:RIS

[Dor69] Fred W. Dorr. Remarks on the iterative solution of the Neumann problem on a rectangle by successive line over-relaxation (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):177–179, January 1969.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Derzko:1965:BSR

[DP65] N. A. Derzko and A. M. Pfeffer. Bounds for the spectral radius of a matrix. Mathematics of Computation, 19(89): 62–67, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dranoff:1961:EPA

[Dra61] J. S. Dranoff. An eigenvalue problem arising in mass and heat transfer studies. Mathematics of Computation, 15(76): 403–409, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dahlquist:1968:CMA

[DSS68] G. Dahlquist, B. Sjoberg, and P. Svensson. Comparison of the method of averages with the method of least squares. Mathematics of Computation, 22(104): 833–845, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Duris:1966:SSC

[Dur66] C. S. Duris. A simplex sufficiency condition for quadrature formulas. Mathematics of Computation, 20(93):68–78, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eberlein:1964:TPT

[Ebe64] P. J. Eberlein. A two parameter test matrix (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):296–298, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Edelblute:1966:MIR

[Ede66] David J. Edelblute. Matrix inversion by rank annihilation (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):149–151, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erber:1963:TF

[EG63] Thomas Erber and Alan Gordon. Tabulation of the functions \frac{∂i_{ν}(z)}{∂ν}, ν = ±\frac{1}{3}. Mathematics of Computation, 17(82): 162–169, April 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ehrman:1967:NPD

[Ehr67] John R. Ehrman. The number of prime divisors of certain Mersenne numbers. Mathematics of Computation, 21(100): 700–704, October 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eisen:1968:CCD

[Eis68] Dennis Eisen. Consistency conditions for difference schemes with singular coefficients. Mathematics of Computation, 22(102):347–351, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1964:EEC

[Ell64] David Elliott. The evaluation and estimation of the coefficients in the Chebyshev series expansion of a function. Mathematics of Computation, 18(86): 274–284, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1965:TET

[Ell65] David Elliott. Truncation errors in two Chebyshev series approximations.

Mathematics of Computation, 19(90): 234–248, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1967:TEP

[Ell67] David Elliott. Truncation errors in Padé approximations to certain functions: An alternative approach. Mathematics of Computation, 21(99):398– 406, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Epstein:1962:RCS

[Eps62] Bernard Epstein. A remark concerning the solution of the Dirichlet problem by finite differences (in Technical Notes and Short Papers). Mathematics of Computation, 16(77):110–112, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Erdos:1964:PEN

[Erd64] P. Erdős. On a problem in elementary number theory and a combinatorial problem (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):644–646, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Elliott:1965:SEC

[ES65] David Elliott and George Szekeres. Some estimates of the coefficients in the Chebyshev series expansion of a function. Mathematics of Computation, 19 (89):25–32, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fair:1964:PAS

[Fai64] Wyman Fair. Padé approximation to the solution of the Ricatti equation. Mathematics of Computation, 18(88):

627–634, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fairweather:1969:NGM

[Fai69] Graeme Fairweather. A note on a generalisation of a method of Douglas (in Technical Notes and Short Papers). Mathematics of Computation, 23(106): 407–409, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1967:EEH

[FC67] Henry E. Fettis and James C. Caslin.

Eigenvalues and eigenvectors of Hilbert matrices of order 3 through 10. Mathematics of Computation, 21(99):431– 441, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1960:N

[Fet60a] H. E. Fettis. Note on ∫_{0}^{∞} e^{−x}J_{0} (\frac{ηx}{ϵ}) J_{1} (\frac{x}{ϵ}) x^{−n}dx.

Mathematics of Computation, 14(72): 372–374, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1960:NTN

[Fet60b] Henry E. Fettis. Note on ∫_{0}^{∞} e^{−x}J_{0}(\frac{ηx}{ξ})J_{1}(\frac{x}{ξ})x^{−n}dx (in Technical Notes and Short Papers). Mathematics of Computation, 14(72): 372–374, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1964:NSE

[Fet64] Henry E. Fettis. On the numerical solution of equations of the Abel type (in Technical Notes and Short Papers). Mathematics of Computation, 18(87): 491–496, July 1964. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1965:CEI

[Fet65] Henry E. Fettis. Calculation of elliptic integrals of the third kind by means of Gauss’ transformation. Mathematics of Computation, 19(89):97–104, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1967:MCI

[Fet67] Henry E. Fettis. More on the calculation of the integral I_{n}(b) = \frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos bxdx (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):727–730, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feuchter:1968:NIS

[Feu68] Christopher A. Feuchter. Numerical integration over a sphere. Mathematics of Computation, 22(102):293–297, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fraser:1962:CNL

[FG62] W. Fraser and C. C. Gotlieb. A calculation of the number of lattice points in the circle and sphere. Mathematics of Computation, 16(79):282–290, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fairweather:1967:SSD

[FG67] G. Fairweather and A. R. Gourlay. Some stable difference approximations to a fourth-order parabolic partial differential equation. Mathematics of Computation, 21(97):1–11, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fielder:1960:TCO

[Fie60] Daniel C. Fielder. Tabulation of coefficients for operations on Taylor series. Mathematics of Computation, 14(72): 339–345, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fields:1965:RAG

[Fie65] Jerry L. Fields. Rational approximations to generalized hypergeometric functions. Mathematics of Computation, 19(92): 606–624, October 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fields:1967:CE

[Fie67] Jerry L. Fields. Confluent expansions.

Mathematics of Computation, 21(98): 189–197, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fisher:1964:MET

[Fis64] Donald D. Fisher. Minimum ellipsoids (in Technical Notes and Short Papers). Mathematics of Computation, 18(88): 669–673, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Filler:1961:SAI

[FL61] L. Filler and H. F. Ludloff. Stability analysis and integration of the viscous equations of motion. Mathematics of Computation, 15(75):261–274, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fair:1966:RAS

[FL66] Wyman Fair and Yudell L. Luke. Rational approximations to the solution of the second order Riccati equation (in Technical Notes and Short Papers).

Mathematics of Computation, 20(96): 602–606, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fair:1967:RAI

[FL67] Wyman G. Fair and Yudell L. Luke. Rational approximations to the incomplete elliptic integrals of the first and second kinds. Mathematics of Computation, 21(99):418–422, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Flammer:1968:A

[Fla68] Carson Flammer. Addendum. Mathematics of Computation, 22(101):249, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fleckner:1968:MCF

[Fle68] Oscar L. Fleckner. A method for the computation of the Fresnel integrals and related functions. Mathematics of Computation, 22(103):635–640, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Floyd:1960:NRA

[Flo60] Robert W. Floyd. A note on rational approximation. Mathematics of Computation, 14(69):72–73, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Forsythe:1964:TPA

[For64] George E. Forsythe. Tests of Parlett’s ALGOL eigenvalue procedure eig3 (in Technical Notes and Short Papers). Mathematics of Computation, 18(87): 486–487, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fosdick:1965:ACW

[Fos65] Lloyd D. Fosdick. Approximation of a class of Wiener integrals. Mathematics of Computation, 19(90):225–233, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fosdick:1968:SCF

[Fos68] Lloyd D. Fosdick. A special case of the Filon quadrature formula. Mathematics of Computation, 22(101):77–81, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franklin:1963:DSR

[Fra63] Joel N. Franklin. Deterministic simulation of random processes. Mathematics of Computation, 17(81):28–59, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/
stable/2003733.

Franklin:1964:EMP

[Fra64] Joel N. Franklin. Equidistribution of matrix-power residues modulo one. Mathematics of Computation, 18(88): 560–568, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franke:1965:NEE

[Fra65a] Charles H. Franke. Numerical evaluation of the elliptic integral of the third kind (in Technical Notes and Short Papers). Mathematics of Computation, 19(91): 494–496, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franklin:1965:DMS

[Fra65b] Joel N. Franklin. Difference methods for stochastic ordinary differential equations. Mathematics of Computation, 19 (92):552–561, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frank:1969:CUC

[Fra69] Evelyn Frank. Computer use in continued fraction expansions (in Technical Notes and Short Papers). Mathematics of Computation, 23(106):429–435, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Freudenstein:1961:BVR

[Fre61] Ferdinand Freudenstein. Bi-variate, rectangular, optimum-interval interpolation (in Technical Notes and Short Papers). Mathematics of Computation, 15(75):288–291, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Froese:1962:SBC

[Fro62] Charlotte Froese. On solving y′′ = fy+g with a boundary condition at infinity (in Technical Notes and Short Papers). Mathematics of Computation, 16(80): 492–494, October 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fields:1961:EHF

[FW61] Jerry L. Fields and Jet Wimp. Expansions of hypergeometric functions in hypergeometric functions. Mathematics of Computation, 15(76):390–395, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fyfe:1966:EER

[Fyf66] David J. Fyfe. Economical evaluation of Runge–Kutta formulae. Mathematics of Computation, 20(95):392–398, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

G:1968:BRD

[G.68] W. G. Book review: Donald Greenspan, editor, Numerical Solutions of Nonlinear Differential Equations, John Wiley & Sons, Inc., New York, 1966, x + 347 pp. Mathematics of Computation, 22(104): 896–898, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL https://doi.org/
10.2307/2004598.

Gray:1968:GT

[GA68] H. L. Gray and T. A. Atchison. The generalized G-transform. Mathematics of Computation, 22(103):595–605, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Galant:1969:GQR

[Gal69] David Galant. Gauss quadrature rules for the evaluation of 2π^{−1/2} ∫_{0}^{∞} exp(−x^{2})f(x)dx. Mathematics of Computation, 23(107): 674, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gary:1965:CEO

[Gar65a] John Gary. Computing eigenvalues of ordinary differential equations by finite differences. Mathematics of Computation, 19(91):365–379, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gary:1965:HMA

[Gar65b] John Gary. Hyman’s method applied to the general eigenvalue problem (in Technical Notes and Short Papers). Mathematics of Computation, 19(90):314– 316, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gary:1967:CRL

[Gar67] John Gary. On convergence rates for line overrelaxation (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):220–223, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gargantini:1969:CET

[Gar69] I. Gargantini. On the computation of the eigenvalues of a tridiagonal matrix (in Technical Notes and Short Papers). Mathematics of Computation, 23(106): 403–405, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1961:RCR

[Gau61] Walter Gautschi. Recursive computation of the repeated integrals of the error function. Mathematics of Computation, 15(75):227–232, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1966:CSD

[Gau66] Walter Gautschi. Computation of successive derivatives of f(z)/z. Mathematics of Computation, 20(94):209– 214, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1968:CGC

[Gau68]  Walter Gautschi.  Construction of

Gauss–Christoffel quadrature formulas. Mathematics of Computation, 22(102): 251–270, April 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1969:CMA

[Gau69] Walter Gautschi. On the condition of a matrix arising in the numerical inversion of the Laplace transform. Mathematics of Computation, 23(105):109– 118, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943–1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Galant:1968:HAG

[GB68] D. C. Galant and P. F. Byrd. High accuracy gamma function values for some rational arguments (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):885–887, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Good:1968:RHP

[GC68] I. J. Good and R. F. Churchhouse. The Riemann hypothesis and pseudorandom features of the Möbius sequence. Mathematics of Computation, 22(104):857– 861, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gear:1967:NIO

[Gea67] C. W. Gear. The numerical integration of ordinary differential equations. Mathematics of Computation, 21(98):

146–156, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gear:1969:SST

[Gea69] C. W. Gear. A simple set of test matrices for eigenvalue programs. Mathematics of Computation, 23(105):119–125, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gebhardt:1964:GND

[Geb64] Friedrich Gebhardt. Generating normally distributed random numbers by inverting the normal distribution function (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):302–306, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gebhardt:1967:GPR

[Geb67] Friedrich Gebhardt. Generating pseudorandom numbers by shuffling a Fibonacci sequence (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):708–709, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gerber:1964:FOH

[Ger64] Henry Gerber. First one hundred zeros of J_{0}(x) accurate to 19 significant figures (in Technical Notes and Short Papers). Mathematics of Computation, 18(86): 319–322, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gargantini:1967:CFA

[GH67] I. Gargantini and P. Henrici. A continued fraction algorithm for the computation of higher transcendental functions in the complex plane. Mathematics of Computation, 21(97):18–29, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gillies:1964:TNM

[Gil64] Donald B. Gillies. Three new Mersenne primes and a statistical theory. Mathematics of Computation, 18(85):93–97, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ginsberg:1963:CPO

[Gin63] F. Ginsberg. On the Cauchy problem for the one-dimensional heat equation. Mathematics of Computation, 17(83): 257–269, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1966:ESI

[Gla66] M. L. Glasser. Evaluation of some integrals involving the ψ-function (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):332– 333, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1968:SIA

[Gla68] M. L. Glasser. Some integrals of the arctangent function (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):445–447, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gloden:1960:NFT

[Glo60] A. Gloden. A note on factors of n^{4} + 1 (in Technical Notes and Short Papers). Mathematics of Computation, 14(71): 278–279, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gloden:1962:ACF

[Glo62] A. Gloden. Additions to Cunningham’s factor table of n^{4} + 1 (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):239–241, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gloden:1965:NPF

[Glo65] A Gloden. New primes of the form n^{4}+1 (in Technical Notes and Short Papers). Mathematics of Computation, 19(89): 144–145, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gardiner:1964:SDE

[GLS64] V. L. Gardiner, R. B. Lazarus, and P. R. Stein. Solutions of the Diophantine equation x^{3} + y^{3} = z^{3} − d. Mathematics of Computation, 18(87):408–413, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gourlay:1967:SOM

[GM67] A. R. Gourlay and A. R. Mitchell. Split operator methods for hyperbolic systems in p space variables. Mathematics of Computation, 21(99):351–354, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gourlay:1968:FDMa

[GM68a] A. R. Gourlay and J. Ll. Morris. Finitedifference methods for nonlinear hyperbolic systems. Mathematics of Computation, 22(101):28–39, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gourlay:1968:FDMb

[GM68b] A. R. Gourlay and J. Ll. Morris. Finitedifference methods for nonlinear hyperbolic systems. II. Mathematics of Computation, 22(103):549–556, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gourlay:1968:MFO

[GM68c] A. R. Gourlay and J. Ll. Morris. A multistep formulation of the optimized Lax–Wendroff method for nonlinear hyperbolic systems in two space variables. Mathematics of Computation, 22(104): 715–719, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Godart:1966:IMS

[God66] M. Godart. An iterative method for the solution of eigenvalue problems. Mathematics of Computation, 20(95):399– 406, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1960:RPP

[Gol60] Michael Goldberg. Rotors in polygons and polyhedra. Mathematics of Computation, 14(71):229–239, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Golub:1962:BET

[Gol62] Gene H. Golub. Bounds for eigenvalues of tridiagonal symmetric matrices computed by the LR method. Mathematics of Computation, 16(80):438–445, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Golub:1963:LBR

[Gol63] Gene H. Golub. On a lower bound for the rank of a partitioned square matrix. Mathematics of Computation, 17(82): 186–188, April 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goldberg:1969:SCE

[Gol69] Michael Goldberg. Stability configurations of electrons on a sphere. Mathematics of Computation, 23(108):785– 786, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goodman:1965:NSE

[Goo65] Theodore R. Goodman. The numerical solution of eigenvalue problems (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):462– 466, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1968:NMP

[Gor68] Paul Gordon. A note on a maximum principle for the DuFort–Frankel difference equation (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):437–439, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grace:1963:SLP

[Gra63] Donald W. Grace. Search for largest polyhedra. Mathematics of Computation, 17(82):197–199, April 1963. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gray:1964:CFD

[Gra64] John Gray. On certain finite difference schemes for hyperbolic systems. Mathematics of Computation, 18(85):1– 18, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greenberg:1960:ECP

[Gre60] Irwin Greenberg. The evaluation of certain probability integrals (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):376–378, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Greenberger:1961:PDS

[Gre61] Martin Greenberger. An a priori determination of serial correlation in computer generated random numbers. Mathematics of Computation, 15(76):383– 389, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See also corrigenda, Math. Comp. 16 (1962), p. 126.

Greenstadt:1967:REG

[Gre67] John Greenstadt. On the relative efficiencies of gradient methods. Mathematics of Computation, 21(99):360– 367, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gillis:1962:TPI

[GS62] J. Gillis and M. Shimshoni. Triple product integrals of Laguerre functions. Mathematics of Computation, 16(77): 50–62, January 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gray:1969:SLC

[GS69] H. L. Gray and W. R. Schucany. Some limiting cases of the G-transformation. Mathematics of Computation, 23(108): 849–859, October 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gray:1969:NAC

[GTM69] H. L. Gray, R. W. Thompson, and G. V.

McWilliams. A new approximation for the chi-square integral. Mathematics of Computation, 23(105):85–89, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gross:1968:DFV

[GV68] W. Gross and R. Vacca. Distribution of the figures 0 and 1 in the various orders of binary representations of kth powers of integers (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):423–427, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gillis:1960:PLP

[GW60] Joseph Gillis and George Weiss. Products of Laguerre polynomials. Mathematics of Computation, 14(69):60–63, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Golub:1969:CGQ

[GW69] Gene H. Golub and John H. Welsch. Calculation of Gauss quadrature rules. Mathematics of Computation, 23(106): 221–230, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). Loose microfiche suppl A1– A10.

H:1965:EPD

[H.65] A. S. H. Book review: Error Propagation for Difference Methods, by Peter Henrici. Mathematics of Computation, 19(89):152, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

                                              (electronic). URL https://www.jstor.
                                              org/stable/2004119.

Haber:1964:NSQ

[Hab64] Seymour Haber. A note on some quadrature formulas for the interval (−∞, ∞) (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):

313–314, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1966:MMC

[Hab66] Seymour Haber. A modified Monte-Carlo quadrature. Mathematics of Computation, 20(95):361–368, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1967:MQF

[Hab67a] Seymour Haber. Midpoint quadrature formulas (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):719–721, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1967:MMC

[Hab67b] Seymour Haber. A modified Monte-Carlo quadrature. II. Mathematics of Computation, 21(99):388–397, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1969:SQF

[Hab69] Seymour Haber. Stochastic quadrature formulas. Mathematics of Computation, 23(108):751–764, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hagis:1969:RPO

[Hag69] Peter Hagis, Jr. On relatively prime odd amicable numbers. Mathematics of Computation, 23(107):539–543, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hamming:1964:BRE

[Ham64] R. W. Hamming. Book review: Edward L. Stiefel, An Introduction to Numerical Mathematics, Academic Press, Inc., New York, 1963, x + 286 p., 24 cm. Mathematics of Computation, 18 (88):684, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL https://www.jstor.
org/stable/2002969.

Hamming:1965:ENA

[Ham65] R. W. Hamming. Book review: Elements of Numerical Analysis, by Peter Henrici. Mathematics of Computation, 19(91): 516–517, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.org/
stable/2003710.

Hansen:1968:SSE

[Han68] Eldon R. Hansen. On solving systems of equations using interval arithmetic. Mathematics of Computation, 22(102): 374–384, April 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harary:1962:PRP

[Har62a] Frank Harary. Permutations with restricted position (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):222–226, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harper:1962:QFI

[Har62b] W. M. Harper. Quadrature formulas for infinite integrals. Mathematics of Computation, 16(78):170–175, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hart:1966:CAR

[Har66] Roger G. Hart. A close approximation related to the error function (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):600–602, October 1966. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Haselgrove:1961:MNI

[Has61] C. B. Haselgrove. A method for numerical integration. Mathematics of Computation, 15(76):323–337, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hayashi:1965:CID

[Hay65] Harry S. Hayashi. Computer investigation of difference sets. Mathematics of Computation, 19(89):73–78, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heatley:1965:ETT

[Hea65] A. H. Heatley. An extension of the table of the Toronto function (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):118–123, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Herbold:1969:GCT

[Her69] Robert J. Herbold. A generalization of a class of test matrices. Mathematics of Computation, 23(108):823–826, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hirsch:1968:EOP

[Hir68] P. M. Hirsch. Evaluation of orthogonal polynomials and relationship to evaluating multiple integrals. Mathematics of Computation, 22(102):280–285, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hull:1964:ORK

[HJ64] T. E. Hull and R. L. Johnston. Optimum Runge–Kutta methods (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):306– 310, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hageman:1968:EOO

[HK68] L. A. Hageman and R. B. Kellogg. Estimating optimum overrelaxation parameters. Mathematics of Computation, 22 (101):60–68, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harumi:1960:VTN

[HKW60] Kasaburo Harumi, Shigetoshi Katsura, and John W. Wrench, Jr. Values of \frac{2}{π} ∫_{0}^{∞} (\frac{sin t}{t})^{n}dt (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):379, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hanson:1969:EAH

[HL69] Richard J. Hanson and Charles L. Lawson. Extensions and applications of the Householder algorithm for solving linear least squares problems. Mathematics of Computation, 23(108):787–812, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Haber:1969:TPS

[HO69] Seymour Haber and Charles F. Osgood. On a theorem of Piatetsky-Shapiro and approximation of multiple integrals (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):165–168, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hoffman:1965:NRM

[Hof65] A. J. Hoffman. On the nonsingularity of real matrices. Mathematics of Computation, 19(89):56–61, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Holdeman:1969:MAF

[Hol69] Jonas T. Holdeman, Jr. A method for the approximation of functions defined by formal series expansions in orthogonal polynomials. Mathematics of Computation, 23(106):275–287, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Householder:1960:BRV

[Hou60] A. S. Householder. Book review: L.

V. Kantorovich & V. I. Krylov, Approximate Methods of Higher Analysis, Translated from the third Russian edition by Curtis D. Benster, Interscience Publishers, Inc., New York, 1958, xv + 681 p., 24 cm. Price $17.00. Mathematics of Computation, 14(69):90–91, January 1960. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1960-14-069/S0025-5718-60-99238-
3;   http://www.ams.org/journals/
mcom/1960-14-069/S0025-5718-60-99238-
3/S0025-5718-60-99238-3.pdf.

Householder:1961:BRK

[Hou61] A. S. Householder. Book review: A.

Korganoff, with the collaboration of L. Bossett, J. L. Groboillot & J. Johnson, Méthodes de Calcul Numérique, Tome I: Algèbre non linéaire, Dunod, Paris, 1961, xxvii + 375 p., 25 cm. Price 58 NF. Mathematics of Computation, 15(75):312–313, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/
1961-15-075/S0025-5718-61-99215-
8;   http://www.ams.org/journals/

                                              mcom/1961-15-075/S0025-5718-61-99215-
                                              8/S0025-5718-61-99215-8.pdf.

Householder:1963:BRCa

[Hou63a] A. S. Householder. Book review: Centre Belge de Recherches Mathématiques, Colloque sur l’Analyse Numérique, Gauthier-Villars, Paris, 1961, 214 p., 25 cm. Mathematics of Computation, 17(82):209–210, April 1963. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

                                              http://www.ams.org/journals/mcom/
                                              1963-17-082/S0025-5718-63-99182-
                                              8;   http://www.ams.org/journals/
                                              mcom/1963-17-082/S0025-5718-63-99182-
                                              8/S0025-5718-63-99182-8.pdf.

Householder:1963:BRJ

[Hou63b] A. S. Householder. Book review: John Todd, Editor, A Survey of Numerical Analysis, McGraw-Hill Book Co., 1962, 23.5 cm., xvi + 589 p. Price $12.50. Mathematics of Computation, 17(81):89, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.ams.org/journals/
                                              mcom/1963-17-081/S0025-5718-1963-
                                              1781088-8;   http://www.ams.org/
                                              journals/mcom/1963-17-081/S0025-
                                              5718-1963-1781088-8/S0025-5718-1963-
                                              1781088-8.pdf.

Howell:1962:GPA

[How62] John R. Howell. Generation of permutations by addition (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):243–244, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hansen:1962:SRV

[HP62] Eldon R. Hansen and Merrell L. Patrick. Some relations and values for the generalized Riemann zeta function. Mathematics of Computation, 16(79):265– 274, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hubbard:1966:SLO

[Hub66] Bert Hubbard. Some locally onedimensional difference schemes for parabolic equations in an arbitrary region. Mathematics of Computation, 20(93):53–59, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hummer:1964:EDF

[Hum64] David G. Hummer. Expansions of Dawson’s function in a series of Chebyshev polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):317–319, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:1964:CCB

[Hun64] D. B. Hunter. The calculation of certain Bessel functions (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):123–128, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hunter:1968:ECF

[Hun68] D. B. Hunter. The evaluation of a class of functions defined by an integral (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):440– 444, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hurwitz:1962:NMP

[Hur62] Alexander Hurwitz. New Mersenne primes (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):249–251, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hammer:1960:QFI

[HW60] Preston C. Hammer and Howard H. Wicke. Quadrature formulas involving derivatives of the integrand. Mathematics of Computation, 14(69):3–7, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Habetler:1961:SSO

[HW61] G. J. Habetler and E. L. Wachspress. Symmetric successive overrelaxation in solving diffusion difference equations. Mathematics of Computation, 15(76): 356–362, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isemanger:1960:CF

[Ise60] K. R. Isemanger. The complete factorization of 2^{132} +1. Mathematics of Computation, 14(69):73–74, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isemonger:1961:CFT

[Ise61] K. R. Isemonger. Complete factorization of 2^{159} −1 (in Technical Notes and Short Papers). Mathematics of Computation, 15(75):295–296, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Isemonger:1965:SAF

[Ise65] K. R. Isemonger. Some additional factorizations of 2^{n} ± 1 (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):145–146, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jagerman:1964:AJD

[Jag64] David L. Jagerman. The autocorrelation and joint distribution functions

of the sequences \{\frac{a}{m}j^{2}\}, \{\frac{a}{m}(j + τ)^{2}\}.

Mathematics of Computation, 18(86): 211–232, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jagerman:1965:STC

[Jag65] D. L. Jagerman. Some theorems concerning pseudo-random numbers. Mathematics of Computation, 19(91):418– 426, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jagerman:1966:IMM

[Jag66] D. Jagerman. Investigation of a modified mid-point quadrature formula. Mathematics of Computation, 20(93):79–89, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jamet:1968:FDM

[Jam68a] Pierre Jamet. Finite difference methods for the computation of the “Poisson kernel” of elliptic operators. Mathematics of Computation, 22(103):477–488, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jamet:1968:NME

[Jam68b] Pierre Jamet. Numerical methods and existence theorems for parabolic differential equations whose coefficients are singular on the boundary. Mathematics of Computation, 22(104):721–743, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jarratt:1966:SFO

[Jar66] P. Jarratt. Some fourth order multipoint iterative methods for solving equations (in Technical Notes and Short Papers). Mathematics of Computation, 20(95): 434–437, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

James:1969:CIC

[JC69] Rodney James and John Cannon. Computation of isomorphism classes of pgroups. Mathematics of Computation, 23 (105):135–140, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jennrich:1962:MDC

[Jen62] R. I. Jennrich. Missing data correlation computations (in Technical Notes and Short Papers). Mathematics of Computation, 16(80):496–497, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

James:1965:GSR

[JJ65] Wendy James and P. Jarratt. The generation of square roots on a computer with rapid multiplication compared with division (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):497–500, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jones:1967:SCL

[JLB67] M. F. Jones, M. Lal, and W. J. Blundon. Statistics on certain large primes (in technical notes and short notices). Mathematics of Computation, 21(97): 103–107, January 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1963:GPA

[Joh63] Selmer M. Johnson. Generation of permutations by adjacent transposition (in Technical Notes and Short Papers). Mathematics of Computation, 17(83): 282–285, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1965:ERM

[Joh65] Selmer M. Johnson. An elementary remark on maximal gaps between successive primes (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):675–676, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Johnson:1968:UAF

[Joh68] Lee Johnson. Unicity in approximation of a function and its derivatives (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):873– 875, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jones:1961:NSC

[Jon61] J. G. Jones. On the numerical solution of convolution integral equations and systems of such equations. Mathematics of Computation, 15(74):131–142, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jordan:1968:EEG

[Jor68] T. L. Jordan. Experiments on error growth associated with some linear least-squares procedures. Mathematics of Computation, 22(103):579–588, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kahng:1969:OI

[Kah69] S. W. Kahng. Osculatory interpolation. Mathematics of Computation, 23(107): 621–629, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kammerer:1961:PAF

[Kam61] William J. Kammerer. Polynomial approximations to finitely oscillating functions. Mathematics of Computation, 15(74):115–119, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kammler:1969:NSD

[Kam69] David W. Kammler. Numerical solution of the Dirichlet problem for systems of circular conductors between parallel ground lines. Mathematics of Computation, 23(105):29–36, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaniel:1966:ESC

[Kan66] Shmuel Kaniel. Estimates for some computational techniques in linear algebra. Mathematics of Computation, 20(95): 369–378, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karst:1961:NFM

[Kar61] Edgar Karst. New factors of Mersenne numbers (in Technical Notes and Short Papers). Mathematics of Computation, 15(73):51, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Katsanis:1968:NSS

[Kat68] Theodore Katsanis. Numerical solution of symmetric positive differential equations. Mathematics of Computation, 22 (104):763–783, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kravitz:1964:LTM

[KB64] Sidney Kravitz and Murray Berg. Lucas’ test for Mersenne numbers, 6000 < p < 7000 (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):148–149, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1967:CTE

[KB67] Donald E. Knuth and Thomas J. Buckholtz. Computation of tangent, Euler, and Bernoulli numbers. Mathematics of Computation, 21(100):663– 688, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/
stable/2005010.

Keedwell:1965:SPP

[Kee65] A. D. Keedwell. A search for projective planes of a special type with the aid of a digital computer (in Technical Notes and Short Papers). Mathematics of Computation, 19(90):317–322, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kellogg:1964:DEM

[Kel64] R. B. Kellogg. Difference equations on a mesh arising from a general triangulation. Mathematics of Computation, 18(86):203–210, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kellogg:1969:NAD

[Kel69] R. B. Kellogg. A nonlinear alternating direction method. Mathematics of Computation, 23(105):23–27, January 1969.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kershaw:1969:EIT

[Ker69] D. Kershaw. The explicit inverses of two commonly occurring matrices (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):189– 191, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

KampedeFeriet:1962:CSN

[KF62] J. Kampe de Feriet and François N. Frenkiel. Correlations and spectra for non-stationary random functions. Mathematics of Computation, 16(77):1–21, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kim:1968:EKA

[KG68] Chul-Hee Kim and R. E. Gilbert. Evaluation of a kernel associated with laminar flow tubular catalytic reactors. Mathematics of Computation, 22(103):517– 523, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Killgrove:1960:NNC

[Kil60] Raymond B. Killgrove. A note on the nonexistence of certain projective planes of order nine. Mathematics of Computation, 14(69):70–71, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1966:RKM

[Kin66] Richard King. Runge–Kutta methods with constrained minimum error bounds. Mathematics of Computation, 20(95):386–391, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kirkpatrick:1960:TVM

[Kir60] E. T. Kirkpatrick. Tables of values of the modified Mathieu functions. Mathematics of Computation, 14(70):118– 129, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karst:1964:FPC

[KK64] E. Karst and U. Karst. The first power of 2 with 8 consecutive zeros (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):508, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kilpatrick:1967:CIP

[KKI67] J. E. Kilpatrick, Shigetoshi Katsura, and Yuji Inoue. Calculations of integrals of products of Bessel functions. Mathematics of Computation, 21(99): 407–412, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kreiss:1968:DAW

[KL68] Heinz-Otto Kreiss and Einar Lundqvist.

On difference approximations with wrong boundary values. Mathematics of Computation, 22(101):1–12, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Klopfenstein:1964:CLS

[Klo64] R. W. Klopfenstein. Conditional least squares polynomial approximation (in Technical Notes and Short Papers). Mathematics of Computation, 18(88): 659–662, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Konheim:1967:NEW

[KM67] Alan G. Konheim and Willard L. Miranker. Numerical evaluation of Wiener integrals. Mathematics of Computation, 21(97):49–65, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Klopfenstein:1968:NSO

[KM68] R. W. Klopfenstein and R. S. Millman. Numerical stability of a oneevaluation predictor-corrector algorithm for numerical solution of ordinary differential equations. Mathematics of Computation, 22(103):557–564, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1962:ECP

[Knu62] Donald E. Knuth. Euler’s constant to 1271 places. Mathematics of Computation, 16(79):275–281, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Konheim:1965:MDT

[Kon65] Alan G. Konheim. Mantissa distributions (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):143–144, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kelisky:1968:RAL

[KR68] R. P. Kelisky and T. J. Rivlin. A rational approximation to the logarithm. Mathematics of Computation, 22(101): 128–136, January 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kravitz:1960:CTN

[Kra60] Sidney Kravitz. The congruence 2^{p−1} ≡ 1(mod p^{2})forp < 100, 000 (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):378, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kravitz:1961:DMN

[Kra61] Sidney Kravitz. Divisors of Mersenne numbers 10, 000 < p < 15, 000 (in Technical Notes and Short Papers). Mathematics of Computation, 15(75):292– 293, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kravitz:1966:DMD

[Kra66]  Sidney Kravitz.  Distributions of

Mersenne divisors (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):448–449, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kreiss:1968:STD

[Kre68] Heinz-Otto Kreiss. Stability theory for difference approximations of mixed initial boundary value problems. I. Mathematics of Computation, 22(104):703– 714, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Krogh:1967:NEC

[Kro67] Fred T. Krogh. A note on the effect of conditionally stable correctors (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):717– 719, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Keller:1963:ELP

[KS63] H. B. Keller and J. R. Swenson. Experiments on the lattice problem of Gauss. Mathematics of Computation, 17(83): 223–230, July 1963. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Kellogg:1965:OAD

[KS65] R. B. Kellogg and J. Spanier. On optimal alternating direction parameters for singular matrices (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):448–452, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LaBudde:1963:RAR

[La 63] C. Donald La Budde. The reduction of an arbitrary real square matrix to tridiagonal form using similarity transformations (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):433–437, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

LaBudde:1964:NAD

[La 64] C. Donald La Budde. A new algorithm for diagonalizing a real symmetric matrix (in Technical Notes and Short Papers). Mathematics of Computation, 18 (85):118–123, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1967:PFT

[Lal67] M. Lal. Primes of the form n^{4} + 1 (in Technical Notes and Short Papers). Mathematics of Computation, 21(98): 245–247, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1969:INT

[Lal69] Mohan Lal. Iterates of a numbertheoretic function (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):181–183, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lang:1960:ECC

[Lan60] H. A. Lang. On the evaluation of certain complex elliptic integrals (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):195–199, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Langel:1967:ACD

[Lan67] R. Langel. An algorithm for certain double sums of polynomial series (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):230– 232, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Larkin:1964:SSE

[Lar64] Bert K. Larkin. Some stable explicit difference approximations to the diffusion equation. Mathematics of Computation, 18(86):196–202, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lavoie:1964:IFS

[Lav64a] Jean L. Lavoie. On inverses of finite segments of the generalized Hilbert matrix (in Technical Notes and Short Papers). Mathematics of Computation, 18(85): 141–143, January 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lavoie:1964:ECD

[Lav64b] Jean L. Lavoie. On the evaluation of certain determinants (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):653–659, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Layman:1964:FDE

[Lay64] J. W. Layman. A finite difference exponential approximation method (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):113–118, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Layman:1967:MAE

[Lay67]  J. W. Layman.  Mixed algebraic-

exponential interpolation using finite differences (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):498–500, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lazarkiewicz:1961:SEC

[LB61] A. Lazarkiewicz and W. Balasinski. A simple experimental computer with negative basis. Mathematics of Computation, 15(75):275–285, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1966:SDE

[LB66] M. Lal and W. J. Blundon. Solutions of the Diophantine equations x^{2} + y^{2} = l^{2}, y^{2} + z^{2} = m^{2}, z^{2} + x^{2} = n^{2} (in Technical Notes and Short Papers). Mathematics of Computation, 20(93): 144–147, January 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1961:EHF

[LC61] Yudell L. Luke and Richard L. Coleman. Expansion of hypergeometric functions in series of other hypergeometric functions. Mathematics of Computation, 15(75):233–237, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levine:1962:MPM

[LD62] Jack Levine and R. E. Dalton. Minimum periods, modulo p, of first-order Bell exponential integers. Mathematics of Computation, 16(80):416–423, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1967:NCI

[LD67] J. N. Lyness and L. M. Delves. On numerical contour integration round a closed contour. Mathematics of Computation, 21(100):561–577, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1968:SDE

[LD68] Mohan Lal and James Dawe. Solutions of the Diophantine equation x^{2}−dy^{4} = k (in Technical Notes and Short Papers). Mathematics of Computation, 22(103): 679–682, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leavitt:1966:MAP

[Lea66] Jay A. Leavitt. Methods and applications of power series. Mathematics of Computation, 20(93):46–52, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leavitt:1967:PAC

[Lea67] Jay A. Leavitt. A polynomial approximation converging in a lens-shaped region (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):491–493, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ledsham:1961:NIU

[Led61] F. C. Ledsham. Numerical integration using sums of exponential functions (in Technical Notes and Short Papers). Mathematics of Computation, 15(73): 48–51, January 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lees:1962:NCA

[Lee62] Milton Lees. A note on the convergence of alternating direction methods. Mathematics of Computation, 16(77): 70–75, January 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lees:1966:LTL

[Lee66] Milton Lees. A linear three-level difference scheme for quasilinear parabolic equations. Mathematics of Computation, 20(96):516–522, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lee:1968:ANB

[Lee68] Elvin J. Lee. Amicable numbers and the bilinear Diophantine equation. Mathematics of Computation, 22(101):181– 187, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lee:1969:DNS

[Lee69] Elvin Lee. On divisibility by nine of the sums of even amicable pairs. Mathematics of Computation, 23(107):545– 548, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1960:LF

[Leh60] R. Sherman Lehman. On Liouville’s function. Mathematics of Computation, 14(72):311–320, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehman:1966:SZR

[Leh66] R. Sherman Lehman. Separation of zeros of the Riemann zeta-function. Mathematics of Computation, 20(96):523–541, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Leszczynski:1965:SST

[Les65] Bogumil Leszczynski. On the sum of the series ∑_{ν=0}^{∞}(t^{ν}/[uν + m]!) (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):677–679, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lether:1966:URE

[Let66] F. G. Lether. The use of Richardson extrapolation in one-step methods with variable step size. Mathematics of Computation, 20(95):379–385, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Levine:1961:SAH

[Lev61] Jack Levine. Some applications of highspeed computers to the case n = 2 of algebraic cryptography. Mathematics of Computation, 15(75):254–260, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1969:FII

[Lew69] Richard L. Lewis. On finite integrals involving trigonometric, Bessel, and Legendre functions. Mathematics of Computation, 23(106):259–273, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Law:1968:RBE

[LF68] V. J. Law and R. H. Fariss. On the relations between eigenvalues and eigenvectors for matrices resulting from pre and post multiplication by the transpose (in Technical Notes and Short Papers). Mathematics of Computation, 22(104): 867–868, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1965:CTT

[LFiCM65] Yudell L. Luke, Wyman Fair, Gerald ine Coombs, and Rosemary Moran. On a constant in the theory of trigonometric series (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):501–502, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1960:EHP

[Lin60] Chih-Bing Ling. Evaluation at half periods of Weierstrass’ elliptic function with rectangular primitive periodparallelogram. Mathematics of Computation, 14(69):67–70, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1965:EHP

[Lin65a] Chih-Bing Ling. Evaluation at half periods of Weierstrass’ elliptic functions with double periods 1 and e^{iα} (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):658–661, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1965:TVR

[Lin65b] Chih-Bing Ling. Tables of values of σ_{2s} relating to Weierstrass’ elliptic function (in Technical Notes and Short Papers). Mathematics of Computation, 19(89): 123–127, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Linz:1969:MSN

[Lin69] Peter Linz. A method for solving nonlinear Volterra integral equations of the second kind. Mathematics of Computation, 23(107):595–599, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1966:NSD

[LJB66] M. Lal, M. F. Jones, and W. J. Blundon.

Numerical solutions of the Diophantine equation y^{3} − x^{2} = k (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):322–325, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laslett:1962:EZC

[LL62a] L. Jackson Laslett and William Lewish.

Evaluation of the zeros of cross-product Bessel functions (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):226–232, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lin:1962:EJE

[LL62b] Y. K. Lin and F. A. Lee. Expansions of Jacobian elliptic functions in powers of the modulus (in Technical Notes and Short Papers). Mathematics of Computation, 16(79):372–375, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1962:MPT

[LLMS62] D. H. Lehmer, E. Lehmer, W. H. Mills, and J. L. Selfridge. Machine proof of a theorem on cubic residues. Mathematics of Computation, 16(80):407–415, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lo:1965:DME

[LLS65] Y. T. Lo, S. W. Lee, and B. Sun. On Davis’ method of estimating quadrature errors (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):133–138, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1967:NQA

[LN67] J. N. Lyness and B. W. Ninham. Numerical quadrature and asymptotic expansions. Mathematics of Computation, 21(98):162–178, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lovass-Nagy:1969:RFS

[LNP69] Victor Lovass-Nagy and David L. Powers. Reduction of functions of some partitioned matrices. Mathematics of Computation, 23(105):127–133, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Longman:1960:MNE

[Lon60a] I. M. Longman. A method for the numerical evaluation of finite integrals of oscillatory functions. Mathematics of Computation, 14(69):53–59, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Longman:1960:UNM

[Lon60b] I. M. Longman. On the utility of Newton’s method for computing complex roots of equations (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):187–189, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lord:1960:IPL

[Lor60] R. D. Lord. Integrals of products of Laguerre polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):375–376, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lotto:1962:GAP

[Lot62] Gary Lotto. On the generation of all possible stepwise combinations (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):241– 243, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowan:1960:NTH

[Low60a] Arnold N. Lowan. On the numerical treatment of heat conduction problems with mixed boundary conditions (in Technical Notes and Short Papers). Mathematics of Computation, 14(71): 266–270, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowan:1960:PEI

[Low60b] Arnold N. Lowan. On the propagation of errors in the inversion of certain tridiagonal matrices. Mathematics of Computation, 14(72):333–338, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowan:1960:PREa

[Low60c] Arnold N. Lowan. On the propagation of round-off errors in the numerical integration of the heat equation. Mathematics of Computation, 14(70):139–146, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowan:1960:PREb

[Low60d] Arnold N. Lowan. On the propagation of round-off errors in the numerical treatment of the wave equation. Mathematics of Computation, 14(71):223–228, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lowan:1961:SCE

[Low61] Arnold N. Lowan. On stability criteria of explicit difference schemes for certain heat conduction problems with uncommon boundary conditions. Mathematics of Computation, 15(74):179–185, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Low:1966:FPZ

[Low66] R. D. Low. On the first positive zero of P_{ν−1/2}^{(−m)}(cos θ), considered as a function of ν (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):421–424, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1966:ESB

[LP66] L. J. Lander and T. R. Parkin. Equal sums of biquadrates (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):450–451, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1967:CPA

[LP67a] L. J. Lander and T. R. Parkin. Consecutive primes in arithmetic progression (in Technical Notes and Short Papers). Mathematics of Computation, 21 (99):489, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1967:CES

[LP67b] L. J. Lander and T. R. Parkin. A counterexample to Euler’s sum of powers conjecture (in technical notes and short notices). Mathematics of Computation, 21(97):101–103, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1967:FAP

[LP67c] L. J. Lander and T. R. Parkin. On first appearance of prime differences (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):483– 488, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lander:1967:SES

[LPS67] L. J. Lander, T. R. Parkin, and J. L.

Selfridge. A survey of equal sums of like powers. Mathematics of Computation, 21(99):446–459, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lynch:1968:CRA

[LR68] Robert E. Lynch and John R. Rice.

Convergence rates of ADI methods with smooth initial error. Mathematics of Computation, 22(102):311–335, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lal:1969:NSF

[LRB69] M. Lal, W. Russell, and W. J. Blundon. A note on sums of four cubes (in Technical Notes and Short Papers). Mathematics of Computation, 23(106): 423–424, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambert:1965:NSC

[LS65a] John D. Lambert and Brian Shaw. On the numerical solution of y′ = f(x, y) by a class of formulae based on rational approximation (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):456–462, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lee:1965:AIF

[LS65b] Nancy Lee and A. H. Stroud. Approximate integration formulas for ellipses (in Technical Notes and Short Papers). Mathematics of Computation, 19(90): 322–327, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambert:1966:MNS

[LS66] J. D. Lambert and B. Shaw. A method for the numerical solution of y′ = f(x, y) based on a self-adjusting nonpolynomial interpolant. Mathematics of Computation, 20(93):11–20, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lear:1967:IRM

[LS67a] James D. Lear and James E. Sturm. An integral representation for the modified Bessel function of the third kind, computable for large, imaginary order (in Technical Notes and Short Papers). Mathematics of Computation, 21(99): 496–498, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1967:CFV

[LS67b] H. R. Lewis, Jr. and E. F. Stovall, Jr. Comments on a floating-point version of Nordsieck’s scheme for the numerical integration of differential equations.

Mathematics of Computation, 21(98): 157–161, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1967:CFP

[LS67c] H. R. Lewis, Jr. and E. J. Stovall, Jr.

Comments on a floating-point version of Nordsieck’s scheme for the numerical integration of differential equations. Mathematics of Computation, 21(98): 157–161, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lietzke:1964:CSM

[LSL64] M. H. Lietzke, R. W. Stoughton, and Marjorie P. Lietzke. A comparison of several methods for inverting large symmetric positive definite matrices. Mathematics of Computation, 18(87):449– 456, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1964:EHP

[LT64] Chih-Bing Ling and Chen-Peng Tsai.

Evaluation at half periods of Weierstrass’ elliptic function with rhombic primitive period-parallelogram. Mathematics of Computation, 18(87):433– 440, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1968:AEI

[Luk68] Yudell L. Luke. Approximations for elliptic integrals. Mathematics of Computation, 22(103):627–634, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lunnon:1968:MFP

[Lun68] W. F. Lunnon. A map-folding problem (in Technical Notes and Short Papers).

Mathematics of Computation, 22(101): 193–199, January 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luther:1968:ESO

[Lut68] H. A. Luther. An explicit sixthorder Runge–Kutta formula (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):434– 436, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1963:JPE

[LW63] Y. L. Luke and J. Wimp. Jacobi polynomial expansions of a generalized hypergeometric function over a semi-infinite ray. Mathematics of Computation, 17 (84):395–404, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1965:TVT

[LW65] Chih-Bing Ling and Hsien-Chueh Wu. Tables of values of three infinite integrals (in Technical Notes and Short Papers). Mathematics of Computation, 19(91): 487–494, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1972:CIM

[LWF72] Yudell L. Luke, Jet Wimp, and Wyman Fair, editors. Cumulative Index to Mathematics of Computation, Volumes 1–24 (1943–1969). American Mathematical Society, Providence, RI, USA, 1972. xvii + 461 pp.

Lynn:1964:REM

[Lyn64] M. Stuart Lynn. On the roundoff error in the method of successive over-relaxation. Mathematics of Computation, 18(85):36–49, January 1964.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1965:IRH

[Lyn65a] J. N. Lyness. Integration rules of hypercubic symmetry over a certain spherically symmetric space (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):471–476, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1965:LNF

[Lyn65b] J. N. Lyness. Limits on the number of function evaluations required by certain high-dimensional integration rules of hypercubic symmetry. Mathematics of Computation, 19(92):638–643, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1965:SIRa

[Lyn65c] J. N. Lyness. Symmetric integration rules for hypercubes I. error coefficients. Mathematics of Computation, 19(90): 260–276, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1965:SIRb

[Lyn65d] J. N. Lyness. Symmetric integration rules for hypercubes II. rule projection and rule extension. Mathematics of Computation, 19(91):394–407, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1965:SIRc

[Lyn65e] J. N. Lyness. Symmetric integration rules for hypercubes III. construction of integration rules using null rules. Mathematics of Computation, 19(92):625–637, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1968:DFA

[Lyn68] J. N. Lyness. Differentiation formulas for analytic functions. Mathematics of Computation, 22(102):352–362, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lyness:1969:QMB

[Lyn69] J. N. Lyness. Quadrature methods based on complex function values. Mathematics of Computation, 23(107):601– 619, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Metropolis:1965:RCU

[MA65] N. Metropolis and R. L. Ashenhurst. Radix conversion in an unnormalized arithmetic system. Mathematics of Computation, 19(91):435–441, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mack:1965:BFI

[Mac65] Lawrence R. Mack. Bessel-function identities needed for the theory of axisymmetric gravity waves (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):654–657, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Magula:1960:CFM

[Mag60] D. A. Magula. A conjugate factor method for the solution of a cubic (in Technical Notes and Short Papers). Mathematics of Computation, 14(71): 281–283, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mann:1961:MCT

[Man61] Henry B. Mann. On modular computation (in Technical Notes and Short Papers). Mathematics of Computation, 15(74):190–192, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mangad:1966:BTD

[Man66] Moshe Mangad. Bounds for the twodimensional discrete harmonic Green’s function. Mathematics of Computation, 20(93):60–67, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mangad:1967:SLV

[Man67] Moshe Mangad. Some limiting values and two error estimation procedures for power series approximations. Mathematics of Computation, 21(99): 423–430, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mapes:1963:FMC

[Map63] David C. Mapes. Fast method for computing the number of primes less than a given limit. Mathematics of Computation, 17(82):179–185, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Markham:1968:IPC

[Mar68] T. L. Markham. An iterative procedure for computing the maximal root of a positive matrix (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):869–871, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mayoh:1965:GTI

[May65] B. H. Mayoh. A graph technique for inverting certain matrices (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):644–646, October 1965. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

McAndrew:1963:PPT

[McA63] M. H. McAndrew. A partition problem (in Technical Notes and Short Papers). Mathematics of Computation, 17(83): 291–295, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McAndrew:1965:ASP

[McA65] M. H. McAndrew. An algorithm for solving a polynomic congruence, and its application to error-correcting cods. Mathematics of Computation, 19(89): 68–72, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McEliece:1969:FPF

[McE69] Robert J. McEliece. Factorization of polynomials over finite fields. Mathematics of Computation, 23(108):861– 867, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McLaren:1963:ONI

[McL63] A. D. McLaren. Optimal numerical integration on a sphere. Mathematics of Computation, 17(84):361–383, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McNamee:1964:EBE

[McN64] John McNamee. Error-bounds for the evaluation of integrals by the Euler– Maclaurin formula and by Gauss-type formulae. Mathematics of Computation, 18(87):368–381, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mechel:1966:CMB

[Mec66] Fr. Mechel. Calculation of the modified Bessel functions of the second kind with complex argument. Mathematics of Computation, 20(95):407–412, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mechel:1968:IRT

[Mec68] Fr. Mechel. Improvement in recurrence techniques for the computation of Bessel functions of integral order (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):202– 205, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meicler:1969:SAM

[Mei69] Marcel Meicler. A steepest ascent method for the Chebyshev problem. Mathematics of Computation, 23(108): 813–817, October 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Melzak:1968:NEI

[Mel68] Z. A. Melzak. Numerical evaluation of an isoperimetric constant. Mathematics of Computation, 22(101):188–190, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Melzak:1969:SPC

[Mel69] Z. A. Melzak. On the solid-packing constant for circles (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):169–172, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Metsankyla:1969:CFF

[Met69] Tauno Metsankyla. Calculation of the first factor of the class number of the cyclotomic field. Mathematics of Computation, 23(107):533–537, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Michels:1963:AWC

[Mic63] H. H. Michels. Abscissas and weight coefficients for Lobatto quadrature. Mathematics of Computation, 17(83):237– 244, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1960:NQRb

[Mil60a] J. C. P. Miller. Numerical quadrature over a rectangular domain in two or more dimensions. Mathematics of Computation, 14(70):130–138, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1960:NQRa

[Mil60b] J. C. P. Miller. Numerical quadrature over a rectangular domain in two or more dimensions: Part 1. quadrature over a square, using up to sixteen equally spaced points. Mathematics of Computation, 14(69):13–20, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1960:NQRc

[Mil60c] J. C. P. Miller. Numerical quadrature over a rectangular domain in two or more dimensions. part 3: Quadrature of a harmonic integrand. Mathematics of Computation, 14(71):240–248, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1963:FIP

[Mil63] James Miller. Formulas for integrals of products of associated Legendre or Laguerre functions (in Technical Notes and Short Papers). Mathematics of Computation, 17(81):84–87, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1968:RLO

[Mil68a] John Miller. On the resolvent of a linear operator associated with a wellposed Cauchy problem. Mathematics of Computation, 22(103):541–548, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Milnes:1968:NCP

[Mil68b] Harold Willis Milnes. A note concerning the properties of a certain class of test matrices. Mathematics of Computation, 22(104):827–832, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mitchell:1961:SMB

[Mit61] A. R. Mitchell. Solution of the von Mises boundary layer equation using a high-speed computer. Mathematics of Computation, 15(75):238–242, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mitchell:1966:NLP

[Mit66] W. C. Mitchell. The number of lattice points in a k-dimensional hypersphere. Mathematics of Computation, 20(94): 300–310, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mitchell:1968:EGC

[Mit68] W. C. Mitchell. An evaluation of Golomb’s constant. Mathematics of Computation, 22(102):411–415, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miranker:1967:PMN

[ML67] Willard L. Miranker and Werner Liniger. Parallel methods for the numerical integration of ordinary differential equations. Mathematics of Computation, 21(99):303–320, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moon:1966:MRP

[MM66] J. W. Moon and L. Moser. A matrix reduction problem (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):328–330, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McQueary:1967:BFI

[MM67] Charles E. McQueary and Lawrence R. Mack. Bessel-function integrals needed for two classes of physical problems. Mathematics of Computation, 21(99): 413–417, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mond:1964:MPA

[Mon64] B. Mond. Multivariate polynomial approximation for equidistant data (in Technical Notes and Short Papers). Mathematics of Computation, 18(86): 298–301, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moore:1966:DAS

[Moo66] D. H. Moore. A division algebra for sequences defined on all the integers. Mathematics of Computation, 20(94): 311–317, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moody:1967:ADF

[Moo67] William T. Moody. Approximations for the Psi (digamma) function (in technical notes and short notices). Mathematics of Computation, 21(97):112, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moursund:1964:CAF

[Mou64] D. G. Moursund. Chebyshev approximations of a function and its derivatives. Mathematics of Computation, 18(87): 382–389, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mitchell:1963:EDM

[MP63] A. R. Mitchell and R. P. Pearce. Explicit difference methods for solving the cylindrical heat conduction equation. Mathematics of Computation, 17(84):426–432, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Medhurst:1965:EI

[MR65] R. G. Medhurst and J. H. Roberts.

Evaluation of the integral I_{n}(b) =

\frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos(bx)dx. Mathematics

of Computation, 19(89):113–117, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muroga:1962:MDF

[MTK62] S. Muroga, I. Toda, and M. Kondo. Majority decision functions of up to six variables. Mathematics of Computation, 16 (80):459–472, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muller:1959:NMG

[Mul59] Mervin E. Muller. A note on a method for generating points uniformly on Ndimensional spheres. Communications of the ACM, 2(4):19–20, April 1959. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Murphy:1967:NAB

[Mur67] W. D. Murphy. Numerical analysis of boundary-layer problems in ordinary differential equations. Mathematics of Computation, 21(100):583–596, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mustard:1964:NID

[Mus64] D. Mustard. Numerical integration over the n-dimensional spherical shell. Mathematics of Computation, 18(88):578– 589, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muskat:1966:DOP

[Mus66] Joseph B. Muskat. On divisors of odd perfect numbers (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):141–144, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nashed:1965:GIM

[Nas65] M. Z. Nashed. On general iterative methods for the solutions of a class of nonlinear operator equations. Mathematics of Computation, 19(89): 14–24, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Navot:1963:EMF

[Nav63] Israel Navot. The Euler–Maclaurin functional for functions with a quasi-step discontinuity. Mathematics of Computation, 17(84):337–345, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nellis:1966:REE

[NC66] W. J. Nellis and B. C. Carlson. Reduction and evaluation of elliptic integrals. Mathematics of Computation, 20(94): 223–231, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ng:1969:CPE

[NDT69] Edward W. Ng, C. J. Devine, and R. F.

Tooper. Chebyshev polynomial expansion of Bose–Einstein functions of orders 1 to 10. Mathematics of Computation, 23(107):639–643, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nelson:1961:NTH

[Nel61] C. W. Nelson. New tables of Howland’s and related integrals. Mathematics of Computation, 15(73):12–18, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1963:MIF

[New63] A. C. R. Newbery. Multistep integration formulas (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):452–455, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1966:IAT

[New66] A. C. R. Newbery. Interpolation by algebraic and trigonometric polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 20(96): 597–599, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1967:CSS

[New67a] A. C. R. Newbery. Convergence of successive substitution starting procedures (in Technical Notes and Short Papers). Mathematics of Computation, 21(99): 489–491, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1967:NPM

[New67b] Morris Newman. Note on partitions modulo 5 (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):481–482, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newbery:1969:SEL

[New69] A. C. R. Newbery. Some extensions of Legendre quadrature (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):173–176, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1963:ICS

[NF63] J. N. Newman and W. Frank. An integral containing the square of a Bessel function. Mathematics of Computation, 17(81):64–70, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nikolai:1960:PIM

[Nik60] Paul J. Nikolai. Permanents of incidence matrices (in Technical Notes and Short Papers). Mathematics of Computation, 14(71):262–266, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ninham:1969:FAE

[NL69] B. W. Ninham and J. N. Lyness. Further asymptotic expansions for the error functional. Mathematics of Computation, 23(105):71–83, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nordsieck:1962:NIO

[Nor62] Arnold Nordsieck. On numerical integration of ordinary differential equations. Mathematics of Computation, 16(77): 22–49, January 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Otis:1964:SHE

[OB64] A. B. Otis, Jr. and M. P. Barnett. Surface harmonic expansions of products of Cartesian coordinates. Mathematics of Computation, 18(88):635–643, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ormonde:1965:MCI

[OD65] Stephan Ormonde and Jerald S. Dick.

A method of calculation of integrals of products of radial atomic wave functions (in Technical Notes and Short Papers). Mathematics of Computation, 19(92): 647–650, October 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oldham:1968:AFT

[Old68] K. B. Oldham. Approximations for the x exp x^{2} erfc x function (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):454, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olunloyo:1965:STD

[Olu65] V. O. S. Olunloyo. A special technique for the determination of eigenvalues (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):466–471, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1964:EAM

[Olv64] F. W. J. Olver. Error analysis of Miller’s recurrence algorithm. Mathematics of Computation, 18(85):65–74, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osborn:1968:IBF

[OM68] David Osborn and Richard Madey. The incomplete beta function and its ratio to the complete beta function. Mathematics of Computation, 22(101):159– 162, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osborne:1962:NFD

[Osb62] M. R. Osborne. A note on finite difference methods for solving the eigenvalue problems of second-order differential equations. Mathematics of Computation, 16(79):338–346, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osborne:1967:MTE

[Osb67] M. R. Osborne. Minimising truncation error in finite difference approximations to ordinary differential equations. Mathematics of Computation, 21(98): 133–145, April 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1969:SDE

[Osh69a] Stanley Osher. On systems of difference equations with wrong boundary conditions. Mathematics of Computation, 23(107):567–572, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Osher:1969:SDA

[Osh69b] Stanley Osher. Stability of difference approximations of dissipative type for mixed initial-boundary value problems. Mathematics of Computation, 23(106): 335–340, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ostrowski:1961:TML

[OV61] R. T. Ostrowski and K. D. Van Duren.

On a theorem of Mann on Latin squares (in Technical Notes and Short Papers). Mathematics of Computation, 15(75): 293–295, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1964:NBA

[Par64a] Beresford Parlett. A note on La Budde’s algorithm (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):505–506, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1964:LMA

[Par64b] Beresford N. Parlett. Laguerre’s method applied to the matrix eigenvalue problem. Mathematics of Computation, 18(87):464–485, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1966:SIM

[Par66] Beresford N. Parlett. Singular and invariant matrices under the QR transformation (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):611–615, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1967:CDH

[Par67] Beresford Parlett. Canonical decomposition of Hessenberg matrices (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):223– 227, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parlett:1968:GCB

[Par68] Beresford N. Parlett. Global convergence of the basic QR algorithm on Hessenberg matrices. Mathematics of Computation, 22(104):803–817, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pathria:1962:SSR

[Pat62] R. K. Pathria. A statistical study of randomness among the first 10, 000 digits of π. Mathematics of Computation, 16(78): 188–197, April 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Patterson:1968:SGL

[Pat68a] T. N. L. Patterson. On some Gauss and Lobatto based integration formulae (in Technical Notes and Short Papers). Mathematics of Computation, 22(104): 877–881, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Patterson:1968:OAP

[Pat68b] T. N. L. Patterson. The optimum addition of points to quadrature formulae. Mathematics of Computation, 22(104): 847–856, October 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Patterson:1969:IFI

[Pat69] T. N. L. Patterson. Integration formulae involving derivatives (in Technical Notes and Short Papers). Mathematics of Computation, 23(106):411–412, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Paxson:1961:CTF

[Pax61] G. A. Paxson. The compositeness of the thirteenth Fermat number. Mathematics of Computation, 15(76):420, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pennell:1961:NSC

[PD61] M. M. Pennell and L. M. Delves. The numerical solution of coupled integrodifferential equations. Mathematics of Computation, 15(75):243–249, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Podolsky:1964:CMC

[PD64] Boris Podolsky and Harry H. Denman. Conditions on minimization criteria for smoothing. Mathematics of Computation, 18(87):441–448, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pearson:1963:C

[Pea63] Erna H. Pearson. On the congruences (p − 1)! ≡ −1 and 2^{p−1} ≡ (mod p^{2}). Mathematics of Computation, 17(82):

194–195, April 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pearson:1965:IEC

[Pea65] Carl E. Pearson. Impulsive end condition for diffusion equation. Mathematics of Computation, 19(92):570–576, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pearson:1969:ABS

[Pea69] Carl E. Pearson. Asymptotic behavior of solutions to the finite-difference wave equation. Mathematics of Computation, 23(108):711–715, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petryshyn:1963:GIM

[Pet63] W. V. Petryshyn. On a general iterative method for the approximate solution of linear operator equations. Mathematics of Computation, 17(81):1–10, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petryshyn:1965:EJS

[Pet65] W. V. Petryshyn. On the extrapolated Jacobi or simultaneous displacements method in the solution of matrix and operator equations. Mathematics of Computation, 19(89):37–55, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Perlin:1960:HPC

[PG60] I. E. Perlin and J. R. Garrett. High precision calculation of Arcsin x, Arccos x, and Arctan x (in Technical Notes and Short Papers). Mathematics of Computation, 14(71):270–274, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pine:1969:NTG

[Pin69] Milton R. Pine. Notes on the tables for Gaussian quadrature of N. M. Steen, G. D. Byrne and E. M. Gelbard. Mathematics of Computation, 23(107):673, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parker:1964:NPP

[PK64] E. T. Parker and R. B. Killgrove. A note on projective planes of order nine (in Technical Notes and Short Papers). Mathematics of Computation, 18(87): 506–508, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pearce:1962:FDM

[PM62] R. P. Pearce and A. R. Mitchell. On finite difference methods of solution of the transport equation. Mathematics of Computation, 16(78):155–169, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pope:1960:MAC

[Pop60] David A. Pope. A method of “alternating corrections” for the numerical solution of two-point boundary value problems. Mathematics of Computation, 14 (72):354–361, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Porsching:1963:MAA

[Por63] T. A. Porsching. Matrix assignments and an associated min max problem (in Technical Notes and Short Papers). Mathematics of Computation, 17(81): 81–84, January 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Preisendorfer:1963:BFE

[PR63] R. W. Preisendorfer and B. W. Roos. On the best fit to experimental data curves using homogeneous two-parameter functions. Mathematics of Computation, 17(83):231–236, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Price:1965:PAB

[Pri65] H. S. Price. A practical application of block diagonally dominant matrices (in Technical Notes and Short Papers). Mathematics of Computation, 19(90): 307–313, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Price:1968:MOM

[Pri68] Harvey S. Price. Monotone and oscillation matrices applied to finite difference approximations. Mathematics of Computation, 22(103):489–516, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Preisendorfer:1963:DMS

[PRS63] R. W. Preisendorfer, B. W. Roos, and W. C. Sangren. Determinants by means of solutions of linear equations (in Technical Notes and Short Papers). Mathematics of Computation, 17(83):285– 289, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pall:1960:PAG

[PS60] Gordon Pall and Esther Seiden. A problem in Abelian groups, with application to the transposition of a matrix on an electronic computer (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):189–192, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parkin:1967:DPP

[PS67] Thomas R. Parkin and Daniel Shanks.

On the distribution of parity in the partition function. Mathematics of Computation, 21(99):466–480, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Polachek:1963:TME

[PWMD63] H. Polachek, T. S. Walton, R. Mejia, and C. Dawson. Transient motion of an elastic cable immersed in a fluid. Mathematics of Computation, 17(81): 60–63, January 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1960:AWL

[Rab60]  Philip Rabinowitz.  Abscissas and

weights for Lobatto quadrature of high order. Mathematics of Computation, 14(69):47–52, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1968:EBG

[Rab68] Philip Rabinowitz. Error bounds in Gaussian integration of functions of loworder continuity (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):431–434, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Radok:1960:TES

[Rad60] J. R. M. Radok. Transcendental equation for the Schrödinger equation (in Technical Notes and Short Papers). Mathematics of Computation, 14(71): 276–278, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ralston:1962:RKM

[Ral62] Anthony Ralston. Runge–Kutta methods with minimum error bounds. Mathematics of Computation, 16(80):431–437, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Razelos:1965:NIR

[Raz65] P. Razelos. A new integral representation of the Bessel coefficients (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):130– 133, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riley:1967:NIV

[RBM67] James D. Riley, Morris M. Bennett, and Emily McCormick. Numerical integration of variational equations. Mathematics of Computation, 21(97):12–17, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ream:1961:NAC

[Rea61] N. Ream. Note on “Approximation of Curves by Line Segments”. Mathematics of Computation, 15(76):418–419, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1967:EES

[Rei67] Alex Reichel. Error estimates in simple quadrature with Voigt functions. Mathematics of Computation, 21(100):647– 651, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1969:IPV

[Rei69] Alex Reichel. The integral of the nth power of the Voigt function. Mathematics of Computation, 23(107):645– 649, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1964:CAE

[Ric64a] John R. Rice. On the computation of L_{1} approximations by exponentials, rationals, and other functions. Mathematics of Computation, 18(87):390–396, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1964:WA

[Ric64b] John R. Rice. On the L_{∞} Walsh arrays for Γ(x) and Erfc(x). Mathematics of Computation, 18(88):617–626, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1965:BRG

[Ric65] John R. Rice. Book review: Günter Meinardus, Approximation von Funktionen und ihre numerische Behandlung, Springer-Verlag, Berlin, 1964, viii + 180 pp., 23 cm. Price DM 49. Mathematics of Computation, 19(91): 517, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.org/
stable/2003711.

Rice:1966:EGS

[Ric66] John R. Rice. Experiments on Gram– Schmidt orthogonalization (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):325–328, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Richman:1969:CFF

[Ric69] P. L. Richman. Compressible fluid flow and the approximation of iterated integrals of a singular function. Mathematics of Computation, 23(106):355–372, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1962:AFA

[Rie62] H. Riesel. All factors q < 10^{8} in all Mersenne numbers 2^{p} − 1, pPrime < 10^{4} (in Technical Notes and Short Papers). Mathematics of Computation, 16(80): 478–482, October 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1963:FFN

[Rie63] Hans Riesel. A factor of the Fermat number F_{19} (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):458, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1964:NCT

[Rie64] Hans Riesel. Note on the congruence a^{p−1} ≡ 1(mod p^{2}) (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):149–150, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1969:LCP

[Rie69a] Hans Riesel. Lucasian criteria for the primality of N = h · 2^{n} − 1. Mathematics of Computation, 23(108):869– 875, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Riesel:1969:SFN

[Rie69b] Hans Riesel. Some factors of the numbers G_{n} = 6^{2^{n}} + 1 and H_{n} = 10^{2^{n}} + 1 (in Technical Notes and Short Papers). Mathematics of Computation, 23(106): 413–415, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rigler:1964:ACR

[Rig64] A. K. Rigler. An application of cyclic reduction to Ritz type difference equations (in Technical Notes and Short Papers). Mathematics of Computation, 18(86): 292–296, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rigler:1965:ESR

[Rig65] A. K. Rigler. Estimation of the successive over-relaxation factor (in Technical Notes and Short Papers). Mathematics of Computation, 19(90):302–307, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1964:AES

[Rob64] Raphael M. Robinson. Algebraic equations with span less than 4. Mathematics of Computation, 18(88):547–559, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Roberts:1965:CMT

[Rob65a] Jerry Allan Roberts. Computation of moments of K_{ν}(t)/I_{ν}(t) (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):651–654, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1965:SCA

[Rob65b] Raphael M. Robinson. Some conjectures about cyclotomic integers. Mathematics of Computation, 19(90):210–217, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Roman:1964:CQ

[Rom64] Irwin Roman. Controlled quadrature. Mathematics of Computation, 18(86): 254–263, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rose:1960:MCS

[Ros60] Milton E. Rose. A method for calculating solutions of parabolic equations with a free boundary. Mathematics of Computation, 14(71):249–256, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rose:1964:FDS

[Ros64] Milton E. Rose. Finite difference schemes for differential equations. Mathematics of Computation, 18(86):179– 195, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rotenberg:1960:CTH

[Rot60] A. Rotenberg. The calculation of toroidal harmonics (in Technical Notes and Short Papers). Mathematics of Computation, 14(71):274–276, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rothmann:1961:GQW

[Rot61] Harry A. Rothmann. Gaussian quadrature with weight function x^{n} on the interval (−1, 1). Mathematics of Computation, 15(74):163–168, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rodemich:1968:PTH

[RR68] Eugene R. Rodemich and Howard Rumsey, Jr. Primitive trinomials of high degree (in Technical Notes and Short Papers). Mathematics of Computation, 22 (104):863–865, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rabinowitz:1969:PST

[RR69] Philip Rabinowitz and Nira Richter.

Perfectly symmetric two-dimensional integration formulas with minimal numbers of points. Mathematics of Computation, 23(108):765–779, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Robinson:1961:ASI

[RS61] Stephen M. Robinson and A. H. Stroud.

The approximate solution of an integral equation using high-order Gaussian quadrature formulas (in Technical Notes and Short Papers). Mathematics of Computation, 15(75):286–288, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1968:LAE

[RU68] John R. Rice and Karl H. Usow. The Lawson algorithm and extensions. Mathematics of Computation, 22(101):118– 127, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rudnick:1966:NCF

[Rud66] Philip Rudnick. Note on the calculation of Fourier series (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):429–430, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Roberts:1966:CDS

[RW66] K. V. Roberts and N. O. Weiss. Convective difference schemes. Mathematics of Computation, 20(94):272–299, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

S:1964:BRT

[S.64] D. S. Book review: The Theory of Linear Graphs with Applications to the Theory of the Virial Development of the Properties of Gases by G. E. Uhlenbeck, G. W. Ford. Mathematics of Computation, 18(86):338, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842

                                              (electronic). URL https://www.jstor.
                                              org/stable/2003333.

Salzer:1960:AFO

[Sal60a] Herbert E. Salzer. Alternative formulas for osculatory and hyperosculatory inverse interpolation. Mathematics of Computation, 14(71):257–261, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1960:NSQ

[Sal60b] Herbert E. Salzer. A note on the solution of quartic equations (in Technical Notes and Short Papers). Mathematics of Computation, 14(71):279–281, July 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1962:MQC

[Sal62a] Herbert E. Salzer. Multiple quadrature with central differences on one line (in Technical Notes and Short Papers). Mathematics of Computation, 16(78): 244–248, April 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1962:NOR

[Sal62b] Herbert E. Salzer. Note on osculatory rational interpolation (in Technical Notes and Short Papers). Mathematics of Computation, 16(80):486–491, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1962:OPF

[Sal62c] Herbert E. Salzer. Optimum-point formulas for osculatory and hyperosculatory interpolation. Mathematics of Computation, 16(79):356–362, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sarma:1968:EMM

[Sar68] V. L. N. Sarma. Eberlein measure and mechanical quadrature formulae. I: Basic theory. Mathematics of Computation, 22(103):607–616, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sato:1966:CSC

[Sat66] Chikara Sato. Correction of stability curves in Hill–Meissner’s equation. Mathematics of Computation, 20(93): 98–106, January 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steen:1969:GQI

[SBG69] N. M. Steen, G. D. Byrne, and E. M. Gelbard. Gaussian quadratures for the integrals ∫_{0}^{∞} exp(−x^{2})f(x)dx and ∫_{0}^{b} exp(−x^{2})f(x)dx. Mathematics of Computation, 23(107):661–671, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shao:1964:TZG

[SCF64] T. S. Shao, T. C. Chen, and R. M. Frank.

Tables of zeros and Gaussian weights of certain associated Laguerre polynomials and the related generalized Hermite polynomials. Mathematics of Computation, 18(88):598–616, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schroder:1962:CEB

[Sch62] J. Schröder. Computing error bounds in solving linear systems. Mathematics of Computation, 16(79):323–337, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schinzel:1963:RPB

[Sch63] A. Schinzel. A remark on a paper of Bateman and Horn (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):445–447, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonheim:1967:CMN

[Sch67] J. Schonheim. Conversion of modular numbers to their mixed radix representation by matrix formula (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):253–257, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scraton:1969:SIE

[Scr69] R. E. Scraton. The solution of integral equations in Chebyshev series. Mathematics of Computation, 23(108):837– 844, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sternberg:1965:CRF

[SD65] H. M. Sternberg and J. B. Diaz. Computation of the Riemann function for the operator ∂^{n}/∂x_{1}∂x_{2} ⋯ ∂x_{n} + a(x_{1}, x_{2}, ⋯, x_{n}). Mathematics of Computation, 19(92):562–569, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schulz-DuBois:1969:IRA

[SD69] E. O. Schulz-DuBois. Integral relations among Bessel functions. Mathematics of Computation, 23(108):845–847, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Secrest:1965:BAI

[Sec65] Don Secrest. Best approximate integration formulas and best error bounds. Mathematics of Computation, 19(89): 79–83, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Senechalle:1964:ITM

[Sen64] Lester J. Senechalle. Iteration of triangular matrices (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):651–653, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sterbenz:1969:OSA

[SF69] P. H. Sterbenz and C. T. Fike. Optimal starting approximations for Newton’s method. Mathematics of Computation, 23(106):313–318, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schucany:1968:NAR

[SG68] W. R. Schucany and H. L. Gray. A new approximation related to the error function (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):201–202, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Selfridge:1964:FNM

[SH64] J. L. Selfridge and Alexander Hurwitz.

Fermat numbers and Mersenne numbers (in Technical Notes and Short Papers). Mathematics of Computation, 18(85): 146–148, January 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sprung:1965:GWO

[SH65] Donald W. L. Sprung and D. J. Hughes. Gauss weights and ordinates for ∫_{0}^{1} f(x)x^{2} dx (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):139–142, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1960:NGT

[Sha60a] Daniel Shanks. A note on Gaussian twin primes (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):201–203, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1960:CHL

[Sha60b] Daniel Shanks. On the conjecture of Hardy & Littlewood concerning the “Number of Primes of the Form n^{2} + a”. Mathematics of Computation, 14(72): 321–332, October 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1961:NF

[Sha61] Daniel Shanks. On numbers of the form n^{4} + 1. Mathematics of Computation, 15(74):186–189, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shapiro:1962:SRM

[Sha62] H. S. Shapiro. Some remarks on modular arithmetic and parallel computation (in Technical Notes and Short Papers). Mathematics of Computation, 16(78): 218–222, April 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1963:SDR

[Sha63a] Daniel Shanks. Supplementary data and remarks concerning a Hardy–Littlewood conjecture. Mathematics of Computation, 17(82):188–193, April 1963. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shapiro:1963:GES

[Sha63b] George Shapiro. Gauss elimination for singular matrices (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):441–445, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1964:MGB

[Sha64a] Daniel Shanks. On maximal gaps between successive primes (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):646–651, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1964:PDS

[Sha64b] Daniel Shanks. Polylogarithms, Dirichlet series, and certain constants (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):322– 324, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1964:SOT

[Sha64c] Daniel Shanks. The second-order term in the asymptotic expansion of B(x). Mathematics of Computation, 18(85): 75–86, January 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1965:QFU

[Sha65a] Lawrence F. Shampine. Quadrature formulas using derivatives (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):481–482, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1965:ROM

[Sha65b] Daniel Shanks. Review of: Marshall Hall, Jr. and James K. Senior, “The Groups of Order 2^{n}(n ≤ 6)”. Mathematics of Computation, 19(90):335– 337, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1965:RPB

[Sha65c] Daniel Shanks. Review of Paul T. Bateman and Roger A. Horn, “Primes Represented by irreducible polynomials in one variable”, in Theory of Numbers, pp. 119–132, in particular Tables II– V. Mathematics of Computation, 19(92): 684–686, October 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1966:SDE

[Sha66] E. Baylis Shanks. Solutions of differential equations by evaluations of functions. Mathematics of Computation, 20(93):21–38, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1967:GEC

[Sha67a] Daniel Shanks. Generalized Euler and class numbers. Mathematics of Computation, 21(100):689–694, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1967:LCG

[Sha67b] Daniel Shanks. Lal’s constant and generalizations (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):705–707, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1969:GCN

[Sha69] Daniel Shanks. On Gauss’s class number problems. Mathematics of Computation, 23(105):151–163, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shisha:1960:CPB

[Shi60] O. Shisha. The Chebyshev polynomial of best approximation to a given function on an interval. Mathematics of Computation, 14(94):266–271, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shisha:1966:CPB

[Shi66] O. Shisha. The Chebyshev polynomial of best approximation to a given function on an interval. Mathematics of Computation, 20(94):266–271, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shortley:1960:BRJ

[Sho60] George Shortley. Book review: The 3j and 6-j Symbols, by Manuel Rotenberg, R. Bivins, N. Metropolis, and John K. Wooten, Jr. Mathematics of Computation, 14(72):382–383, 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://links.jstor.
org/sici?sici=0025-5718%28196010%
2914%3A72%3C382%3ATAS%3E2.0.CO%3B2-

Siegel:1965:DCM

[Sie65] Irving H. Siegel. Deferment of computation in the method of least squares (in Technical Notes and Short Papers). Mathematics of Computation, 19(90): 329–331, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Simon:1965:NTS

[Sim65] W. E. Simon. Numerical technique for solution and error estimate for the initial value problem. Mathematics of Computation, 19(91):387–393, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Simpson:1967:FSB

[Sim67] R. Bruce Simpson. A fundamental solution for a biharmonic finite-difference operator. Mathematics of Computation, 21(99):321–339, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1961:IFC

[SK61] Herbert E. Salzer and Genevieve M. Kimbro. Improved formulas for complete and partial summation of certain series. Mathematics of Computation, 15(73): 23–39, January 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1967:DMD

[SK67] Daniel Shanks and Sidney Kravitz. On the distribution of Mersenne divisors (in technical notes and short notices). Mathematics of Computation, 21(97): 97–101, January 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1961:TWC

[SL61] Herbert E. Salzer and Norman Levine.

Table of a Weierstrass continuous nondifferentiable function. Mathematics of Computation, 15(74):120–130, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salzer:1968:PEI

[SL68] Herbert E. Salzer and Norman Levine.

Proof that every integer ≤ 452, 479, 659 is a sum of five numbers of the form 𝒬_{x} ≡ (x^{3} + 5x)/6, x ≥ 0 (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):191– 192, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1969:AHT

[SL69] W. E. Smith and J. N. Lyness. Applications of Hilbert transform theory to numerical quadrature. Mathematics of Computation, 23(106):231–252, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sankar:1963:NRM

[SM63] R. Sankar and V. Malini. A note on the relative merits of Padé and Maehly’s diagonal convergents in computing e^{x}. Mathematics of Computation, 17(84): 414–418, October 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1965:ASO

[Smi65] Francis J. Smith. An algorithm for summing orthogonal polynomial series and their derivatives with applications to curve-fitting and interpolation. Mathematics of Computation, 19(89):33– 36, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scheffler:1960:NEE

[SO60] D. Scheffler and R. Ondrejka. The numerical evaluation of the eighteenth perfect number (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):199–200, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Solomon:1966:SRS

[Sol66] Alan Solomon. Some remarks on the Stefan problem. Mathematics of Computation, 20(95):347–360, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Solomon:1967:SSP

[Sol67] A. Solomon. A steady state phase change problem. Mathematics of Computation, 21(99):355–359, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spielberg:1961:ECF

[Spi61] Kurt Spielberg. Efficient continued fraction approximations to elementary functions. Mathematics of Computation, 15 (76):409–417, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spielberg:1962:PCF

[Spi62]  Kurt Spielberg.  Polynomial and

continued-fraction approximations for logarithmic functions. Mathematics of Computation, 16(78):205–217, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1966:ZSZ

[Spi66] Robert Spira. Zeros of sections of the zeta function. I. Mathematics of Computation, 20(96):542–550, October 1966.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1967:ZAF

[Spi67] Robert Spira. Zeros of approximate functional approximations. Mathematics of Computation, 21(97):41–48, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1968:ZSZ

[Spi68] Robert Spira. Zeros of sections of the zeta function. II. Mathematics of Computation, 22(101):163–173, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1969:CDF

[Spi69] Robert Spira. Calculation of Dirichlet L-functions. Mathematics of Computation, 23(107):489–497, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spohn:1969:LC

[Spo69] W. G. Spohn. On the lattice constant for |x^{3} + y^{3} + z^{3}| ≤ 1. Mathematics of Computation, 23(105):141–149, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Srivastava:1969:EHH

[Sri69] H. M. Srivastava. An extension of the Hille–Hardy formula. Mathematics of Computation, 23(106):305–311, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stancu:1963:QFS

[SS63a] D. D. Stancu and A. H. Stroud. Quadrature formulas with simple Gaussian nodes and multiple fixed nodes. Mathematics of Computation, 17(84):384–394, October 1963. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1963:AIF

[SS63b] A. H. Stroud and Don Secrest. Approximate integration formulas for certain spherically symmetric regions. Mathematics of Computation, 17(82):105– 135, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sag:1964:NEH

[SS64] T. W. Sag and G. Szekeres. Numerical evaluation of high-dimensional integrals. Mathematics of Computation, 18(86): 245–253, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stein:1965:ERA

[SS65] M. L. Stein and P. R. Stein. Experimental results on additive 2-bases. Mathematics of Computation, 19(91): 427–434, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1966:VTL

[SS66] Daniel Shanks and Larry P. Schmid. Variations on a theorem of landau. part I. Mathematics of Computation, 20(96): 551–569, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spitznagel:1968:CSO

[SS68] Edward L. Spitznagel, Jr. and Stephen A. Szygenda. A computer study of the orders of finite simple groups (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):669–671, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sarma:1969:EMM

[SS69] V. L. N. Sarma and A. H. Stroud. Eberlein measure and mechanical quadrature formulae. II. numerical results. Mathematics of Computation, 23(108):781– 784, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stalley:1962:ASC

[Sta62] Robert D. Stalley. An algorithm for solving certain polynomial equations with coefficient parameters (in Technical Notes and Short Papers). Mathematics of Computation, 16(77):106–109, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stancu:1963:ERT

[Sta63] D. D. Stancu. Evaluation of the remainder term in approximation formulas by Bernstein polynomials. Mathematics of Computation, 17(83):270–278, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Steck:1964:DPS

[Ste64a] Richard Steck. A dynamic programming strategy for the two machine problem. Mathematics of Computation, 18(86): 285–291, April 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stemmler:1964:IWT

[Ste64b] Rosemarie M. Stemmler. The ideal Waring theorem for exponents 401–200, 000 (in Technical Notes and Short Papers). Mathematics of Computation, 18(85): 144–146, January 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stetter:1965:SPW

[Ste65] Hans J. Stetter. Stabilizing predictors for weakly unstable correctors. Mathematics of Computation, 19(89): 84–89, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stetter:1968:EBG

[Ste68a] Franz Stetter. Error bounds for Gauss– Chebyshev quadrature (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):657–659, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stetter:1968:GMR

[Ste68b] Franz Stetter. On a generalization of the midpoint rule (in Technical Notes and Short Papers). Mathematics of Computation, 22(103):661–663, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stengle:1969:UAS

[Ste69a] Gilbert Stengle. Uniform asymptotic solution of second order linear differential equations without turning varieties. Mathematics of Computation, 23(105): 1–22, January 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stewart:1969:LMF

[Ste69b] G. W. Stewart, III. On Lehmer’s method for finding the zeros of a polynomial. Mathematics of Computation, 23(108):829–835, s24–s30, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Ste71].

Stewart:1971:CLM

[Ste71] G. W. Stewart, III. Corrigendum: “On Lehmer’s Method for Finding the Zeros of a Polynomial”. Mathematics of Computation, 25(113):203, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Ste69b].

Stone:1961:ACL

[Sto61] Henry Stone. Approximation of curves by line segments. Mathematics of Computation, 15(73):40–47, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1960:NIF

[Str60a] A. H. Stroud. Numerical integration formulas of degree two. Mathematics of Computation, 14(69):21–26, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Struble:1960:TUQ

[Str60b] George Struble. Tables for use in quadrature formulas involving derivatives of the integrand. Mathematics of Computation, 14(69):8–12, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1961:BAI

[Str61a] A. H. Stroud. A bibliography on approximate integration (in surveys and expository papers). Mathematics of Computation, 15(73):52–80, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1961:NIF

[Str61b] A. H. Stroud. Numerical integration formulas of degree 3 for product regions and cones. Mathematics of Computation, 15(74):143–150, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1963:CQF

[Str63] A. H. Stroud. Coefficients in quadrature formulas (in Technical Notes and Short Papers). Mathematics of Computation, 17(83):289–291, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1964:AIF

[Str64] A. H. Stroud. Approximate integration formulas of degree 3 for simplexes. Mathematics of Computation, 18(88): 590–597, October 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Strand:1965:MCE

[Str65] Otto Neall Strand. A method for the computation of the error function of a complex variable (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):127–129, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1966:SFD

[Str66] A. H. Stroud. Some fifth degree integration formulas for symmetric regions. Mathematics of Computation, 20(93): 90–97, January 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Strecok:1968:CIE

[Str68a] Anthony J. Strecok. On the calculation of the inverse of the error function. Mathematics of Computation, 22(101): 144–158, January 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stroud:1968:ESI

[Str68b] A. H. Stroud. Extensions of symmetric integration formulas. Mathematics of Computation, 22(102):271–274, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shanks:1962:CD

[SW62] Daniel Shanks and John W. Wrench, Jr. Calculation of π to 100,000 decimals. Mathematics of Computation, 16(77):76–99, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.jstor.org/stable/2003813. A

note added in proof says: “J. M. Gerard of IBM United Kingdom Limited, who was then unaware of the computation described above, computed π to 20,000 D on the 7090 in the London Data Centre on July 31, 1961. His program used Machin’s formula, (1) [π = 16 arctan(1/5) − 4 arctan(1/239)], and required 39 minutes running time. His result agrees with ours to that number of decimals.”.

Shanks:1963:CCD

[SW63] Daniel Shanks and John W. Wrench, Jr. The calculation of certain Dirichlet series. Mathematics of Computation, 17(82):136–154, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Shampine:1969:BIO

[SW69a] L. F. Shampine and H. A. Watts. Block implicit one-step methods. Mathematics of Computation, 23(108):731–740, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swartz:1969:GFD

[SW69b] Blair Swartz and Burton Wendroff. Generalized finite-difference schemes. Mathematics of Computation, 23(105):37–49, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swartz:1968:EBE

[Swa68] Blair K. Swartz. Explicit O(h^{2}) bounds on the eigenvalues of the half-L. Mathematics of Computation, 22(101):40–59, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sweeney:1963:CEC

[Swe63] Dura W. Sweeney. On the computation of Euler’s constant. Mathematics of Computation, 17(82):170–178, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Swift:1960:CGF

[Swi60] J. D. Swift. Construction of Galois fields of characteristic two and irreducible polynomials. Mathematics of Computation, 14(70):99–103, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tait:1967:EAR

[Tai67] R. Tait. Error analysis of recurrence equations. Mathematics of Computation, 21(100):629–638, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Takenaga:1966:EIG

[Tak66] Roy Takenaga. On the evaluation of the incomplete gamma function (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):606–610, October 1966. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic).

Tamir:1962:CEM

[Tam62] T. Tamir. Characteristic exponents of Mathieu functions (in Technical Notes and Short Papers). Mathematics of Computation, 16(77):100–106, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tausworthe:1965:RNG

[Tau65] Robert C. Tausworthe. Random numbers generated by linear recurrence modulo two. Mathematics of Computation, 19(90):201–209, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2003345.

Thacher:1961:ISR

[Tha61] Henry C. Thacher, Jr. Iterated square root expansions for the inverse cosine and inverse hyperbolic cosine. Mathematics of Computation, 15(76):399–403, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thompson:1965:AEI

[Tho65] G. T. Thompson. The asymptotic expansion of the integrals psi and chi in terms of Tchebycheff polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 19(92): 661–663, October 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thompson:1966:ESI

[Tho66] Rory Thompson. Evaluation of I_{n}(b) = 2π^{−1} ∫_{0}^{∞} (\frac{sinx}{x})^{n} cos(bx)dx and of similar integrals (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):330–332, April 1966.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thuraisamy:1969:ASM

[Thu69a] V. Thuraisamy. Approximate solutions for mixed boundary value problems by finite-difference methods. Mathematics of Computation, 23(106):373–386, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thuraisamy:1969:MTD

[Thu69b] V. Thuraisamy. Monotone type discrete analogue for the mixed boundary value problem. Mathematics of Computation, 23(106):387–394, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tooper:1968:SCP

[TM68] Robert F. Tooper and John Mark. Simplified calculation of ei(x) for positive arguments, and a short table of Shi(x) (in Technical Notes and Short Papers). Mathematics of Computation, 22(102): 448–449, April 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Traub:1965:GSA

[Tra65] J. F. Traub. Generalized sequences with applications to the discrete calculus. Mathematics of Computation, 19(90): 177–200, April 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Traub:1966:CGC

[Tra66] J. F. Traub. A class of globally convergent iteration functions for the solution of polynomial equations. Mathematics of Computation, 20(93):113–138, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Treanor:1966:MNI

[Tre66] Charles E. Treanor. A method for the numerical integration of coupled firstorder differential equations with greatly different time constants. Mathematics of Computation, 20(93):39–45, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Theilheimer:1961:FSL

[TS61] Feodor Theilheimer and William Starkweather. The fairing of ship lines on a high-speed computer. Mathematics of Computation, 15(76):338–355, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuck:1967:SFT

[Tuc67]  E. O. Tuck.  A simple “Filon-

trapezoidal” rule (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):239–241, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tung:1966:BDK

[Tun66] Shih-Hsiung Tung. Boundedness of difference kernels of Bessel and Fourier series (in Technical Notes and Short Papers). Mathematics of Computation, 20 (93):157–163, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Uchiyama:1966:NTJ

[Uch66] Saburo Uchiyama. A note on a theorem of J. N. Franklin (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):139–140, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Valentin:1968:UHI

[Val68] Richard A. Valentin. The use of the hypercircle inequality in deriving a class of numerical approximation rules for analytic functions. Mathematics of Computation, 22(101):110–117, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tuyl:1964:ESD

[Van64] A. H. Van Tuyl. The evaluation of some definite integrals involving Bessel functions which occur in hydrodynamics and elasticity. Mathematics of Computation, 18(87):421–432, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VandeVel:1969:SEM

[Van69] H. Van de Vel. On the series expansion method for computing incomplete elliptic integrals of the first and second kinds. Mathematics of Computation, 23(105): 61–69, January 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1966:CPA

[Var66] J. M. Varah. Certification of Parlett’s ALGOL eigenvalue procedure Eig 3 (in Technical Notes and Short Papers). Mathematics of Computation, 20(95): 437–438, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1968:CEG

[Var68a] J. M. Varah. The calculation of the eigenvectors of a general complex matrix by inverse iteration. Mathematics of Computation, 22(104):785–791, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Varah:1968:RMB

[Var68b] J. M. Varah. Rigorous machine bounds for the eigensystem of a general complex matrix. Mathematics of Computation, 22(104):793–801, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1965:CEF

[Ver65] Arun Verma. A class of expansions of G-functions and the Laplace transform (in Technical Notes and Short Papers). Mathematics of Computation, 19(92): 664–666, October 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1966:CEB

[Ver66a] Arun Verma. Certain expansions of the basic hypergeometric functions (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):151–157, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1966:EIH

[Ver66b] Arun Verma. Expansions involving hypergeometric functions of two variables. Mathematics of Computation, 20(96): 590–596, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1966:NEH

[Ver66c] Arun Verma. A note on an expansion of hypergeometric functions of two variables. Mathematics of Computation, 20(95):413–417, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1967:NEI

[Ver67a] Arun Verma. A note on expansions involving Meijer’s G-functions (in technical notes and short notices). Mathematics of Computation, 21(97):107–112, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1967:NSG

[Ver67b] Arun Verma. A note on the summation of the generalised hypergeometric functions (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):232–236, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vonHoldt:1963:DIP

[vH63] R. E. von Holdt and R. J. Howerton. The definite integral of the product of linear functions. Mathematics of Computation, 17(84):419–425, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Voigt:1966:NSI

[Voi66] Robert G. Voigt. On a numerical solution of an integral equation with singularities (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):163–166, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vonRosenberg:1967:NSF

[von67] Dale U. von Rosenberg. Numerical solution for flux components in potential flow. Mathematics of Computation, 21 (100):620–628, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vandeVooren:1966:NCI

[vv66] A. I. van de Vooren and H. J. van Linde. Numerical calculation of integrals with strongly oscillating integrand. Mathematics of Computation, 20(94): 232–245, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vandeLune:1969:NSD

[vW69] J. van de Lune and E. Wattel. On the numerical solution of a differentialdifference equation arising in analytic number theory (in Technical Notes and Short Papers). Mathematics of Computation, 23(106):417–421, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wadsworth:1964:IAE

[vZW64] D. van Z. Wadsworth. Improved asymptotic expansion for the error function with imaginary argument (in Technical Notes and Short Papers). Mathematics of Computation, 18(88):662–664, October 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wadsworth:1965:IAE

[vZW65] Donald van Zelm Wadsworth. Improved asymptotic expansion for the exponential integral with positive argument (in Technical Notes and Short Papers). Mathematics of Computation, 19(90):327–328, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wagstaff:1967:SIN

[Wag67] S. S. Wagstaff, Jr. On sequences of integers with no 4, or no 5 numbers in arithmetical progression. Mathematics of Computation, 21(100):695–699, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walters:1969:NME

[Wal69] Jerry A. Walters. Nonnegative matrix equations having positive solutions. Mathematics of Computation, 23(108): 827, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watson:1962:PPM

[Wat62a] E. J. Watson. Primitive polynomials (mod 2). Mathematics of Computation, 16(79):368–369, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Watson:1962:PPT

[Wat62b] E. J. Watson. Primitive polynomials (Mod 2) (in Technical Notes and Short Papers). Mathematics of Computation, 16(79):368–369, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

White:1964:CMC

[WB64] Paul A. White and Robert R. Brown. A comparison of methods for computing the eigenvalues and eigenvectors of a real symmetric matrix. Mathematics of Computation, 18(87):457–463, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werner:1961:CAG

[WC61] Helmut Werner and Robert Collinge. Chebyshev approximations to the gamma function (in Technical Notes and Short Papers). Mathematics of Computation, 15(74):195–197, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weingarten:1961:TGC

[WD61] Harry Weingarten and A. R. Di Donato.

A table of generalized circular error. Mathematics of Computation, 15(74): 169–173, April 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wells:1961:GPT

[Wel61] Mark B. Wells. Generation of permutations by transposition (in Technical Notes and Short Papers). Mathematics of Computation, 15(74):192–195, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wendroff:1965:BES

[Wen65] Burton Wendroff. Bounds for eigenvalues of some differential operators by the Rayleigh–Ritz method. Mathematics of Computation, 19(90):218–224, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walmsley:1960:FTA

[WG60] Charles Walmsley and Arthur S. G.

Grant. The Fourier transform of [c^{2k} + (x − x^{−1})^{2k}]^{−1} arising from study of tuned circuit spectra (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):193–195, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wang:1964:RAR

[WG64] H. H. Wang and R. T. Gregory. On the reduction of an arbitrary real square matrix to tridiagonal form (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):501–505, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1965:SCP

[WGZ65] H. C. Williams, R. A. German, and C. R. Zarnke. Solution of the Cattle Problem of Archimedes (in Technical Notes and Short Papers). Mathematics of Computation, 19(92):671–674, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Whittlesey:1963:IGF

[Whi63] John R. B. Whittlesey. Incomplete gamma functions for evaluating Erlang process probabilities. Mathematics of Computation, 17(81):11–17, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Whitney:1965:EWG

[Whi65] E. L. Whitney. Estimates of weights in Gauss-type quadrature. Mathematics of Computation, 19(90):277–286, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Widlund:1965:SPD

[Wid65] Olof B. Widlund. On the stability of parabolic difference schemes. Mathematics of Computation, 19(89):1–13, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                              URL http://www.jstor.org/stable/
                                              2004090.

Widlund:1966:RCA

[Wid66] Olof B. Widlund. On the rate of convergence of an alternating direction implicit method in a noncommutative case. Mathematics of Computation, 20(96): 500–515, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wigley:1969:MSS

[Wig69] Neil M. Wigley. On a method to subtract off a singularity at a corner for the Dirichlet or Neumann problem. Mathematics of Computation, 23(106):395– 401, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilf:1963:CRC

[Wil63] Herbert S. Wilf. Calculations relating to a conjecture of Polya and Schoenberg. Mathematics of Computation, 17(82): 200–201, April 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilcox:1964:ACR

[Wil64] W. R. Wilcox. Asymptotic computation of the repeated integrals of the error function complement. Mathematics of Computation, 18(85):98–105, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Willner:1967:EMC

[Wil67] Leopold B. Willner. An elimination method for computing the generalized inverse (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):227–229, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilcox:1968:ZTN

[Wil68] Peter H. Wilcox. The zeros of P_{ν}^{1}(cos θ)

and \frac{∂}{∂θ} P_{μ}^{1}(cos θ) (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):205–208, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilson:1969:GAN

[Wil69] M. Wayne Wilson. A general algorithm for nonnegative quadrature formulas. Mathematics of Computation, 23(106):253–258, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1961:PAI

[Wim61] Jet Wimp. Polynomial approximations to integral transforms. Mathematics of Computation, 15(74):174–178, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1962:PEB

[Wim62] Jet Wimp. Polynomial expansions of Bessel functions and some associated functions. Mathematics of Computation, 16(80):446–458, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Wim72].

Wimp:1967:ARC

[Wim67] Jet Wimp. The asymptotic representation of a class of G-functions for large parameter. Mathematics of Computation, 21(100):639–646, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1968:RFH

[Wim68] Jet Wimp. Recursion formulae of hypergeometric functions. Mathematics of Computation, 22(102):363–373, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1972:CPE

[Wim72] Jet Wimp. Corrigendum: Polynomial expansions of Bessel functions and some associated functions (Math. Comp. 16 (1962), 446–458). Mathematics of Computation, 26(117):309, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/journals/
mcom/1972-26-117/S0025-5718-1972-
0400659-X;   http://www.ams.org/

journals/mcom/1972-26-117/S0025-
5718-1972-0400659-X/S0025-5718-1972-
0400659-X.pdf; https://www.ams.org/
mathscinet-getitem?mr=400659;   https:
//www.ams.org/mathscinet/search/
authors.html?authorName=Wimp%2C%

20Jet. See [Wim62].

Witty:1964:NMN

[Wit64] W. H. Witty. A new method of numerical integration of differential equations (in Technical Notes and Short Papers). Mathematics of Computation, 18(87): 497–500, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1966:DBI

[WKG66] Van E. Wood, R. P. Kenan, and M. L.

Glasser. Doppler broadening integrals (in Technical Notes and Short Papers). Mathematics of Computation, 20(96): 610–611, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wells:1965:NSC

[WM65] Jacqueline Wells and Joseph Muskat.

On the number of solutions of certain trinomial congruences (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):483–487, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weil:1967:ZTN

[WMR67] Joyce Weil, Tadepalli S. Murty, and Desiraju B. Rao. Zeros of J_{n}(λ)Y_{n}(ηλ) − J_{n}(ηλ)Y_{n}(λ) (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):722–727, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1966:SIR

[Woo66] Van E. Wood. Some integrals of Ramanujan and related contour integrals (in Technical Notes and Short Papers). Mathematics of Computation, 20(95): 424–429, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1967:CEI

[Woo67] Van E. Wood. Chebyshev expansions for integrals of the error function (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):494– 496, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1968:ECC

[Woo68] Van E. Wood. Efficient calculation of Clausen’s integral (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):883–884, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Walton:1960:CTM

[WP60] Thomas S. Walton and Harry Polachek. Calculation of transient motion of submerged cables. Mathematics of Computation, 14(69):27–46, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werner:1963:AFI

[WR63] H. Werner and G. Raymann. An approximation to the Fermi integral F_{1/2}(x). Mathematics of Computation, 17(82): 193–194, April 1963. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrathall:1964:NFF

[Wra64] Claude P. Wrathall. New factors of Fermat numbers (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):324–325, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1960:FEK

[Wre60] John W. Wrench, Jr. Further evaluation of Khintchine’s constant (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):370–371, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1961:EAC

[Wre61] John W. Wrench, Jr. Evaluation of Artin’s constant and the twin-prime constant. Mathematics of Computation, 15 (76):396–398, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1968:CTS

[Wre68] John W. Wrench, Jr. Concerning two series for the gamma function. Mathematics of Computation, 22(103):617– 626, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1966:QCK

[WS66] John W. Wrench, Jr. and Daniel Shanks.

Questions concerning Khintchine’s constant and the efficient computation of regular continued fractions (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):444–448, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wunderlich:1962:CPP

[Wun62] M. Wunderlich. Certain properties of pyramidal and figurate numbers (in Technical Notes and Short Papers). Mathematics of Computation, 16(80): 482–486, October 1962. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wunderlich:1963:NEF

[Wun63] M. Wunderlich. On the non-existence of Fibonacci squares (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):455–457, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wynn:1960:RAF

[Wyn60] Peter Wynn. The rational approximation of functions which are formally defined by a power series expansion. Mathematics of Computation, 14(70): 147–186, April 1960. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wynn:1961:EAO

[Wyn61] P. Wynn. The epsilon algorithm and operational formulas of numerical analysis. Mathematics of Computation, 15(74): 151–158, April 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wynn:1962:ATI

[Wyn62] P. Wynn. Acceleration techniques for iterated vector and matrix problems. Mathematics of Computation, 16(79): 301–322, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1968:RPN

[WZ68] H. C. Williams and C. R. Zarnke. A report on prime numbers of the forms M = (6a + 1)2^{2m−1} − 1 and M′ = (6a − 1)2^{2m} − 1 (in Technical Notes and Short Papers). Mathematics of Computation, 22(102):420–422, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yamaguti:1967:SRL

[Yam67] Masaya Yamaguti. Some remarks on the Lax–Wendroff finite-difference scheme for nonsymmetric hyperbolic systems. Mathematics of Computation, 21(100): 611–619, October 1967. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1965:NMV

[Yan65] C. H. Yang. A note on the maximum value of determinants over the complex field (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):455–456, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1966:SDM

[Yan66] C. H. Yang. Some designs for maximal (+1, −1)-determinant of order n ≡ 2(mod 4) (in Technical Notes and Short Papers). Mathematics of Computation, 20(93):147–148, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1968:DMM

[Yan68] C. H. Yang. On designs of maximal (+1, −1)-matrices of order n ≡ 2(mod 4). Mathematics of Computation, 22(101):174–180, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yang:1969:DMM

[Yan69] C. H. Yang. On designs of maximal (+1, −1)-matrices of order n ≡ 2(mod 4). II (in Technical Notes and Short Papers). Mathematics of Computation, 23(105):201–205, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Yamabe:1961:CAF

[YP61] Hidehiko Yamabe and David Pope. A computational approach to the fourcolor problem. Mathematics of Computation, 15(75):250–253, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Zassenhaus:1969:PH

[ZL69] H. Zassenhaus and J. Liang. On a problem of hasse. Mathematics of Computation, 23(107):515–519, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).